A Complete Bibliography of Publications in Foundations of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 February 2024 Version 1.13

Title word cross-reference

3 [797]. 4n + 1 [776]. 4π [645]. Aπαταω [734]. δ [437]. E = mc^{2} [525]. k [437]. n [961]. Z [793].

-Animal-Technology-Human- [449]. -Model [437]. -Party [797]. -Quantum [437]. -Space [793]. -Symmetry [645].

17th [327]. 19 [1007].

20th [35].

322 [851]. 3PKE [797]. 3PKE-NEMO [797].

’97 [34].

Abduction [489, 117, 118, 229, 247, 234, 231, 425, 932, 233, 954]. Abductive [301, 303, 724, 306, 172, 511]. Absence [226]. Abstract [995, 745, 882]. Abstraction [245, 527, 542, 647]. Academies [840]. Acceptance [101]. Accepting [924]. Access [510]. According [325]. Account [249, 343, 854]. Accusations [907]. Acknowledgements [73]. Acquire [930]. Acquisition [233, 191]. Act [772]. Action [374]. activity [94]. Actor [930]. Acts [923]. Actuality [620]. Actually [699]. adaptationist [761]. Adaptive [489, 165, 119]. Adequately [760]. Adolescents [909]. Adults [909]. Advances [847]. Advantages [102]. Adversarial [990]. Adversus [459]. Advertising [322]. Aerts [77]. Aesthetic [679]. Aesthetics [633, 673, 706, 828]. Affective [386]. AFOS [11, 38, 10, 20, 36, 58]. Afraid [467]. After [876, 1007]. Again [723, 871, 462]. Against [146, 818, 862]. Age [962, 625, 626]. Agency [138, 965, 624]. agent [16]. Agents [1006, 748]. Aggregativity [145]. Agnostic [372]. Agonism [942]. Agonistic [895]. Aharonov [538]. AI [951]. aiming [118]. Akkerhuis [393]. Albert [910]. Albinism [908, 904]. Alchemist [51]. Alcohol [906]. Alexander [696]. Alfred [658]. Algebra [313]. Algebraic [237, 504]. Algorithimic [1001]. Algorithm [792]. Algorithmic [940]. Algorithms [1001]. Aliquid} [963]. All-embracing [221]. Alleged [598]. Allies [304]. Already [463]. Also [913]. Alterity [894]. Alternative [716, 212, 835, 981, 965]. Alternatives [139]. Always [595]. Ambiguities [424]. Ambiguity [809]. America [1015]. Amir [696]. Among [350, 810]. ampliative [119]. Analogical [116, 248, 657]. Analogies [506, 600, 370]. analogue [1003]. Analysing [907]. Analysis [87, 438, 610, 962, 987, 424, 1021, 958, 483, 1013, 838, 372, 1014, 891, 598, 713, 905, 412, 645, 879, 266, 988]. Anamnesis [675]. Ancient [901]. Andrzej [55]. Anesthetics [112]. ANFIS [801]. Animal [449, 188]. Animals [425]. Anomaly [315]. Anselm [739]. Answer [865]. Answers [26]. Antenna [863]. Anthropic [194, 217, 369]. Anthropocene [921, 888, 997, 868, 866, 871, 899, 873]. Anthropocentrism [997]. Anthropomorphic [359]. Anthropomorphism [994]. Anti [277]. Anti-philosophical [277]. Anticipative [349]. Antimatter [814, 946]. Antinomicity [98]. Antique [1013]. Antirealism [972]. Antirealists [644]. Anywhere [474]. Aphorisms [273]. Apostelian [90]. Apparent [272]. Apparitions [452]. Applicability [702, 490]. Application [943, 360, 343, 503]. Applications [5, 365, 357, 809]. Applied [41, 761, 839, 294]. Approach [805, 310, 992, 510, 949, 267, 170, 596, 729, 790, 360, 740, 1018, 778, 816, 153, 799, 90, 1]. Appropriation [825]. Appropriations [675]. Approximation [677, 118]. April [926]. Apriorics [756]. Architect [295]. Architectural [388]. Architecture [138, 454, 455]. Area [162]. Argentinian [825]. Argument [146, 207, 846, 335, 279, 818, 763]. Argumentation [336, 334]. Arguments [972, 333, 337, 331]. Argumentum [739]. Aristotelian [882, 324, 360]. Aristotelism [318]. Aristotle [322, 325, 323, 316, 321, 319, 320, 317, 327]. Arithmetic [535]. Arithmetics [807]. Arithmetization [570]. Arrow [291, 286]. Art [322, 633, 630, 729, 626, 753, 621, 622, 616, 617, 444, 951, 635]. Artefacts [1020]. article [82, 77]. Artifacts [736, 308]. Artifactual [876]. Artificial [973, 1006, 828, 51, 486, 409, 410, 165]. Artistic [635]. ArtScience [815]. Ascriptions [731]. Aspects [355, 993]. Assemblage [665]. Asserting [713]. Assess [801]. Assessment [1015, 822]. Association [37, 793, 38, 10, 20, 36, 58]. Associations [798]. Astronomy [346]. Asymmetry [774]. Atechnogenesis [609]. Atom [912, 155]. Attack [362]. Attacker [801]. Attempt [840]. Attention [376]. Attitudes [809]. Attractors [247]. August [926]. Authors [157, 174]. Autonomous [675, 16]. Autonomy [625]. Autopoietic [605]. Availability [420]. Awareness [153]. Away [897]. Axiomatic [85]. Axiomatization [400, 965, 681, 758]. Axiomatize [649].

Back [878]. Background [115, 1023]. Baconian [835, 477]. Barriers [341]. Based [716, 864, 229, 794, 511, 797, 713, 191, 243, 360, 657, 450]. Basic [270]. Basis [590, 518]. Bayesian [979, 956]. Bayesianism [986]. Be [403, 644, 978, 1001, 344, 896, 225, 727]. Beauty [550, 45]. Becoming [436, 548]. Bedford [285]. Been [730]. Before [690, 867, 889]. Behavior [501, 749]. Being [948, 559, 581, 578, 384, 473]. Being-with [581]. beings [365]. Belief [737, 572, 821, 566, 827, 569, 568, 571, 565, 860]. Beliefs [567, 50, 296, 570]. Bell [683, 789, 848]. Bell-Type [789, 848]. Belt [124, 1023]. Bendectin [557]. Bending [97]. Bergson [730]. Bernard [926]. Bernhard [912]. Bertini [864]. Bertolt [452]. Best [986, 373, 954, 128]. Better [634, 672, 527, 542, 647]. Between [881, 707, 1001, 999, 137, 908, 381, 1022, 959, 463, 436, 291, 771, 712, 334, 933, 453, 120, 388, 856, 886, 563, 938, 380, 940, 442, 377, 378, 385]. Beyond [428, 997, 707, 180, 510, 532, 377, 378, 468, 526, 698]. Bialkowski [75, 73]. Bias [217]. Biases [831, 145]. Bicentenary [832]. Big [643, 609, 871, 792, 940]. Bijective [552]. Bilinear [797]. Biographical [9, 19, 56, 71, 106, 114, 135, 143, 151, 156, 166, 173, 183, 189, 197, 205, 209, 213, 218, 223, 227, 236, 250, 253, 260, 274, 280]. Biological [495, 845, 1006, 825, 819, 730, 370]. Biology [101, 389, 970, 496, 925, 341, 299, 605, 351, 697, 931, 163, 731]. Biophysics [816]. Birth [482]. Bit [762]. Black [759, 946, 689, 406]. Blades [1016]. Blame [8]. Blessings [983]. Blind [716]. Blindness [658]. Blue [104]. Bodies [771, 905]. Body [999, 379, 991, 847]. Bohm [538]. Bohmian [206, 826, 980]. Bohr [808]. Bolzano [700]. Book [981, 1018, 696]. Boomerangs [1016]. Borders [892]. Borges [739]. Bos [470]. Bose [957]. Boson [780]. Bostrom [217]. Bottom [440]. Bottom-Up [440]. Boundaries [340]. Box [826]. Brain [750, 954, 953]. Bram [470]. Branching [352]. Brazil [983]. Breaking [959, 62]. Breakings [396]. Brecht [452]. Bricks [544]. Bricmont [195]. Bridge [436]. Bridging [334]. Broader [516]. Broglie [436]. Bruce [105]. Bruno [506]. Brussels [805]. Building [991, 190, 527, 542, 544, 647]. Bunge [687]. Burgessian [414].

Burggraeve [581]. Butterflies [435].

C [990]. Calculus [651, 421, 965]. Call [72, 99]. Called [948]. Calvin [829]. Can [978, 1001, 344, 513, 760, 738, 623, 141, 600, 900, 394]. Canadian [812]. Candidate [879]. Can’t [533]. Canterbury [739]. Capacity [781]. Capital [871]. Care [629, 447]. Career [656]. Carnes [320]. Cart [690]. Cartesianism [429]. Case [123, 185, 962, 298, 124, 661, 564, 840, 111, 315, 966, 831, 401, 635, 126, 125, 128]. Case-Study [126]. Cassirer [822]. Casting [221]. Catastrophe [444]. Catastrophic [457]. Categorizations [831]. Caterpillars [435]. Cauchy [680, 421, 441]. Causal [344, 212, 907, 85, 222, 299, 975, 283, 259]. Causality [367, 963, 33]. Causally [978]. Causally-Explanatorily [978]. Causation [247, 975, 374, 865, 847]. Cause [212]. Causes [220, 662, 877]. Celebrated [645]. Century [424, 653, 838, 327, 838, 215]. Challenge [131, 931, 118]. Challenges [366, 577]. Chance [207]. Change [653, 775, 62, 623, 443, 503, 562, 527, 542, 647, 585]. Changes [737, 758]. Changing [192, 563, 697]. Chaos [338, 195]. Characteristic [505]. Characteristics [813]. Characterization [85]. Charge [669]. Charles [124]. Chemical [513, 512, 710, 754, 32]. Chemistry [984, 86, 366, 292, 31]. Chief [750]. Children [825, 233, 191]. Chimeras [458]. Chomskyan [490]. Choose [801, 304]. Chose [954]. Chutes [983]. Circle [461, 563]. Citizenship [573, 591, 590]. Civic [586, 592, 589, 587]. Civicness [588]. Civilization [609]. Civilizations [412]. Claes [588, 587]. Claims [642]. Classic [758]. Classical [442, 771, 658, 932, 961, 788, 599, 754]. Classicality [359, 719]. Classification [342, 795]. Climate [62, 276]. Closer [846]. Closure [393]. Cloud [669]. Cluster [801]. Clustering [795, 794]. Co [690]. Co-evolution [690]. Coal [484]. Cobot [1008]. Cobweb [339]. Cochrane [962]. Code [881]. Codex [1014]. Coefficient [753]. Coffee [978]. Cognition [682, 780, 37, 170, 339, 546, 994, 164, 147, 38, 10, 20, 36, 58, 547]. Cognitive [509, 196, 140, 170, 97, 857, 169, 736, 781, 779, 508, 163, 850, 843, 615]. Coherence [373]. Coherent [701]. Coherentism [298]. Cold [320]. Collaborative [306, 1008]. Collapse [679, 959, 515]. Collapsing [946]. Collective [913]. Collectively [626]. Collectives [935]. Collectivity [703]. Collision [771]. Collisions [504]. Colonization [1015]. Combination [590]. Combine [729]. Come [230]. Coming [543]. Comment [777, 588, 848, 818]. Commentary [13, 921]. Comments [610, 46, 25, 82, 696, 405, 598, 713, 77, 14]. Commercialization [561]. Common [820, 808, 879]. Communication [186, 766, 1007, 188, 571]. Communicative [111, 555]. Comparative [1021, 994, 879]. Comparing [169]. Comparison [908]. Compatible [177]. Competing [392]. Complementarism [298]. Complementarity [650, 764]. Completed [208]. Completée [208]. Completing [182]. Completion [974]. Complex [310, 403, 498, 659, 817, 566, 568, 501]. Complexity [494, 210, 211, 149, 497, 364, 541, 105, 730, 493, 754, 548, 1002]. Components [726]. Composed [282]. Compositional [864]. Comprehensible [766]. Comprehension [185]. Comprehensive [988, 958]. Compromises [923]. Computation [238, 791, 659, 1010, 1005]. Computational [705, 7, 50, 746, 571, 501, 370, 191, 81, 118]. Computer [934, 671, 927, 52, 670, 126, 481]. Computer-Supported [126]. Computing [307, 799]. Conceivable [96]. Concept [98, 842, 836, 323, 117, 195, 697, 317, 482, 144]. Conception [255, 343, 993]. Conceptions [240]. Concepts [957, 424, 221, 604, 735, 474, 872]. Conceptual [347, 137, 761, 596, 540, 341, 933, 245, 847, 527, 542, 544, 602, 647, 652, 420, 645, 879]. Conceptuality [743]. Concerning [412]. Concrete [843]. Concretisation [245]. Condensed [288]. Condensed-matter [288]. Condition [929]. Conditions [974, 408, 947]. Conductors [837]. Conferences [558]. Confidence [913, 22]. Configuring [319]. Conflicts [610, 126]. Conform [329]. confrontational [621]. Confusion [291]. Congo [909]. Congress [108]. Connecting [936, 960, 813, 246]. Connection [847]. Connes [458]. Conscious [376, 748]. Consciousness [912, 967, 112, 692, 169, 750, 153, 81]. Consensus [557, 558]. Consequences [605]. Conservation [358, 838, 834]. Considerations [867, 604, 646]. Considering [908]. Consistent [766]. Constant [584, 785]. Constituent [93]. Constitution [37, 997, 549, 431]. Constraints [221, 498]. Construct [765, 212]. Construction [196, 170, 338, 152, 245, 155, 165, 694]. Constructive [343, 251]. Constructivism [523, 164, 168, 163, 162]. Constructivist [777, 353, 159, 171, 721, 165, 160]. Contact [908]. Contemporary [883, 322, 488, 772, 84, 414, 840, 195, 727]. Content [935]. Contents [129, 158, 175, 198, 219, 261, 281]. Context [695, 185, 970, 184, 497, 186, 64, 738, 994, 528, 188, 530, 190, 850, 331, 191]. Context-Based [191]. Context-dependence [188]. Context-Dependency [497]. Context-Sensitiveness [970]. Contextual [842, 664]. Contextuality [747, 187, 744]. Contextualize [679]. Contextures [936]. Continental [511]. Contingency [226]. Continuity [266, 120]. Continuous [251]. Contribution [876, 759]. Contributions [918, 68]. Contributor [411]. Control [660]. Control-Theory [660]. Controls [239]. Controversies [610]. Conundrum [273]. conventionalism [18]. Conversation [171]. Converse [234]. Conveyor [124]. Conveyors [1023]. Cooperate [600]. Copenhagen [719]. Copernican [182]. Coping [633]. Corpuses [717]. Correcting [595]. Correction [803, 988, 914, 1010, 860]. Correctness [971]. Correlations [643]. Correspondences [959]. Cosmic [480, 609, 360]. Cosmological [785, 689, 409, 410]. Cosmology [915, 408, 284, 784, 655, 831]. Cosmos [929, 948]. Cosmotechnical [869, 938]. Coulomb [752]. Counterfactuals [285, 85, 222]. Coupled [495]. Course [291]. Courts [557]. Covariance [140]. Covariance/Invariance [140]. Covid [1007]. Covid-19 [1007]. Cracow [28]. Crane [1017]. Creating [16]. Creation [409]. Creationism [362]. Creative [138, 123, 124, 139]. Creativity [107, 738, 115, 753, 119]. Crisis [880]. Criteria [558]. Criterion [782]. Critical [989, 987, 66, 630, 27]. Criticism [968, 444]. Critique [989, 414, 710, 195, 411, 564]. Critiques [761]. Cross [192]. Cross-Cultural [192]. Crowd [220]. Crushing [1016]. Cubist [650]. Cultural [192, 364, 767, 417, 418, 605, 905]. Cultural-Biology [605]. Culture [148, 822]. Cultures [633]. Current [656]. Curves [1022, 915]. Cybernetica [154]. Cybernetics [943]. Cyborgs [435].

D. [77]. Dancing [750]. Dark [807]. Darwinian [101, 267, 141]. Darwinism [359, 360]. Dasein [929]. Data [643, 483, 457, 792, 372]. David [685, 3, 696]. Davidson [761]. Dealing [452]. Death [677, 564, 482, 363]. Debate [808, 952, 940, 754]. Debunking [438]. Decision [805, 962, 615, 956, 5]. Decision-Making [805, 962]. Decision-Oriented [615]. Decohere [679]. Decolonizing [812]. Deconstructing [952]. Deduction [480]. Deductive [119]. Deep [609, 216, 104]. Defamiliarizing [949]. Defeasible [375]. Defense [1011, 733, 258]. Defies [81]. Definition [973, 83, 363, 395]. Definitions [392, 394, 391, 153]. Deflection [910]. Degree [659, 784]. Déjà [676]. Déjà-Vu [676]. Dekker [589]. Delta [441]. Deluge [643]. Delvaux [674]. Demarcation [782]. Demise [730]. Democratic [559]. Democratizing [560]. Demodulate [801]. Demolishing [782]. Demonstrate [738]. Density [669]. dependence [188]. Dependency [497]. Depends [691]. Descartes [991, 475]. Descent [776, 526]. Descriptional [179]. Descriptions [775, 485]. Design [362]. Desperate [923]. Details [776]. Detection [794]. Determinism [997, 48]. Detour [944]. Developing [371]. Development [232, 885, 283, 368, 570, 356, 144, 162]. Developments [202, 10]. Develops [168]. Dewey [319]. Diagrammatic [311, 232, 332]. Diagrams [732, 242, 337]. Dialectical [925, 646, 898]. Dialectics [921, 635, 928, 944, 873]. Dialetheism [490]. Dialogue [553]. Diamonds [484, 486]. Dice [536]. Dichotomies [932]. Dichotomy [844]. Dictatorship [825]. Did [695, 670]. Difference [560, 351, 318]. Different [583]. Differential [299, 722]. Differentiating [584]. Difficult [321]. Digestive [345, 350]. Digital [945, 998, 1004, 1003, 625, 887, 1000, 1009, 951, 872, 626]. Dilemma [612, 695]. Dimension [592, 232, 574]. Dimensional [795, 282]. dinosaurs [17]. Diplomacy [320]. Dirac [981, 441]. Direct [212]. Directed [349]. Direction [689]. Directionality [939]. Directions [682, 683]. Dis [625]. Discarding [941]. Discipline [890]. Discontinuous [454]. Discourse [480]. Discourses [944]. discovered [722]. discoverer [16]. Discoveries [141]. Discovering [705]. Discovery [235, 550, 137, 768, 738, 115, 994, 231, 127, 120, 568, 121, 481, 531, 14, 13, 450, 15, 12]. discussants [26]. Discussing [932]. Discussion [992, 3, 51, 697]. Disempowerment [913]. Disentangling [662]. Disorders [906]. Displacing [384]. Dispositionalist [255]. Dispositions [419]. Disproportionate [904]. Disputes [845]. Dissent [560]. Dissolute [932]. Distinction [710]. Distinguishing [712]. Distributed [308]. Disunity [853]. Diverse [767]. Diversity [751, 854, 142].

Divide [201]. DNA [531]. Do [329, 95, 783, 230, 749, 536, 394]. Document [795]. Documentary [678, 674, 673]. Documents [717]. Does [538, 365, 930, 715, 465, 622, 474, 262, 947]. Dogma [257]. Doing [451]. Domain [171, 712]. Domains [137]. Dominelli [592, 591]. Donald [761]. Don’t [62]. Double [716]. Double-Blind [716]. Doubling [910]. Doubt [642]. Down [440]. DR [909]. Drama [327]. Drawing [1022]. Drift [661, 511, 605]. Dual [421]. Dualistic [787]. Duality [237]. Duhem [200, 952]. During [424, 1007, 921, 873]. Dwelling [388]. Dynamic [138, 839]. Dynamical [971, 498]. Dynamically [495]. Dynamics [884, 169, 959, 247, 417, 418, 401, 514, 585]. Dyscalculia [844]. Dyscalculia/Dyslexia [844]. Dyslexia [844]. Dystopia [902].

Ear [912]. Early [656, 233, 503, 563, 1017]. Earth [876, 361]. Ecological [726, 681]. Ecology [754]. Economic [200, 325, 202, 321, 822]. Economics [648]. Economist [201]. Economist/Economy [201]. Economy [252, 201]. Econophysics [849]. Ecosystem [881]. Edition [933]. Editorial [87, 107, 100, 130, 184, 39]. Editors [866]. Education [833, 137, 908, 192, 1019]. Effect [538, 212, 983, 456]. Effective [527, 542, 647]. Effects [217, 915, 399, 222, 657, 662, 979]. Efficiency [801, 828]. Efficient [792]. Einstein [957, 140, 808, 910, 221, 525, 111]. Elastic [745]. Elche [1020]. Eleatic [995, 443]. Electric [314]. Electricity [837]. Electrodynamics [752, 788]. Electron [669]. Electronic [243]. Elementary [985]. Elements [985, 513, 512, 710, 1014, 90]. Elgin [818]. Eliminative [155]. Emanation [477]. Embeddable [651]. Embedding [661, 516]. Embodied [1001, 913, 315]. Embodiment [428, 378, 1007, 379, 430]. Embracing [470, 221]. Emergence [240, 955, 313, 659, 712, 283, 730, 611, 554, 145]. Emergent [210, 211, 752, 823, 720]. Emmer [674]. Empirical [897, 987, 971, 187, 911, 239, 460, 264, 867, 889, 938, 616, 121, 928, 821, 118]. Empiricism [207, 249, 7]. Empiricist [517]. Employed [123]. Empowerment [625]. Emptines [288]. Encoded [943]. Energy [801, 841, 836, 807, 358, 838, 525, 834]. Engaging [560]. Engineering [311, 407, 1015, 731]. Enjoyable [513]. Enough [593, 395]. Enriques [995]. Ens [273]. Entailment [942]. Entangle [679]. Entangled [780, 665]. Entanglement [682, 742, 803, 802, 704, 717, 750, 959, 806, 753]. Entanglements [916]. Entangling [991]. Entities [347, 701, 269, 437, 32, 31]. Entitled [262]. Entity [930]. Entropocene [877]. Entropy [220, 392, 759]. Environmental [399, 64]. Environments [1006, 1004, 711]. Enzyme [345, 350]. Ephemeral [402]. Epistemic [978, 962, 488, 747, 751, 284, 507, 919, 731, 718]. Epistemological [989, 707, 325, 768, 829, 505]. Epistemology [812, 761, 552, 556, 813, 546, 810, 658, 384, 547]. Equality [148]. Equations [711, 722]. Equivalence [525, 820]. Ergodic [275]. Erik [588, 587]. Erkki [676]. Ernst [822]. Errata [113]. Erratic [312]. Erratum [647, 547]. Error [349]. Error-Statistics [349]. Errors [60, 843]. Eschatology [220]. Especially [251]. Essential [40, 42]. Estimation [425]. Eternalism [611]. Ether [482]. Ethical [881, 579, 325, 386, 308]. Ethics [633, 629, 631, 61, 880]. Euclid [820]. Euler [722]. European [695]. Evaluating [636, 633]. Event [835, 869, 578]. eventruth [118]. Events [84, 691]. Every [776, 51]. Everyday [190, 585, 582]. Everywhere [812, 474]. Evidence [716, 817, 460]. Evidence-Based [716]. Evolution [500, 365, 342, 170, 364, 541, 884, 358, 169, 609, 49, 283, 730, 417, 418, 147, 368, 356, 1023, 690]. Evolutionary [761, 154, 341, 605, 931, 153]. Evolved [837]. evolving [141]. Ex [963, 340]. Exaggerated [730]. Examining [641]. Example [360, 710]. Examples [93, 950]. Exception [870]. Exceptional [924]. Exchange [797]. Exemplars [305]. Exist [474, 947]. Existence [269, 773, 993, 31]. Existential [719]. Expanding [545, 433]. Expectations [749]. Experience [840, 439, 153]. Experiencing [580]. Experiment [841, 835, 859, 826]. Experimental [1006, 215, 847]. experimentalism [31]. Experiments [724, 836, 302, 540, 837, 645]. Expert [557]. Explain [663, 760, 688]. Explained [344]. Explaining [395]. Explains [915]. Explanation [359, 857, 237, 986, 855, 975, 854, 856, 258, 954, 259, 853, 27]. Explanations [496, 931, 655, 214]. Explanatorily [978]. Explanatory [347, 690, 858, 139, 247, 854, 856]. Explication [642]. Exploitative [590]. Exploration [702]. Explorative [908]. Exploring [781, 786, 824, 767, 940]. Exposition [768]. expression [102]. Extended [389]. Extensions [666]. External [246]. Externalism [885]. Externalized [887]. Externalizing [308]. extinct [17]. Extra [831]. Extraction [1021, 824].

F. [77]. Fabric [97, 396]. Fabricated [727]. Fabulations [893]. Face [524]. Facing [967]. Fact [535]. Fact-Arithmetic [535]. Facts [296, 694]. Faith [827]. Fall [983]. Fallacies [741, 713]. Falling [601]. Falsifiable [594]. Falsification [594]. Falsificationism [271, 155]. Falsified [271]. Family [305]. Fantasy [951]. Far [412]. Farming [64]. Faster [830]. Fe [849]. Feature [795, 792, 794]. Features [834]. Feedback [599]. Fermat [695]. Fermi [790]. Fetal [906]. Fiction [738]. Field [211, 297]. Fields [176]. Fighting [1016]. Film [674]. Final [589, 585]. Finality [705]. Finance [919]. Finch [974]. Finding [532, 808]. Findings [1012]. Fine [370, 410]. Fine-Tuned [370]. Fine-Tuning [410]. First [651, 152, 813, 838, 1014, 719, 57, 365]. First-Person [813]. Fish [64]. Fish-Farming [64]. Fitting [575]. Fix [62]. Fixed [520]. Fleck [935]. Flow [396]. Fluctuations [194]. Fluid [600]. Flunks [640]. Focus [970]. Focused [654]. Follow [365]. Food [665]. Forces [752]. Forensic [656]. Foresee [670]. Forgotten [512]. Form [776]. Formal [100, 329, 510, 865, 375, 265, 102, 214]. Formalism [206]. Formalization [835, 785]. Formation [140, 824]. Formative [909]. Forms [259]. Formula [525]. Formulating [815]. Formulation [664, 601]. Forum [65]. Fossil [365]. Foucault [630]. Foundation [10, 723, 492, 2]. Foundational [275, 680, 776].

Foundationalism [298]. Foundations [87, 101, 38, 20, 36, 37, 58, 610, 67, 782, 960, 154, 815, 824, 791, 609, 225, 148, 767, 839, 832, 809, 251, 3, 1]. Founder [995]. Four [944]. Fraassen [269]. Fractal [357, 755]. Fragile [580]. FRAME [884]. Frames [491]. Framework [347, 344, 342, 216, 375]. Frameworks [426, 839, 933]. Franca [751]. France [327]. Francia [54, 25]. Free [789, 848, 601, 964]. Freedman [3]. Freedom [659, 607]. French [208]. Friends [304, 404]. Frisch [988, 958]. Fulfillment [575]. Function [441, 795, 346, 963, 731]. Functions [505, 817, 251]. Fundamental [424, 408, 709, 975, 701]. Fundamentality [782]. Further [442, 463]. Future [196, 648, 609, 691, 413, 412].

G. [545]. G3 [965]. GA [801]. GA-ANFIS [801]. Galactic [831]. Galaxies [915]. Galilean [415]. Galileo [601]. Games [913, 214]. Gap [334]. Gas [780]. Gastrointestinal [817]. Gate [864]. Gauge [663]. Gave [421]. Gene [374]. Genealogy [849]. General [98, 45, 584, 884, 918, 625, 770, 366, 604, 90]. Generalised [278]. generalization [332]. Generalizations [694]. Generative [741, 990]. Genesis [600]. Genetic [792]. Genetics [315]. Geneva [933]. Geology [661]. Geometric [736, 953]. Geometrization [707]. Geometry [901, 707, 995, 506, 283, 850]. Get [782, 232, 465]. Ghost [452]. GHZ [285]. Gift [580]. Gigantism [875]. Gionta [767]. Giuliano [54]. Global [53, 657]. God [47, 317, 536]. Gödel [221, 360, 52]. Gold [716, 818]. Good [939, 574]. Gordon [171]. Gould [348]. Grace [580]. Grand [79]. Gravitation [279]. Gravitational [686, 687]. Gravity [369, 755]. Great [800]. Greek [901]. Greenhouse [657]. Gregory [668]. Grey [781, 786, 824]. Grossone [713]. Grossone-Based [713]. Ground [808, 563]. Grounded [648]. Groundwater [664]. Group [656]. Groups [663]. Growth [364]. Grzegorz [75]. Guest [866]. Guises [263]. Guliano [25]. Gut [350].

h [580, 578]. Habituation [949]. Hadrons [720]. Haecceity [597]. Hameroff [112]. Happens [976]. Hard [967]. Harm [394]. Harmony [1022]. Hate [757]. Head [801]. Health [903, 904]. Health-Related [903, 904]. Heart [704]. Heat [837]. Hegel [944]. Hegelian [898, 928]. Heidegger [619, 883, 620]. Heideggerian [875]. Heisenberg [125]. Henri [674]. Here [812]. Here/There/Everywhere [812]. Hermeneutics [204]. Heterogeneous [797]. Heteronomy [625]. Heuristic [140, 664, 346, 600, 125]. Heuristics [725]. Hidden [641]. Hierarchical [363]. High [908, 795]. Hilbert [809]. Historical [707, 776, 609, 133, 1018, 283, 767, 839, 120, 614, 1023, 24]. Historiographical [885]. Historiography [414, 741]. History [1016, 438, 654, 660, 421, 1012, 345, 350, 914, 861, 252, 1019, 832, 598, 1017, 562, 482, 10]. Hole [759, 406]. Holes [946, 689]. Holism [103]. Holistic [876, 864, 928]. Holons [799]. Honohan [589]. Horse [690]. Hub [921]. Huhtamo [676]. Hull [348]. Human [780, 805, 957, 829, 787, 781, 449, 559, 810, 148, 188, 147, 904, 171, 843, 536, 104]. Human-Machine [810]. Humanities [556]. Humanity [413]. Humanoid [881]. Humans [541, 435]. Hybrid [629, 792, 796]. Hyper [535]. Hyper-Slingshot [535]. Hyperreals [458]. Hyperuniverse [969]. Hypotheses [110, 247, 230]. Hypothesis [690, 669, 714, 142]. Hypothetical [737].

ICT [633, 625, 635]. ICTs [633]. idea [31]. Ideas [520, 195]. Identification [754]. Identifying [682, 560, 544]. Identity [825, 930, 715, 709, 562]. Ideologies [519]. IDTC [832]. if [463, 597, 62]. Ignorance [978, 724]. Ignorance-Preserving [724]. Ihde [377, 872]. II [726, 646, 683, 211, 345, 60, 418]. Illusion [402]. Illusions [868]. Ilya [343, 195]. Im [379]. Image [704, 292]. Imagery [760]. Imagination [945, 307, 887, 951, 872]. Imaging [439]. Imitation [828]. Immanent [916]. Immobility [699]. Impact [904]. Implications [633, 369, 325, 295, 499, 59, 666, 451, 994]. Implies [719]. Import [498]. Importance [231, 940]. Imposed [360]. Impossibility [239, 534]. Imprecise [986]. Impulse [988, 958]. Impure [522, 726, 681]. In-Between [381, 463, 453, 388, 378, 385, 377]. Incisive [790]. Incommensurability [642, 120]. Incompatibility [442]. Incomplete [491]. Incompleteness [299, 360, 52]. Inconsistencies [93, 92]. Inconsistency [97, 415]. Increasing [801]. Independence [713]. Independent [171]. Indeterminism [48]. Indispensability [207]. Indistinguishable [957]. Individual [749]. Individuality [355]. Individuation [348]. Indivisibles [685, 696, 684]. Induced [855]. Induction [891, 491]. Inductivism [155]. Industrial [875]. Industry [192]. Inequality [148]. Inessential [44]. Infants [191]. Inference [986, 511, 954, 128]. Inferences [243]. Inferential [725]. Inferring [222]. Infinite [776, 242, 43, 655, 700]. Infinitesimal [242]. Infinitesimals [685, 695, 680, 696]. Infinities [708]. Infinity [769, 713, 816, 604, 646, 763]. Inflation [360]. Influence [977, 749]. Informal [100, 329]. Information [765, 852, 943, 962, 967, 392, 824, 517, 659, 224, 360, 391, 890, 417, 418, 865, 243]. Information-based [360]. Informational [546, 352, 547]. Informons [799]. Infosphere [621]. Inheritance [671]. Initial [408]. Inquiries [601]. Inquiry [813, 818, 867]. Insight [830, 527, 542, 647]. Inspired [806]. Instance [491]. Instead [807, 39]. Institute [849]. Institutions [905]. Instructions [157, 174]. Instruments [934, 703]. Integrated [967, 865]. Integrating [979]. Integration [390]. Integrity [61]. Intellect [625]. Intellectual [457]. Intelligence [828, 51, 165]. Intelligencer [640]. Intelligent [362, 412]. Intentional [1011]. Intentionality [990, 265]. Inter [712]. Inter-domain [712]. Interaction [192]. Interactional [212]. Interactions [499, 663, 396]. Interdependence [810]. Interdisciplinarity [87]. Interdisciplinary [729]. Interest [793]. Interfaces [748]. Interferometry [645]. Intermedial [387]. Internal [246]. Internalism [885]. International [803, 35, 742, 802].

Interpersonal [966]. Interpretation [743, 306, 901, 278, 747, 771, 269, 719, 551, 980, 1010, 1005]. Interpretations [836, 963]. Interpreter [995]. Interpretive [886]. Intertheoretical [823]. Interval [874, 916]. Interventions [737, 906, 444]. Interview [55, 54, 105, 112]. Intra [712]. Intra-domain [712]. Intrinsicality [865]. Introducing [841]. Introduction [199, 1012, 573, 369, 998, 556, 960, 65, 59, 866, 832, 804, 262, 356, 89, 487, 73]. Introductory [122]. Intrusion [794]. Intuition [345, 350]. Invaders [446]. Invariance [140]. Invariances [169]. Inventing [314]. Invention [748]. Inventioni [767]. Inverse [728]. Investigating [544, 853]. Investigation [322, 999, 164]. Investigations [563]. iron [94]. iron-reducing [94]. Irrational [301]. Irreversibility [774]. Isis [524]. isn’t [996]. Issue [742, 802, 1012, 903, 998, 791, 832, 262, 356, 393, 960, 803]. Issues [72, 99, 295, 919, 3, 2]. Italian [840].

J [545]. Jacobs [675, 676]. Jagers [393]. Jean [195]. Jesuits [685, 684]. JFK [677]. Jorge [739]. Joule [841, 835, 838, 840, 837, 832]. Journey [443]. Just [473]. Justification [66, 507, 834].

Keep [695]. Ken [675, 676]. Kepler [234]. Ketland [846]. Key [830, 797]. Kingdom [819]. Know [606, 62]. Knowing [772, 171]. Knowledge [998, 968, 232, 824, 560, 740, 308, 534, 827, 563, 79, 514, 944, 16, 34, 57]. Knowledges [429]. Kolmogorovian [667]. Krishna [793]. Kuhn [885, 526].

Lacan [944]. Lack [642]. Lake [905]. Landmark [914, 861]. Landscape [324]. Language [780, 38, 10, 20, 36, 37, 58, 957, 876, 186, 187, 225, 913, 191]. Languages [249]. Lanka [665]. Lasota [55]. Latent [798]. Later [734]. Latour [464]. Law [814, 601, 766, 838]. Laws [255, 365, 343, 477, 121, 8]. Layers [945, 282]. leading [118]. Leaps [283]. Learn [305, 745]. Learning [349, 141, 615]. Left [976]. Legal [766]. Leggenda [674]. Leibniz [479, 478, 505, 474]. Leibnizian [651]. Lena [592, 591]. Leonardo [740]. Lessons [991, 557]. Level [819]. Levels [549, 823]. Liar [109, 278]. Liar-paradox [109]. Liberal [718]. Libro [767]. Life [878, 1001, 270, 541, 394, 580, 691, 411, 190, 472, 371, 413, 714, 665, 577, 574, 363, 395, 585, 582]. Light [910, 871, 639, 215]. Limitations [822, 102]. Limits [63, 28, 533, 360]. Lindfelt [581]. Lines [526]. Lingua [751]. Linguistic [323, 415, 534]. Linguistics [741, 127]. Linking [342]. Links [164]. Lipton [954]. Lissajous [1022]. Listeners [304]. Litigation [557]. Little [142]. Living [365, 633, 342, 634, 96, 630, 216, 899, 621, 618, 616, 617, 635]. Living-beings [365]. Local [53, 788]. locality [285]. Localization [94]. Localizations [902]. Location [549]. Lockean [488]. Locker [658]. Logic [651, 970, 488, 489, 360, 178, 490, 316, 231, 283, 147, 961, 746, 735, 898, 544, 214]. Logical [355, 178, 827, 891, 713, 18]. Logics [235, 864, 965, 119]. Logos [992]. Lombardo [840]. Long [330, 169]. Long-Term [169]. Look [146, 846].

Loop [301]. Loophole [789, 848]. Loophole-Free [789, 848]. Lord [320]. Lorentz [863]. Loss [659]. Lost [774]. Love [576, 757, 628]. Loving [626]. Luciano [674]. Lucio [840]. Luddite [444]. Ludwik [935]. Luhmannian [652]. Luis [739]. Luk [666]. Lyons [733].

Mach [221]. Machine [14, 810, 1019, 127, 1014, 12, 437, 13]. Machines [705]. Macro [186, 823]. Mad [677]. made [141]. Madrid [1014]. Magic [996]. Magritte [729]. Main [332]. Major [366]. Make [513, 141]. Makes [883]. Making [805, 962, 636, 177, 216, 243, 956]. Man [881, 432, 51, 740]. Managed [714]. Managed-Metabolism [714]. Mandatory [688]. Maneuvering [916]. Manipulation [128]. Manipulative [229]. Many [553, 588, 920, 961]. Many-Measurements [553]. Many-Valued [961]. Many-Worlds [553]. Margins [968]. Mario [687]. Marriage [758]. Mars [439]. Mass [525]. Material [543]. Materialism [405]. Materialistic [403]. Materials [126]. Mathematical [707, 424, 550, 335, 958, 640, 353, 775, 49, 398, 337, 931, 604, 735, 773, 331, 263, 88, 698, 988]. Mathematician [249]. Mathematics [336, 207, 654, 41, 440, 613, 732, 648, 807, 267, 40, 47, 882, 42, 333, 458, 414, 779, 48, 702, 242, 346, 264, 334, 933, 268, 570, 646, 502, 492, 251, 638, 262]. Mathematization [702]. Matter [985, 807, 518, 326, 834, 269, 288]. Maximality [969]. Maxwellian [555]. May [730, 536]. MDL [139]. Mean [783, 262]. Meaning [323, 584, 580, 216, 411, 371, 286, 583, 577, 582]. Meaningful [588, 590, 532]. Meaningfulness [579, 586, 592, 589, 587, 578, 583, 574, 573]. Means [413, 489]. Measure [187, 793]. Measurement [477, 126, 698, 744]. Measurements [553, 710, 745]. Measuring [399, 465]. Mecha [1016]. Mechanical [93, 109, 1015, 1020, 963]. Mechanics [978, 992, 762, 206, 181, 285, 419, 180, 176, 777, 238, 723, 984, 1022, 516, 729, 813, 132, 838, 785, 982, 639, 1019, 886, 719, 551, 721, 600, 410, 826, 980, 6, 981]. Mechanism [1012, 137, 1021]. Mechanisms [495, 922, 1018, 141, 256, 662, 422]. Mechanistic [496, 549]. Media [895, 445, 434]. Mediating [302, 439, 703]. Mediation [629, 632, 840, 380, 940, 433]. Medicine [716]. Medieval [1018, 503]. Meduso [369]. Meeting [679, 577]. Meets [970, 772, 556]. Membranes [745]. memetics [500]. Memorization [513]. Memory [926, 506]. Men [677]. Mending [97]. mensura} [340]. Mental [724, 152, 760]. Merely [1011]. Meta [838]. Meta-law [838]. Metabolism [714]. Metallic [837]. Metametaphysical [992]. Metamodel [521]. Metaphor [124, 540]. Metaphors [724, 935]. Metaphysical [972, 534]. Metaphysics [976, 475]. Method [685, 776, 399, 299, 568, 684]. Methodological [362, 919, 115, 615]. Methodology [206, 200, 67, 925, 154, 202, 231, 713, 375]. Methods [303, 716]. Mexico [1013]. Michael [772, 768]. Michel [467]. Micro [186]. Microbiomes [697]. Microscopes [461]. Microscopic [402]. Microstructures [725]. Microtubules [112]. Mid [424]. Mid-nineteenth [424]. Midst [384]. Might [896]. Mikael [581]. Milieu [953]. Milk [978]. Mill [1013]. Mimesis [130, 133]. Mind [196, 945, 47, 991, 518, 847]. Mind-Body [847]. Minds [268]. Mines [1021, 1013]. Minimal [500]. Mining [457, 793]. Miracle [972]. Misconceptions [438, 362, 598]. Mismatch [807]. Misrepresentation [528, 530]. Missing [834]. Modal [489, 827]. Model [683, 278, 846, 212, 988, 958, 787, 324, 855, 666, 229, 990, 521, 245, 551, 171, 571, 191, 437]. Model-Based [229]. Model-Induced [855]. Model-Theoretic [846]. Modeling [116, 805, 711, 354, 276, 569]. Modelling [310, 123, 304, 149, 489, 302, 886, 118]. Models [724, 496, 311, 812, 510, 248, 971, 302, 313, 353, 483, 400, 726, 351, 501, 375, 681, 401, 599, 92, 155, 514, 243, 23, 30]. Modern [991, 935, 927, 321, 503, 366, 481]. Molecular [943, 970, 496, 984, 299, 605, 422, 315]. Monde [674]. Monotonic [492]. Moral [307, 319]. Morality [308]. Morals [268]. morendo} [888]. Morphodynamical [247]. Mother [457]. Mothering [904]. Motion [699, 601]. Moved [578]. Movements [450]. Moves [361]. Multi [917, 746]. Multi-Quantificational [917]. Multiagent [310]. Multidimensional [870]. Multilevel [138]. Multiple [671, 600]. Multiplicity [857]. Multiscale [390]. Mum [695]. Musical [304, 935]. Must [403]. Mutually [212]. my [26]. Mystery [1020]. Mysticism [758]. Myth [432, 283].

Names [512]. Nanochemistry [823]. Nanomaterials [823]. Nanoscale [875]. Narrow [983]. Nastrój [935]. Natiello [777]. Natural [101, 249, 432, 772, 541, 186, 475, 774, 708, 511, 605, 502, 607]. Natural-Language [186]. Naturalism [362, 67, 718, 931, 263]. Nature [682, 692, 408, 84, 386, 559, 525, 457, 343, 383, 832, 834, 477, 954, 215, 843, 8]. Navigating [874]. Near [905]. Necessarily [897, 939]. Necessity [942]. Need [881, 949, 95]. Needs [919]. Neglected [932]. Negotiating [387]. Neither [474, 22, 203]. NEMO [797, 797]. Neo [476]. Neo-Platonic [476]. nessing [580, 578]. Network [930]. Networks [801, 494, 495, 797, 796, 141]. neural [141]. Neuroeducation [564]. Neuropharmacy [128]. Neuroscience [178]. Neuroscientific [999]. Neutron [645]. News [939]. Newton [478, 777, 835, 475, 480, 933, 476, 721, 474]. Newtonian [807, 715, 226, 477]. Next [162]. Nick [217]. Nietzschean [727]. Nihilism [617]. Nihilo [963]. nineteenth [424, 838]. No [972, 963, 770]. No-Thing [963]. Nodes [801]. Nominalistic [414]. Nomograms [1019]. Non [494, 285, 718, 807, 342, 761, 882, 667, 813, 355, 324, 621, 713, 963, 615, 754, 492, 145, 437, 811]. Non- [355]. Non-abstract [882]. Non-adaptationist [761]. Non-Aggregativity [145]. Non-classical [754]. Non-cognitive [615]. Non-confrontational [621]. Non-epistemic [718]. Non-Living [342]. Non-locality [285]. Non-Monotonic [492]. Non-Newtonian [807]. Non-ordinary [437]. Non-Spatial [811]. Non-standard [713]. Non-trivial [813]. Non-Uniqueness [494]. Noncontextual [516]. Nonfoundational [637]. Nonlinear [150, 399, 339]. Nonlocality [131]. Nor [474, 22]. Normal [61]. Norms [329, 749]. Not-is [328]. Note [122, 39, 29, 21]. Notes [56, 71, 106, 114, 143, 151, 156, 166, 173, 183, 189, 197, 205, 209, 213, 218, 223, 227, 236, 250, 253, 260, 274, 280, 193, 637, 9, 19, 135]. Notions [820, 51, 735]. Novel [141]. Novelty [738, 730]. Nuclear [193]. Number [776]. Numbers [416, 268]. Nurturing [469]. Nut [983]. NZT [635].

Object [670, 900]. Object-Oriented [670]. Objections [733]. Objective [517]. Objectivity [918, 561, 179, 427]. Objects [152, 1011, 458, 398, 831]. Obliterating [804]. Observability [361]. Observation [217]. observations [1000]. Observer [456]. Observers [690, 536]. Observing [161]. Occur [538]. Oceanography [124]. Offer [86]. Offered [153]. Offs [856]. Old [622]. Olympics [640]. Omelette [744]. Once [723]. One [835, 688, 465, 889]. Online [386, 1007, 966, 796]. Only [95, 394]. Ontic [955, 947]. Onto [812]. Onto-Epistemology [812]. Ontological [426, 955, 922, 984, 917, 202, 554, 763, 831, 293]. Ontologicum [739]. Ontologies [463]. Ontology [206, 997, 270, 1008, 471, 917, 546, 806, 769, 297, 937, 701, 459, 611, 554, 476, 328, 77, 547]. op [393]. Opacity [380]. Opening [377]. Operation [668]. Operational [181, 507, 740, 778]. Operationalism [901]. operationalization [821]. Operationism [6]. Operator [393]. Opposite [713]. Opposition [298]. Optical [242]. Oracle [532]. Ordeal [877]. Order [651, 641, 642, 979]. ordinary [437]. Organic [364]. Organism [697, 363]. Organisms [282]. Organization [367, 819, 754]. Organizational [210, 211, 212]. Organized [936]. Organizing [981, 959]. Oriental [758]. Oriented [615, 670]. Origin [367, 147, 395]. Original [696]. Origins [477, 714, 263]. Ornithologicum [739]. Orthogonal [810]. Orthomodular [974]. Oscillators [495]. Osculating [461]. Other [350, 600, 176]. Our [101, 629, 68, 373, 605]. Outcomes [698, 744]. Outline [706]. Outlining [716]. Overview [805].

Paddle [841]. paddy [94]. Painless [272]. Painting [674]. Paintings [729]. Pairing [797]. Pairing-Based [797]. Panel [73]. Paper [598, 13, 25]. Papers [108]. Paraconsistency [91]. Paraconsistent [490]. Paradigm [389, 786]. Paradigms [680, 120]. Paradox [278, 859, 109, 964]. Paradoxes [699, 344]. Paradoxical [993]. Parameters [408, 370]. Parmenides [995, 423]. Part [803, 802, 211, 345, 332, 417, 418, 167, 744]. Partial [572, 375]. Participating [411]. Particles [347, 985, 288, 402]. Particular [366]. Party [797]. Pask [171]. Past [244, 691]. Patching [97]. Patented [500]. Paternity [927]. Path [725]. Pathos [727]. Pathway [879]. Pattern [667, 795]. Patterns [793]. Paul [674]. Pauli [125]. Peirce [124, 117, 1010, 1005]. Penrose [606]. Perceiving [242]. Perception [704, 428, 376, 1006, 427, 1000]. Perceptual [1004, 439]. Performative [675]. Performing [451, 387]. Periodic [710]. Permanence [420]. Persistence [342]. Person [813]. Personal [825]. Persons [152]. Perspective [109, 895, 660, 1001, 192, 97, 353, 781, 656, 652, 703]. Perspectives [990, 899, 1009, 583, 165]. Persuasiveness [333]. Pessimistic [644]. Pessimists [644]. Phantasmagorical [452]. Phantom [952]. Pharmaceutical [192]. Pharmakon [625]. Phenomena [310, 212, 971, 480]. Phenomenological [1001, 270, 990, 1000]. Phenomenology [878, 1006, 999, 998, 471, 893, 1003, 451, 637, 1009, 387, 1002]. Phenomenon [225, 607]. Philip [470]. Philosophical [193, 362, 1006, 295, 786, 824, 702, 525, 710, 867, 317, 604, 646, 599, 277, 1]. Philosophically [648]. Philosophies [882]. Philosophy [895, 901, 883, 440, 217, 407, 757, 671, 556, 499, 960, 880, 751, 918, 42, 919, 475, 545, 977, 372, 334, 822, 203, 86, 950, 493, 870, 670, 616, 502, 292, 548, 727, 482, 125, 108]. Phonemic [240]. Physical [220, 170, 787, 396, 460, 778, 551, 898, 698]. Physicalism [939]. Physics [389, 74, 707, 193, 930, 807, 939, 392, 859, 532, 288, 936, 408, 169, 471, 42, 838, 391, 296, 977, 346, 208, 839, 933, 351, 804, 168, 215, 816, 370, 79, 482]. Physiologic [817]. Physique [208]. Piper [675]. Places [708]. Planck [296]. Planck-Scale [296]. Planetary [929]. Plate [661]. Plato [670]. Platonic [476]. Play [536]. Plea [468]. Please [62]. Plimpton [851]. Pluralism [917, 276]. Pluralistic [77]. Poetics [327, 444]. Poiêsis} [880]. Poincaré [18, 694]. Point [1016, 239, 134]. Points [510, 596, 401]. Polanyi [772, 768]. Policy [64, 544, 879]. Political [592, 325, 467, 587, 252, 470, 466, 562, 627]. Politics [381, 382, 463, 327, 447]. Politization [468]. Polycontextural [936]. Polyphonic [665]. Polyrhythmic [665]. Popper [67, 66, 68]. Popperian [65, 69]. Porphyrian [671]. Posets [974]. Positions [889, 599]. Positivist [473]. positivists [18]. Possibilities [379, 451]. Possibility [947]. Possible [347, 716, 369, 984, 535, 612, 164, 554, 849]. Possibly [759]. Post [893, 466, 473, 61, 997]. Post-Normal [61]. Post-Phenomenology [893]. Post-Political [466]. Post-Positivist [473]. Postgraduate [192]. Posthuman [945]. Posthumanism [631]. Postmodern [786, 324]. Postphenomenological [430, 990, 427]. Postphenomenologist [465]. Postphenomenology [377, 382, 911]. Postulates [442]. Potential [314]. Potentiality [397]. Power [913]. Powers [992]. Practical [290]. Practice [641, 277, 398, 384, 450]. practice-based [450]. Practices [886]. Practicing [873, 921]. Practise [105]. Pragmatic [103, 232, 710, 492]. pragmatics [450]. Pragmatism [204, 199, 290, 200, 932, 203, 201]. Praxis [925, 287]. Pre [839, 867]. Pre-Philosophical [867]. Pre-Service [839]. Prebiotic [422]. Precaution [63]. Precautionary [64, 59, 60, 62, 61]. Predicate [316]. Predict [688, 501]. Predominant [313]. Preexistent [744]. Preface [742, 803, 802, 903, 136, 228, 241, 300, 309, 167, 254]. Preferences [66]. Prejudices [782]. Preliminary [322, 999]. Presence [387, 448]. Present [539]. Preserving [724]. Presuppositions [641]. Prigogine [343, 195]. Primary [985]. Prime [776]. Primordial [946]. Principal [272]. Principia [154, 933]. Principle [194, 272, 369, 705, 64, 59, 60, 659, 62, 305, 142, 61, 139]. Principles [833, 981, 442, 969]. Prism [472]. Pro [464]. Pro-Latour [464]. Probability [181, 747, 996, 269, 6]. Probable [520]. Probably [896].

Problem [101, 613, 967, 836, 248, 815, 785, 702, 865, 870, 770, 599, 332, 31]. Problems [728, 117, 332, 792, 622, 1009]. Procedures [654]. Process [509, 354, 245, 15]. Processes [306, 922, 824, 774, 50, 485, 398, 5]. Produced [50]. Production [406, 944]. Professor [80]. profile [94]. Profiled [798, 793]. Program [716]. Programme [969]. Programming [299, 670]. Programs [312, 375]. Progress [830, 341, 614, 602, 118]. Project [154]. Projects [72, 636, 312]. Promises [448]. Proof [49, 266]. Proof-analysis [266]. Proofs [329, 330, 858, 533]. Propagation [988, 958]. Properties [53, 402, 744]. Property [597]. Proposal [270, 648, 829, 784, 701]. Proposals [99]. Prospects [532, 127, 263]. Protection [796]. Protist [819]. Provide [715]. Providing [551]. Proving [613]. Proximity [727]. Psy [564]. Psyche [150]. psychological [5, 821]. Psychologist [239]. Public [320]. Publisher [29, 21]. Puech [467]. Pulleys [1017]. Punches [1016]. Pure [759, 862, 294]. Purpose [101]. Purposes [406]. Putnam [146]. Putting [690]. Puzzle [540].

QBism [672]. QBist [937]. QED [344]. QM [344]. Qualitative [909, 128]. Quality [898]. Quantification [316]. Quantificational [917]. Quantique [208]. Quantities [700]. Quantity [898]. Quantum [93, 109, 130, 347, 683, 716, 743, 780, 805, 704, 992, 762, 181, 717, 957, 211, 285, 419, 812, 278, 679, 180, 176, 272, 369, 238, 896, 981, 723, 359, 442, 984, 112, 516, 747, 177, 653, 692, 471, 959, 729, 436, 813, 132, 355, 715, 838, 785, 982, 806, 360, 297, 178, 639, 977, 752, 131, 886, 719, 551, 804, 650, 746, 788, 850, 963, 410, 980, 420, 437, 745, 956, 698, 744, 862, 5, 6, 208]. Quantumbit [915]. Quarks [720]. Quasi [264, 877]. Quasi-Causes [877]. Quasi-Empirical [264]. Quaternions [725]. Quesiti [767]. Quest [561, 316]. Question [751, 871, 277]. Questions [284, 230]. Quine [200].

Radical [829, 580, 159, 163, 160]. Radice [840]. Ramsey [846]. Random [982]. Rare [691]. Ratchet [210, 211]. Rate [711]. Rather [485]. Rational [50, 568]. Rationalism [989]. Rationality [768, 111, 555, 979, 121]. Re [379, 430]. Re-embodiment [379, 430]. Readers [405]. Real [678, 673, 796]. Real-Time [796]. Realism [972, 204, 199, 146, 955, 976, 177, 132, 952, 628, 783, 277, 131, 603, 733, 203, 614, 201, 788, 826, 32, 31]. Realist [882, 343, 980]. Realistic [146, 711, 551, 963, 627]. Reality [852, 707, 999, 976, 543, 882, 97, 523, 353, 416, 434, 133, 522, 534, 687, 165, 811, 698]. Really [709]. Realm [806]. Reason [632, 593]. Reasoning [116, 306, 172, 232, 736, 332, 657, 230, 375, 243, 128]. Reasons [576]. Reassembling [930]. Reciprocal [374]. Recognition [667]. Reconcile [984, 854]. Reconciling [955]. Reconsidering [871]. Reconstruction [1018]. Rectangles [851]. Reducible [752]. reducing [94]. Reduction [76, 78, 81]. Reductionism [845, 84, 79, 103]. Reductionisms [145]. Redux [429]. Reexamination [669]. Reference [323]. Reflection [829]. Reflections [707, 887, 916]. Reformational [918]. Refounded [693].

Reframing [455]. Refuted [761]. Regarding [598]. Regularity [549]. Rejoinder [671, 4]. relate [900]. Related [908, 903, 907, 904]. Relation [101, 454, 569, 233, 940, 251, 155, 91]. Relational [554]. Relationism [611]. Relations [1004, 823]. Relationship [707, 1001, 999, 757, 321, 933, 886]. Relationships [590]. Relativism [761]. Relativistic [930]. Relativities [179]. Relativity [693, 743, 389, 140, 365, 45, 436, 838, 357, 390]. Releasement [618, 617]. Relevance [290, 322, 287]. Relevant [978]. Reloaded [423, 462]. Remakes [676]. Remarks [846, 70]. Remedial [434]. Remedies [843]. Remodeling [244]. René [729]. Renunciation [618]. Replies [405]. Reply [14, 592, 589, 46]. Report [602]. Representation [760, 740, 133, 529, 850, 698, 462]. Representational [710]. Representations [100, 503, 246]. Representing [809]. Research [716, 636, 312, 561, 540, 813, 954, 697, 375, 821, 450]. Researchers [656]. Resemblance [305]. Resolution [315, 126]. Respect [559]. Response [685, 576, 825, 591, 467, 587, 581, 470, 696, 80, 585]. Responses [907]. Responsibilities [559]. Results [909]. Resurrection [564]. Retention [210, 211]. Rethinking [866, 134]. Retributions [533]. Retrieval [890]. Retrocausal [771]. Revealing [524]. Reveals [642]. Reversibility [397]. Review [442, 154]. Reviewed [208]. Reviewing [966]. Revisions [876]. Revisited [298, 200, 589, 764, 979, 257, 928, 144]. Revisiting [675]. Revival [457, 482]. Revolution [111, 555, 182]. Revue [208]. Rhetoric [322]. Rhetorical [202]. Riemann [912, 707]. Rights [904]. Rigid [771]. Rigor [570]. Rigorous [421]. RNAi [315]. Road [243]. Robert [467]. Robot [165]. Robotics [1016, 165]. Robots [881, 1008]. Roger [581]. Rohrlich [77]. Role [275, 997, 376, 825, 907, 358, 458, 222, 337, 638, 125]. Roles [810, 600]. Roman [1021]. Rope [881]. Rotating [1016]. Rotation [915, 361]. Route [490]. Routing [801]. Rumors [730]. Rupture [991]. Russell [964]. Russellian [993].

S [124, 279]. sacred [22]. Said [920]. Salmon [269]. Santa [849]. Sant’Orsola [674]. Sarah [533]. Save [386]. Sawhill [105]. Scale [389, 365, 1018, 296, 357, 390]. Scharff [467]. School [34, 57]. Schools [908]. Schrag [829]. Schrödinger [669]. Schubring [610]. Schwinger [297]. Science [303, 107, 100, 833, 38, 10, 20, 36, 37, 58, 150, 74, 196, 962, 636, 1012, 176, 217, 595, 825, 757, 671, 772, 835, 830, 381, 882, 991, 97, 489, 499, 857, 960, 161, 751, 22, 287, 345, 350, 935, 463, 729, 561, 64, 59, 84, 415, 68, 559, 791, 738, 641, 642, 840, 229, 247, 914, 861, 343, 656, 148, 52, 195, 372, 511, 264, 164, 832, 662, 159, 163, 670, 366, 602, 652, 61, 482, 126, 879, 799, 160, 294, 28, 214, 292, 1, 508]. Sciences [116, 130, 973, 311, 829, 806, 203, 615, 731, 134, 821]. Scientific [348, 972, 235, 934, 426, 636, 137, 976, 768, 149, 32, 289, 952, 400, 738, 60, 642, 354, 608, 666, 783, 890, 277, 239, 557, 148, 994, 131, 601, 427, 373, 245, 120, 706, 856, 439, 27, 563, 246, 758, 853, 81, 162, 840]. Scientificity [784]. Scientist [142]. Scientists [606, 559]. Scope [822, 571, 293]. Screen [385].

Script [502]. Search [95, 774]. Second [803, 802, 641, 642, 814]. Secretion [345, 350]. Secrets [1018]. Securing [797]. Seeing [658]. Seeking [304, 16]. Seen [225]. Selection [101, 210, 211, 217, 541, 792, 409, 831, 410]. Selective [210, 211, 788]. Selectivism [733]. Self [595, 349, 237, 367, 959, 319]. Self-Correcting [595]. Self-Directed [349]. Self-Duality [237]. Self-Organization [367]. Self-Organizing [959]. Selves [629]. Semantic [850, 604, 646, 599]. Semanticists [612]. Seminal [840]. Semiotic [519, 1010, 1005]. Sense [883]. Sensitive [307]. Sensitiveness [970]. Sensor [801]. Sentence [846]. Separability [431]. Serendipitous [481]. Serendipity [531]. Service [839]. Sesto [767]. Set [759, 492]. Sets [301]. Setting [960]. Shared [563, 942]. Sherry [685, 696]. Shift [313]. Shifting [584, 788]. Shoah [678]. Shortcomings [164]. Should [644, 462]. Sign [679, 323, 243]. Significance [185]. Similarity [795, 793]. Simon [13]. Simulated [999]. Simulation [303]. Simulations [934]. Simultaneity [650]. Single [491]. Singularitates [459]. Singularities [459]. Situated [429, 165]. Situation [612, 33]. Situations [186]. Sixth [668]. Skype [434]. Slingshot [535]. Small [940]. Sobot [1008]. Social [130, 196, 1008, 922, 556, 68, 806, 625, 203, 906, 571, 749, 527, 542, 647, 796, 134]. Sociality [1007]. Socially [307]. Society [563]. Sociology [563]. Socrates [462]. Software [791, 770]. soil [94]. Solari [777]. Soliton [711]. Solution [770]. Solutions [964]. Solving [613, 248]. Some [707, 846, 2, 284, 919, 115, 506, 713, 412, 10, 32, 3]. Something [409]. Son [675]. Sorcery [907]. Space [479, 478, 707, 552, 596, 386, 715, 324, 545, 357, 778, 459, 476, 870, 809, 474, 793]. Spaces [572]. Spacetime [686, 443, 611, 687, 720]. Spatial [340, 246, 811]. Spatiality [437]. Spatio [417, 418]. Spatio-Cultural [417, 418]. Special [742, 803, 802, 99, 1012, 903, 998, 791, 832, 262, 356, 960]. Species [365]. Spectrum [906]. Speculations [609, 412]. Speculative [893, 655]. Spin [682, 683]. Spinoza [763]. Spinozism [477]. Spontaneous [482]. Spurious [643]. Squares [776]. Sri [665]. Stability [754]. Stake [560]. Stamp [1013]. Stance [277]. Stances [886]. Standard [716, 818, 963, 713]. Stapp [285]. State [683, 812, 656, 862]. States [978, 169]. Statistic [664]. Statistical [978]. Statisticalist [931]. Statistics [957, 349, 5, 2, 3]. Status [193, 922, 984]. STEM [839]. Steps [312, 367]. Stevin [416]. Stiegler [926, 872]. Stigma [908, 903, 904, 906]. Stigmatisation [909]. Stigmatized [905]. Still [486]. Stimmung [935]. Stimmung/Nastrój [935]. Stop [62]. Storck [674]. Stories [907]. Strange [948]. Strategies [123, 231, 600]. Strategy [908, 104]. Stretched [881]. Stretching [378]. Structural [955, 1020]. Structuralism [756]. Structuralist [949]. Structure [780, 240, 984, 460, 748, 778, 731]. Structured [312]. Structures [943, 249, 186, 663, 667, 791, 754, 544, 698]. Struggle [627]. Stuart [112]. Studies [999, 470, 422, 5, 94]. Study [123, 660, 124, 1006, 364, 909, 354, 608, 666, 851, 115, 990, 111, 315, 635, 126, 125, 128, 582, 119]. Style [415]. Styles [935]. Sub [544]. Sub-systemic [544]. Subjective [824]. Subjectivity [868].

Substances [985]. Substantivalism [611]. Substantivism [938]. Sudarsana [793]. Suffice [622]. Sufficient [593]. Suggestions [153]. Suicide [272]. Sum [680, 776]. Summary [11]. Summer [34, 57]. Superorganisms [819]. Superpose [679]. Superposition [704]. Superpositions [698]. Supertasks [515]. Suppesian [69]. Support [954, 527, 542, 647]. Supported [126]. Surely [536]. Surface [216]. Surreal [674]. Surrealism [603]. Surround [388]. survey [27]. Surveyability [330]. Survival [881, 342]. Sustainability [469, 383, 468]. Sustainable [382, 652, 144]. Switch [66]. Symbolic [249, 313, 502]. Symbolism [507]. Symbolization [505]. Symbols [758]. Symmetric [961]. Symmetries [396]. Symmetry [180, 397, 959, 396, 638, 645]. Symmetry-Breakings [396]. Symposium [742, 803, 802, 35]. Synthesis [87]. Synthetic [1006]. System [303, 817, 710, 794, 571, 401, 758, 796, 514]. Systemic [842, 605, 544]. Systems [496, 572, 342, 295, 403, 567, 498, 781, 786, 824, 715, 711, 659, 273, 658, 691, 178, 522, 566, 726, 216, 569, 605, 422, 1017, 501, 565, 681, 860, 265, 527, 542, 544, 602, 647, 652, 879].

T [871]. Tacit [679]. Take [533]. Takes [889]. Taking [870]. Tale [421, 767]. Talking [473]. Taming [43]. Tanzania [904]. Tanzanian [908]. Target [746]. Tartaglia [767]. Task [513]. Taxonomy [243]. Teach [513]. Teaching [841]. Teams [810]. Technical [878, 629, 632, 431, 894]. Techniques [483]. Technocultural [609]. Technological [308, 990, 380]. Technologies [997, 469, 875, 439]. Technologizing [941]. Technology [500, 454, 997, 883, 949, 876, 379, 432, 880, 382, 634, 449, 866, 871, 929, 383, 470, 839, 950, 453, 920, 637, 618, 870, 616, 617, 627, 468, 947, 900, 874, 913, 616]. Technoscience [921, 727, 928, 873]. Technoscientific [384]. Technosphere [621]. Tectonics [661]. Teleology [845]. Teleportation [896]. Temporal [798, 340, 939, 793]. Ten [438]. Tendencies [414]. Tension [42]. Tensions [40, 44]. Term [169]. Terms [940]. Testimony [557]. Text [795]. Texts [713]. Textual [566]. Thales [779]. Theatre [324, 387]. Their [304, 438, 971, 503, 295]. Theological [695]. Theorem [680, 613, 285]. Theorems [789, 848, 52]. Theoretic [846, 759]. Theoretical [707, 216, 79, 645, 23, 30, 31]. Theories [348, 93, 743, 426, 760, 400, 239, 460, 373, 773, 23, 30, 69, 88]. Theorising [415, 293]. Theorist [967]. Theorizing [773]. Theory [347, 693, 716, 101, 34, 57, 479, 275, 181, 211, 943, 707, 660, 661, 140, 365, 369, 930, 335, 835, 981, 403, 884, 105, 516, 338, 177, 653, 692, 918, 935, 781, 786, 824, 436, 806, 273, 658, 299, 664, 608, 666, 297, 566, 827, 357, 390, 321, 334, 245, 147, 706, 906, 865, 258, 171, 954, 327, 754, 433, 627, 492, 251, 544, 652, 514, 862, 799, 393, 820]. There [812, 769, 770]. Thermodynamics [726, 814]. Thesis [200, 488]. Thing [963, 874]. Thing-Transcendentality [874]. Thingness [804]. Things [960, 350]. Think [843]. Thinking [495, 1001, 880, 653, 449, 673, 871, 899, 472]. Third [765, 152, 574, 257]. Thomas [885]. Thought [724, 995, 935, 540, 148, 763, 938, 104]. Thoughts [273]. Three [220, 282, 558]. Three-dimensional [282]. Throughout [482]. Till [835]. Time [798, 707, 552, 880, 653, 787, 367, 291, 324, 774, 396, 252, 357, 459, 539, 689, 286, 474, 796]. Times [629, 632, 631]. Today [559]. Tom [675, 675, 675]. Too [983]. Tool [232]. Tools [458, 119, 214]. Top [440]. Top-Down [440]. Topic [72]. Topical [465]. Topoi [676]. Topological [566, 778, 816]. Topology [521]. Toraldo [54, 25]. Torsional [863]. Total [420]. Touch [449]. Tracking [301]. Tract [817]. Trade [856]. Trade-Offs [856]. Trading [307]. Tradition [918, 317, 758]. Trans [658]. Trans-Classical [658]. Transcendental [895, 897, 876, 868, 871, 941, 920, 867, 889]. Transcendentality [911, 874]. Transduction [893]. Transfer [137]. Transformation [794]. Transformations [936, 246]. Transition [470]. Transitions [469, 342, 383]. Translation [694]. Transnational [766]. Transparency [380]. Transport [896, 790]. Treatise [1014]. Tree [671]. Triad [850]. trick [22]. Tricks [913]. Triggering [835]. trivial [813]. Tropes [256]. True [95, 783]. Trust [962, 966]. Trusting [350]. Trusts [801]. Truth [842, 290, 523, 971, 775, 818, 961, 877, 23, 30, 24]. Truthlikeness [529]. Truths [727]. Tuned [370]. Tuning [410]. Turing [800, 1010, 1005, 481]. Turn [897, 987, 876, 825, 911, 614, 655, 938, 616, 928]. Turning [897, 616]. Turns [202, 867]. Twentieth [838, 108, 215]. Twenty [838]. Twenty-First [838]. Two [240, 973, 776, 70, 1018, 282, 599, 406, 167]. Two-dimensional [282]. Tylman [671]. Type [789, 848, 60]. Tyranny [862].

U [663]. Ubiquity [210, 211]. Ultimate [237, 367, 798]. Unaffected [168]. Uncertainty [60]. Underdetermination [537, 980]. Undermine [537]. Understanding [550, 633, 78, 808, 858, 518, 909, 394, 367, 781, 729, 672, 68, 289, 354, 216, 511, 601, 276, 159, 906, 370, 544]. Unearthing [798]. Unexpected [490, 563]. Unfalsifiable [110]. Unfolding [945]. Ungrounded [419]. Unification [88]. Unificationist [258]. Unificatory [858]. Unified [854]. Unifying [225]. Uniqueness [494, 580]. Units [636]. Unity [84, 779, 853]. Universal [74, 279, 505]. Universe [765, 352, 690, 358, 545, 226, 368, 371, 356]. Unknown [800]. Unnecessary [394]. Unobservable [269]. Unscrambling [744]. Unseen [676]. Unspeakable [896]. Unveiling [242]. Upgrade [420]. upon [713]. Urgent [881]. Use [630, 234, 950, 947]. Used [1015, 1020]. Uses [632]. Using [801, 797, 305, 745, 799]. UTTAMA [794].

Vacuum [358]. Value [252, 731]. Valued [961]. Values [718, 294, 64, 559, 961, 615]. Variable [568, 879]. Variation [210, 211, 187]. Variation-and-Selective-Retention [210, 211]. Variations [210, 211, 399]. Variety [84]. Vectors [899]. Ventures [463]. Verbal [845]. Verfallenheit [619]. Vergragt [470]. Verification [770]. Veritism [818]. Versa [919]. Version [285]. Versus [908, 392, 584, 480, 855, 603, 199, 214, 125]. via [747, 354]. Vibrantly [665]. Vice [919]. Victoria [905]. Vienna [563]. View [1016, 510, 777, 596, 95, 154, 517, 133, 239, 696, 721, 401, 134, 160]. Viewpoint [729]. Views [93, 90]. Vinci [991, 740]. Violation [863, 272]. Violence [907]. Virosphere [881]. Virtual [999, 446]. Virtues [856]. Vision [675, 519, 646]. Visioning [468]. Visual [704, 248, 333]. Visualizations [246]. Vitruvian [740]. Volume [129, 158, 175, 198, 219, 261, 281]. Voluntarism [591, 590]. Volunteering [579, 584, 578, 582]. Volunteers [573]. vs [194, 975, 69]. Vu [676]. Vulnerability [633].

Waiting [734]. Walker [790]. Walls [388]. War [320]. Warming [657]. Water [1021]. Wave [963]. Waves [686, 436, 687]. Way [981, 172, 465]. Ways [920]. Weak [194]. Weaving [97]. Webcams [386]. Weierstrass [421]. Weighted [801]. Weinberg [79]. Weizsäcker [517]. Well [559, 626]. Well-Being [559]. Wellner [945]. were [463]. Western [758]. Weyl [707]. Whatever [474]. Wheel [841]. Where [230]. Which [688, 975, 623]. Whitrow [545]. Who [421, 467, 598, 465]. Whole [947]. Whose [624]. Will [268]. Williamson [968]. Wind [682, 683]. Wireless [801]. Wisdom [623, 404]. Wit [580, 578]. Withdrawing [110]. Within [461, 341]. Without [976]. Wits [1015]. Wittgenstein [35, 305]. Wójcicki [80]. Wolf [576]. Word [233]. Words [780]. Work [320]. workshop [11]. World [93, 997, 1001, 945, 510, 270, 922, 53, 936, 95, 154, 982, 242, 133, 623, 748, 108, 551, 562, 1000, 1009, 92, 90]. Worldmaking [172]. Worlds [553, 742, 803, 802, 96]. Worldview [97]. Worldviews [903, 907, 582, 91]. Would [727]. Writing [502]. Wrong [939].

Years [734].

Z3 [326]. Zag [964]. Zeilinger [852]. Zeno [699]. Zig [964]. Zones [307]. Zurek [719]. Zwart [921]. Zytkow [82].

References

vanFraassen:1995:PAF

[1] Bas van Fraassen. A philosophical approach to foundations of science. Foundations of Science, 1(1):5–18, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

      com/article/10.1007/BF00208722.

Freedman:1995:SIF

[2] David Freedman. Some issues in the foundation of statistics. Foundations of Science, 1(1):19–39, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/ article/10.1007/BF00208723. See discussion [3] and rejoinder [4].

Berger:1995:DDF

[3] James Berger. Discussion of David Freedman’s “Some issues in the foundations of statistics”. Foundations of Science, 1(1):41–67, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/BF00208724. See [2, 2].

Freedman:1995:R

[4] David Freedman. Rejoinder. Foundations of Science, 1(1):69–83, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic).

URL http://link.springer.com/article/10.1007/BF00208725. See

[3, 2].

Aerts:1995:AQS

[5] Diedrik Aerts and Sven Aerts. Applications of quantum statistics in psychological studies of decision processes. Foundations of Science, 1 (1):85–97, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00208726.

Galavotti:1995:OPQ

[6] Maria Carla Galavotti. Operationism, probability and quantum mechanics. Foundations of Science, 1(1):99–118, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/BF00208727.

Humphreys:1995:CE

[7] Paul Humphreys. Computational empiricism. Foundations of Science, 1(1):119–130, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00208728.

Agassi:1995:BLN

[8] Joseph Agassi. Blame not the laws of nature. Foundations of Science, 1(1):131–154, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00208729.

Anonymous:1995:BNa

[9] Anonymous. Biographical notes. Foundations of Science, 1(1): 155–157, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00208730.

Anonymous:1995:AFSa

[10] Anonymous. On Association for the Foundation of Science, Language and Cognition, AFOS; some history and some recent developments. Foundations of Science, 1(1):159–160, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/BF00208731.

Agassi:1995:SAW

[11] Joseph Agassi. Summary of AFOS workshop, 1994. Foundations of Science, 1(1):161–166, March 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00208732.

Simon:1995:MD

[12] Herbert Simon. Machine discovery. Foundations of Science, 1(2): 171–200, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.

1007/BF00124609. See comment [13] and reply [14].

Boden:1995:CSP

[13] Margaret Boden. Commentary on Simon’s paper on “machine discovery”. Foundations of Science, 1(2):201–224, June 1995. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/BF00124610. See [14, 12].

Anonymous:1995:MDR

[14] Anonymous. Machine discovery: Reply to comments. Foundations of Science, 1(2):225–232, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00124611. See [12, 13].

Shen:1995:PD

[15] Wei-Min Shen. The process of discovery. Foundations of Science, 1 (2):233–251, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00124612.

Zytkow:1995:CDA

[16] Jan M. Zytkow. Creating a discoverer: Autonomous knowledge seeking agent. Foundations of Science, 1(2):253–283, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/BF00124613.

Creath:1995:DE

[17] Richard Creath. Are dinosaurs extinct? Foundations of Science, 1 (2):285–297, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00124614.

Friedman:1995:PCL

[18] Michael Friedman. Poincaré’s conventionalism and the logical positivists. Foundations of Science, 1(2):299–314, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/BF00124615.

Anonymous:1995:BNb

[19] Anonymous. Biographical notes. Foundations of Science, 1(2): 315–316, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00124616.

Anonymous:1995:AFSb

[20] Anonymous. On the Association for Foundations of Science, Language and Cognition, AFOS. Foundations of Science, 1(2):317–318, June 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/BF00124617.

Humphreys:1995:PN

[21] Paul Humphreys and Ryszard Wójcicki. Publisher’s note. Foundations of Science, 1(3):iii, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00145400.

Haack:1995:SNS

[22] Susan Haack. Science is neither sacred nor a confidence trick. Foundations of Science, 1(3):323–335, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/BF00145401.

Wojcicki:1995:TTMa

[23] Ryszard Wójcicki. Theories, theoretical models, truth. Foundations of Science, 1(3):337–406, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00145402.

diFrancia:1995:HT

[24] Guliano Toraldo di Francia. Historical truth. Foundations of Science, 1 (3):407–416, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.

1007/BF00145403. See comments [25].

Eilstein:1995:CGT

[25] Helena Eilstein. Comments on Guliano Toraldo di Francia’s paper. Foundations of Science, 1(3):417–425, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/BF00145404. See [24].

diFrancia:1995:AMD

[26] Guliano Toraldo di Francia. Answers to my discussants. Foundations of Science, 1(3):427–428, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00145405.

Schurz:1995:SEC

[27] Gerhard Schurz. Scientific explanation: a critical survey. Foundations of Science, 1(3):429–465, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00145406.

Heller:1995:LSC

[28] Michal Heller. Limits of science — Cracow, 1996. Foundations of Science, 1(3):467–469, September 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00145407.

Anonymous:1995:PN

[29] Anonymous. Publisher’s note. Foundations of Science, 1(4):ii, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic).

URL http://link.springer.com/article/10.1007/BF00125782.

Wojcicki:1995:TTMb

[30] Ryszard Wójcicki. Theories, theoretical models, truth. Foundations of Science, 1(4):471–516, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00125783.

Zeidler:1995:IRN

[31] Paweł Zeidler and Danuta Sobczyńska. The idea of realism in the new experimentalism and the problem of the existence of theoretical entities in chemistry. Foundations of Science, 1(4):517–535, December 1995. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/BF00125784.

Hacking:1995:SRA

[32] Ian Hacking. Scientific realism about some chemical entities. Foundations of Science, 1(4):537–542, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00125785.

Shafer:1995:SC

[33] Glenn Shafer. The situation of causality. Foundations of Science, 1(4): 543–563, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00125786.

Anonymous:1995:SST

[34] Anonymous. ’97 Summer School for Theory of Knowledge. Foundations of Science, 1(4):565, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/BF00125787.

Anonymous:1995:IWS

[35] Anonymous. The 20th International Wittgenstein Symposium. Foundations of Science, 1(4):567, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/BF00125788.

Anonymous:1995:AFSc

[36] Anonymous. On the Association for Foundations of Science, Language and Cognition, AFOS. Foundations of Science, 1(4):569–570, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic).

URL http://link.springer.com/article/10.1007/BF00125789.

Anonymous:1995:CAF

[37] Anonymous. Constitution of the association for the foundations of science, language and cognition. Foundations of Science, 1(4):571– 576, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/BF00125790.

Anonymous:1995:AAF

[38] Anonymous. AFOS the Association for Foundations of Science, Language and Cognition. Foundations of Science, 1(4):577–588, December 1995. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/BF00125791.

Wojcicki:1997:IEN

[39] Ryszard Wójcicki. Instead of an editorial note. Foundations of Science, 2(1):5–10, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009685021757.

Duda:1997:MET

[40] Roman Duda. Mathematics: Essential tensions. Foundations of Science, 2(1):11–19, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009654622666.

Brown:1997:WAM

[41] James Robert Brown. What is applied mathematics? Foundations of Science, 2(1):21–37, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009658723574.

Heller:1997:ETM

[42] Michael Heller. Essential tension: Mathematics–physics–philosophy. Foundations of Science, 2(1):39–52, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009662824483.

Moore:1997:TI

[43] A. W. Moore. Taming the infinite. Foundations of Science, 2(1):53– 56, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009666925392.

Sankaran:1997:IT

[44] G. K. Sankaran. Inessential tensions. Foundations of Science, 2(1):57– 60, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009671009462.

Chao:1997:BGR

[45] Wu Zhong Chao. The beauty of general relativity. Foundations of Science, 2(1):61–64, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009623126300.

Duda:1997:RC

[46] Roman Duda. Reply to comments. Foundations of Science, 2(1):65– 66, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009675127209.

Fleischhacker:1997:MMG

[47] Louk Fleischhacker. Mathematics and the mind of God. Foundations of Science, 2(1):67–72, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009679311279.

Lasota:1997:DIM

[48] Andrezej Lasota. Determinism, indeterminism and mathematics. Foundations of Science, 2(1):73–75, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009631428118.

Mrozek:1997:EMP

[49] Marian Mrozek and Jacek Urbaniec. Evolution of mathematical proof. Foundations of Science, 2(1):77–85, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009683412188.

Marciszewski:1997:RBP

[50] Witold Marciszewski. Rational beliefs as produced by computational processes. Foundations of Science, 2(1):87–106, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009635529026.

Flasinski:1997:EMH

[51] Mariusz Flasiński. “Every man in his notions” or alchemists’ discussion on artificial intelligence. Foundations of Science, 2(1):107–121, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009687513096.

Murawski:1997:GIT

[52] Roman Murawski. Gödel’s incompleteness theorems and computer science. Foundations of Science, 2(1):123–135, ???? 1997. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009639629935.

Demaret:1997:LGP

[53] Jacques Demaret, Michael Heller, and Dominique Lambert. Local and global properties of the world. Foundations of Science, 2(1):137–176, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009691614005.

Anonymous:1997:IGT

[54] Anonymous. An interview with Giuliano Toraldo Di Francia. Foundations of Science, 2(1):177–182, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1017184330843.

Anonymous:1997:IAL

[55] Anonymous. An interview with Andrzej Lasota. Foundations of Science, 2(1):183–189, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1017195714914.

Anonymous:1997:BNa

[56] Anonymous. Biographical notes. Foundations of Science, 2(1):191–194, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017147831752.

Anonymous:1997:FSS

[57] Anonymous. First Summer School for Theory of Knowledge. Foundations of Science, 2(1):195, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1017199815822.

Anonymous:1997:AFS

[58] Anonymous. On the Association for Foundations of Science, Language and Cognition, AFOS. Foundations of Science, 2(1):197–198, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1017151932661.

Kaiser:1997:PPI

[59] Matthias Kaiser. “The precautionary principle and its implications for science” — introduction. Foundations of Science, 2(2):201–205, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009641701933.

Lemons:1997:PPS

[60] John Lemons, Kristin Shrader-Frechette, and Carl Cranor. The precautionary principle: Scientific uncertainty and Type I and Type II errors. Foundations of Science, 2(2):207–236, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009611419680.

Westra:1997:PNS

[61] Laura Westra. Post-normal science, the precautionary principle and the ethics of integrity. Foundations of Science, 2(2):237–262, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009667603750.

Martin:1997:IYD

[62] Philippe H. Martin. “If you don’t know how to fix it, please stop breaking it!” the precautionary principle and climate change. Foundations of Science, 2(2):263–292, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009619720589.

Hansson:1997:LP

[63] Sven Ove Hansson. The limits of precaution. Foundations of Science, 2(2):293–306, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009671722406.

Kaiser:1997:FFP

[64] Matthias Kaiser. Fish-farming and the precautionary principle: Context and values in environmental science for policy. Foundations of Science, 2(2):307–341, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009679923315.

Grobler:1997:PFI

[65] Adam Grobler. Popperian forum: Introduction. Foundations of Science, 2(2):349–350, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009643701045.

Grobler:1997:PSJ

[66] Adam Grobler. Popper’s switch from justification to critical preferences. Foundations of Science, 2(2):350–354, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009684024223.

Childers:1997:PNF

[67] Timothy Childers. Popper on naturalism and the foundations of methodology. Foundations of Science, 2(2):355–360, ???? 1997. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009688108294.

Koertge:1997:PCO

[68] Noretta Koertge. Popper’s contributions to our understanding of social science. Foundations of Science, 2(2):365–370, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009640225132.

Wojcicki:1997:PVS

[69] Ryszard Wójcicki. Popperian vs. Suppesian theories. Foundations of Science, 2(2):375–382, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009692210111.

Eilstein:1997:TR

[70] Helena Eilstein. Two remarks. Foundations of Science, 2(2):389–392, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009648511019.

Anonymous:1997:BNb

[71] Anonymous. Biographical notes. Foundations of Science, 2(2):397–399, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017166111928.

Anonymous:1997:CPT

[72] Anonymous. Call for projects of topic issues. Foundations of Science, 2 (2):400, ???? 1997. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017156728766.

Wojcicki:1998:BPI

[73] Ryszard Wójcicki. “Bialkowski Panel” — introduction and acknowledgements. Foundations of Science, 3(1):5–7, ???? 1998. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009604100879.

Bialkowski:1998:PUS

[74] Grzegorz Bialkowski. Is physics an universal science? Foundations of Science, 3(1):9–21, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009621917717.

Pukas:1998:GB

[75] Sergiusz Pukas. Grzegorz Bialkowski. Foundations of Science, 3(1):23– 25, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009673901787.

Aerts:1998:R

[76] Diederik Aerts and Fritz Rohrlich. Reduction. Foundations of Science, 3(1):27–35, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009626018626. See comments [77].

Wojcicki:1998:RPO

[77] Ryszard Wójcicki. Rohrlich’s pluralistic ontology; comments on D. Aerts and F. Rohrlich article. Foundations of Science, 3(1):37–43, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009678002696. See

[76].

Dieks:1998:RU

[78] Dennis Dieks and Henk W. de Regt. Reduction and understanding. Foundations of Science, 3(1):45–59, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009630119534.

Wojcicki:1998:PTK

[79] Ryszard Wójcicki. Physics, theoretical knowledge and Weinberg’s grand reductionism. Foundations of Science, 3(1):61–77, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009682103604.

Weinberg:1998:RPW

[80] Steven Weinberg. A response to Professor Wójcicki. Foundations of Science, 3(1):79–81, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009634220443.

Zytkow:1998:CDS

[81] Jan M. Zytkow. Consciousness defies scientific and computational reduction. Foundations of Science, 3(1):83–101, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http:

//link.springer.com/article/10.1023/A:1009686204513. See com-

ments [82].

Horst:1998:CZA

[82] Steve Horst. Comments on Zytkow’s article. Foundations of Science, 3(1):103–109, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009638321351. See [81].

Brown:1998:WD

[83] James Robert Brown. What is a definition? Foundations of Science, 3(1):111–132, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009690306330.

Kaluszynska:1998:RCS

[84] Elzbieta Kaluszyńska. Reductionism in contemporary science; unity of nature, variety of events. Foundations of Science, 3(1):133–150, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009646624077.

Galles:1998:ACC

[85] David Galles and Judea Pearl. An axiomatic characterization of causal counterfactuals. Foundations of Science, 3(1):151–182, ???? 1998. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009602825894.

Psarros:1998:WPO

[86] Nikos Psarros. What has philosophy to offer to chemistry? Foundations of Science, 3(1):183–202, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009611009965.

Aerts:1998:ESA

[87] Diederik Aerts. Editorial: Synthesis and analysis, interdisciplinarity and foundations. Foundations of Science, 3(2):203–206, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009663005061.

Wojtowicz:1998:UMT

[88] Krzysztof Wójtowicz. Unification of mathematical theories. Foundations of Science, 3(2):207–229, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009665705969.

Weber:1998:I

[89] Erik Weber. Introduction. Foundations of Science, 3(2):231–234, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009669806878.

tBroekaer:1998:WVE

[90] Jan ’t Broekaer. World views. elements of the Apostelian and general approach. Foundations of Science, 3(2):235–258, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009673907786.

Batens:1998:PRW

[91] Diderik Batens. Paraconsistency and its relation to worldviews. Foundations of Science, 3(2):259–283, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009678125533.

Weber:1998:WMI

[92] Erik Weber and Wim Christiaens. World models and inconsistencies. Foundations of Science, 3(2):285–311, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009634309604.

Aerts:1998:ICT

[93] Diederik Aerts, Jan Broekaert, and Sonja Smets. Inconsistencies in constituent theories of world views: Quantum mechanical examples. Foundations of Science, 3(2):313–340, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009686326442.

Maier:1998:LIR

[94] Robert Maier. Localization of iron-reducing activity in paddy soil by profile studies. Foundations of Science, 3(2):341–357, ???? 1998. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009650720124.

Havas:1998:DWN

[95] Katalin G. Havas. Do we need to search for the only true world view? Foundations of Science, 3(2):359–373, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009638410512.

Havel:1998:LCW

[96] Ivan M. Havel. Living in conceivable worlds. Foundations of Science, 3(2):375–394, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009690428259.

Gabora:1998:WBP

[97] Liane Gabora. Weaving, bending, patching, mending the fabric of reality: a cognitive science perspective on worldview inconsistency. Foundations of Science, 3(2):395–428, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009646612330.

Asenjo:1998:GCA

[98] F. Gonzalez Asenjo. The general concept of antinomicity. Foundations of Science, 3(2):429–465, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009698629168.

Anonymous:1998:CPS

[99] Anonymous. Call for proposals for special issues. Foundations of Science, 3(2):467, ???? 1998. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1017218000930.

Aerts:1999:EFI

[100] Diederik Aerts, Jan Broekaert, and Liane Gabora. Editorial: Formal and informal representations of science. Foundations of Science, 4(1):1– 2, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009664701532.

Agutter:1999:FBP

[101] Paul S. Agutter and Denys N. Wheatley. Foundations of biology: On the problem of “purpose” in biology in relation to our acceptance of the Darwinian theory of natural selection. Foundations of Science, 4(1):3– 23, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009634718370.

Heylighen:1999:ALF

[102] Francis Heylighen. Advantages and limitations of formal expression. Foundations of Science, 4(1):25–56, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009686703349.

Edmonds:1999:PHP

[103] Bruce Edmonds. Pragmatic holism (or pragmatic reductionism). Foundations of Science, 4(1):57–82, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009642920187.

diFrancia:1999:DBS

[104] Giuliano Toraldo di Francia. Deep Blue and the strategy of human thought. Foundations of Science, 4(1):83–93, ???? 1999. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009694906075.

Gabora:1999:CTP

[105] Liane Gabora. Complexity from theory to practise: an interview with Bruce Sawhill. Foundations of Science, 4(1):95–105, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009655307892.

Anonymous:1999:BNa

[106] Anonymous. Biographical notes. Foundations of Science, 4(1):107–108, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017295211917.

Aerts:1999:ECS

[107] Diederik Aerts. Editorial: Creativity and science. Foundations of Science, 4(2):111–112, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009610911185.

Rohrlich:1999:PTW

[108] Fritz Rohrlich. Papers from the Twentieth World Congress of Philosophy. Foundations of Science, 4(2):113–114, ???? 1999. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009660925413.

Aerts:1999:LPQ

[109] Diederik Aerts, Jan Broekaert, and Sonja Smets. The liar-paradox in a quantum mechanical perspective. Foundations of Science, 4(2):115– 132, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009610326206.

Magnani:1999:WUH

[110] Lorenzo Magnani. Withdrawing unfalsifiable hypotheses. Foundations of Science, 4(2):133–153, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009667028023.

Nugayev:1999:ERC

[111] Rinat M. Nugayev. Einstein’s revolution: a case study in communicative rationality. Foundations of Science, 4(2):155–204, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009619128932.

Gabora:1999:MAQ

[112] Liane Gabora. Microtubules, anesthetics, and quantum consciousness: an interview with Stuart Hameroff. Foundations of Science, 4(2):205– 223, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009623213002.

Anonymous:1999:E

[113] Anonymous. Errata. Foundations of Science, 4(2):225, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/
A:1017215013911;   http://link.springer.com/content/pdf/10.
1023/A:1017215013911.pdf.

Anonymous:1999:BNb

[114] Anonymous. Biographical notes. Foundations of Science, 4(2):227–229, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017288825298.

Meheus:1999:MSC

[115] Joke Meheus and Thomas Nickles. The methodological study of creativity and discovery — some background. Foundations of Science, 4 (3):231–235, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1009663506333.

Abrantes:1999:ARM

[116] Paulo Abrantes. Analogical reasoning and modeling in the sciences. Foundations of Science, 4(3):237–270, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009623823171.

Hoffmann:1999:PPC

[117] Michael Hoffmann. Problems with Peirce’s concept of abduction. Foundations of Science, 4(3):271–305, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009675824079.

Kuipers:1999:AAE

[118] Theo A. F. Kuipers. Abduction aiming at empirical progress or eventruth approximation leading to a challenge for computational modelling. Foundations of Science, 4(3):307–323, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1009688227714.

Meheus:1999:DAA

[119] Joke Meheus. Deductive and ampliative adaptive logics as tools in the study of creativity. Foundations of Science, 4(3):325–336, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009680025897.

Rebaglia:1999:SDB

[120] Alberta Rebaglia. Scientific discovery: Between incommensurability of paradigms and historical continuity. Foundations of Science, 4(3):337– 355, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1009696428623.

Weber:1999:RDE

[121] Erik Weber. Rationality in the discovery of empirical laws. Foundations of Science, 4(3):357–370, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009600614510.

Meheus:1999:IN

[122] Joke Meheus and Thomas Nickles. Introductory note. Foundations of Science, 4(4):373–374, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009675215728.

Bailer-Jones:1999:CSE

[123] D. M. Bailer-Jones. Creative strategies employed in modelling: a case study. Foundations of Science, 4(4):375–388, ???? 1999. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009687632566.

Bruning:1999:CPC

[124] R. Brüning and G. Lohmann. Charles S. Peirce on creative metaphor: a case study on the conveyor belt metaphor in oceanography. Foundations of Science, 4(4):389–403, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009639700707.

deRegt:1999:PVH

[125] Henk W. de Regt. Pauli versus Heisenberg: a case study of the heuristic role of philosophy. Foundations of Science, 4(4):405–426, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1009643801616.

deJong:1999:CSR

[126] Hidde de Jong, Nicolaas Mars, and Paul van der Vet. Computersupported resolution of measurement conflicts: a case-study in materials science. Foundations of Science, 4(4):427–461, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1009647918454.

Pericliev:1999:PMD

[127] Vladimir Pericliev. The prospects for machine discovery in linguistics. Foundations of Science, 4(4):463–482, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1009699902524.

vandenBosch:1999:IBM

[128] Alexander P. M. van den Bosch. Inference to the best manipulation — a case study of qualitative reasoning in neuropharmacy. Foundations of Science, 4(4):483–495, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1009652019363.

Anonymous:1999:CV

[129] Anonymous. Contents of volume 4. Foundations of Science, 4(4):507– 510, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017296706283.

Aerts:2000:EQM

[130] Diederik Aerts. Editorial: Quantum, mimesis and the social sciences. Foundations of Science, 5(1):1–2, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1026511218440.

Norris:2000:QNC

[131] Christopher Norris. Quantum nonlocality and the challenge to scientific realism. Foundations of Science, 5(1):3–45, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1026447021096.

Kosso:2000:QMR

[132] Peter Kosso. Quantum mechanics and realism. Foundations of Science, 5(1):47–60, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1026563220257.

Mathijs:2000:MRR

[133] Ernest Mathijs and Bert Mosselmans. Mimesis and the representation of reality: a historical world view. Foundations of Science, 5(1):61– 102, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1026473504257.

vanLangenhove:2000:RSS

[134] Luk van Langenhove. Rethinking the social sciences? A point of view. Foundations of Science, 5(1):103–118, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1026550901601.

Anonymous:2000:BNa

[135] Anonymous. Biographical notes. Foundations of Science, 5(1):119–120, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1026518015714.

Magnani:2000:P

[136] Lorenzo Magnani, Nancy J. Nersessian, and Paul Thagard. Preface. Foundations of Science, 5(2):121–127, ???? 2000. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1011327402056;   http:/
/link.springer.com/content/pdf/10.1023/A:1011327402056.pdf.

Cornuejols:2000:NMT

[137] Antoine Cornuéjols, Andrée Tiberghien, and Gérard Collet. A new mechanism for transfer between conceptual domains in scientific discovery and education. Foundations of Science, 5(2):129–155, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1011398802965.

Amigoni:2000:MAC

[138] Francesco Amigoni, Viola Schiaffonati, and Marco Somalvico. A multilevel architecture of creative dynamic agency. Foundations of Science, 5(2):157–184, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011346818894.

Hernandez-Orallo:2000:ECA

[139] José Hernández-Orallo and Ismael García-Varea. Explanatory and creative alternatives to the MDL principle. Foundations of Science, 5 (2):185–207, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011350914776.

Cerroni:2000:CIC

[140] Andrea Cerroni. Covariance/invariance: a cognitive heuristic in Einstein’s relativity theory formation. Foundations of Science, 5(2):209– 224, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1011304431615.

Pennock:2000:CDM

[141] Robert T. Pennock. Can Darwinian mechanisms make novel discoveries?: Learning from discoveries made by evolving neural networks. Foundations of Science, 5(2):225–238, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011356415685.

Viale:2000:DPL

[142] Riccardo Viale and Daniel Osherson. The diversity principle and the little scientist hypothesis. Foundations of Science, 5(2):239–253, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1011308600664.

Anonymous:2000:BNb

[143] Anonymous. Biographical notes. Foundations of Science, 5(2):255–259, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017321706580.

deAvila-Pires:2000:CSD

[144] Fernando Dias de Avila-Pires, Luiz Carlos Mior, Vilênia Porto Aguiar, and Susana Regina de Mello Schlemper. The concept of sustainable development revisited. Foundations of Science, 5(3):261–268, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1011327612400.

Wimsatt:2000:ENA

[145] William C. Wimsatt. Emergence as non-aggregativity and the biases of reductionisms. Foundations of Science, 5(3):269–297, ???? 2000. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1011342202830.

Batitsky:2000:RLP

[146] Vadim Batitsky. A realistic look at Putnam’s argument against realism. Foundations of Science, 5(3):299–321, ???? 2000. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1011323103739.

Redko:2000:ECT

[147] Vladimir G. Red’ko. Evolution of cognition: Towards the theory of origin of human logic. Foundations of Science, 5(3):323–338, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1011327204647.

Mosselmans:2000:HCS

[148] Bert Mosselmans and Ernest Mathijs. Human culture and science: Equality and inequality as foundations of scientific thought. Foundations of Science, 5(3):339–378, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011331321486.

Edmonds:2000:CSM

[149] Bruce Edmonds. Complexity and scientific modelling. Foundations of Science, 5(3):379–390, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1011383422394.

Anonymous:2000:NSP

[150] Anonymous. Nonlinear science and the psyche. Foundations of Science, 5(3):391–393, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1017311611936.

Anonymous:2000:BNc

[151] Anonymous. Biographical notes. Foundations of Science, 5(3):395–398, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017328128774.

Goudsmit:2000:CMO

[152] Arno L. Goudsmit. On the construction of mental objects in third and in first persons. Foundations of Science, 5(4):399–428, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1011319826957.

Vaneechoutte:2000:EAC

[153] Mario Vaneechoutte. Experience, awareness and consciousness: Suggestions for definitions as offered by an evolutionary approach. Foundations of Science, 5(4):429–456, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1011371811027.

Heylighen:2000:FME

[154] Francis Heylighen. Foundations and methodology for an evolutionary world view: a review of the Principia Cybernetica Project. Foundations of Science, 5(4):457–490, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1011389509210.

Weinert:2000:CAM

[155] Friedel Weinert. The construction of atom models: Eliminative inductivism and its relation to falsificationism. Foundations of Science, 5 (4):491–531, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011315710119.

Anonymous:2000:BNd

[156] Anonymous. Biographical notes. Foundations of Science, 5(4):533–534, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017390908302.

Anonymous:2000:IA

[157] Anonymous. Instructions for authors. Foundations of Science, 5(4):535– 544, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017322011234.

Anonymous:2000:CV

[158] Anonymous. Contents of volume 5. Foundations of Science, 5(4):545– 546, ???? 2000. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017214019787.

Riegler:2001:TRC

[159] Alexander Riegler. Towards a radical constructivist understanding of science. Foundations of Science, 6(1–3):1–30, ???? 2001. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1011305022115.

vonGlasersfeld:2001:RCV

[160] Ernst von Glasersfeld. The radical constructivist view of science. Foundations of Science, 6(1–3):31–43, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011345023932.

Glanville:2001:OS

[161] Ranulph Glanville. An observing science. Foundations of Science, 6 (1–3):45–75, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011353225749.

deZeeuw:2001:CNA

[162] Gerard de Zeeuw. Constructivism: a ‘next’ area of scientific development? Foundations of Science, 6(1–3):77–98, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1011361409819.

Stewart:2001:RCB

[163] John Stewart. Radical constructivism in biology and cognitive science. Foundations of Science, 6(1–3):99–124, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1011313532109.

Peschl:2001:CCS

[164] Markus F. Peschl. Constructivism, cognition, and science — an investigation of its links and possible shortcomings. Foundations of Science, 6 (1–3):125–161, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011390200250.

Ziemke:2001:CRR

[165] Tom Ziemke. The construction of ‘reality’ in the robot: Constructivist perspectives on situated artificial intelligence and adaptive robotics. Foundations of Science, 6(1–3):163–233, ???? 2001. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1011394317088.

Anonymous:2001:BNa

[166] Anonymous. Biographical notes. Foundations of Science, 6(1–3):235– 237, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017368201159.

Riegler:2001:PPT

[167] Alexander Riegler. Preface to Part Two. Foundations of Science, 6(4): 239, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1011669518906.

Schwegler:2001:PDU

[168] Helmut Schwegler. Physics develops unaffected by constructivism. Foundations of Science, 6(4):241–253, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011606803885.

Grossing:2001:CLT

[169] Gerhard Grössing. Comparing the long-term evolution of “cognitive invariances” in physics with a dynamics in states of consciousness. Foundations of Science, 6(4):255–272, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011654702976.

Diettrich:2001:PAC

[170] Olaf Diettrich. A physical approach to the construction of cognition and to cognitive evolution. Foundations of Science, 6(4):273–341, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1011610921631.

Scott:2001:GPC

[171] Bernard Scott. Gordon Pask’s conversation theory: a domain independent constructivist model of human knowing. Foundations of Science, 6(4):343–360, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1011667022540.

Fischer:2001:ARW

[172] Hans Rudi Fischer. Abductive reasoning as a way of worldmaking. Foundations of Science, 6(4):361–383, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1011671106610.

Anonymous:2001:BNb

[173] Anonymous. Biographical notes. Foundations of Science, 6(4):385–387, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017398407519.

Anonymous:2001:IA

[174] Anonymous. Instructions for authors. Foundations of Science, 6(4):389– 398, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017462803282.

Anonymous:2001:CV

[175] Anonymous. Contents of volume 6. Foundations of Science, 6(4):399– 400, ???? 2001. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1017369602668.

Chiara:2002:QMO

[176] Maria Luisa Dalla Chiara. Quantum mechanics and other fields of science. Foundations of Science, 7(1–2):1–9, June 2002. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1016087222412.

Ghirardi:2002:MQT

[177] GianCarlo Ghirardi. Making quantum theory compatible with realism. Foundations of Science, 7(1–2):11–47, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1016039307391.

McCollum:2002:SLS

[178] Gin McCollum. Systems of logical systems: Neuroscience and quantum logic. Foundations of Science, 7(1–2):49–72, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1016047510116.

Mugur-Schachter:2002:ODR

[179] Mioara Mugur-Schächter. Objectivity and descriptional relativities. Foundations of Science, 7(1–2):73–180, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1016095424229.

Castellani:2002:SQM

[180] Elena Castellani. Symmetry, quantum mechanics, and beyond. Foundations of Science, 7(1–2):181–196, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1016035105573.

Beltrametti:2002:QMO

[181] E. G. Beltrametti and S. Bugajski. Quantum mechanics and operational probability theory. Foundations of Science, 7(1–2):197–212, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1016007827863.

Voorhees:2002:CCR

[182] Burton Voorhees. Completing the Copernican revolution. Foundations of Science, 7(1–2):213–227, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1016066221503.

Anonymous:2002:BNa

[183] Anonymous. Biographical notes. Foundations of Science, 7(1–2):229– 231, June 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1016063929681.

Edmonds:2002:ECC

[184] Bruce Edmonds and Varol Akman. Editorial: Context in context. Foundations of Science, 7(3):233–238, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/A:1019614708997.

Bazzanella:2002:SCC

[185] Carla Bazzanella. The significance of context in comprehension: The ‘we case’. Foundations of Science, 7(3):239–254, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1019657025835.

Fetzer:2002:MSM

[186] Anita Fetzer. Micro situations and macro structures: Natural-language communication and context. Foundations of Science, 7(3):255–291, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1019609109906.

Heylighen:2002:VCL

[187] Francis Heylighen and Jean-Marc Dewaele. Variation in the contextuality of language: an empirical measure. Foundations of Science, 7(3): 293–340, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1019661126744.

Perconti:2002:CDH

[188] Pietro Perconti. Context-dependence in human and animal communication. Foundations of Science, 7(3):341–362, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1019613210814.

Anonymous:2002:BNb

[189] Anonymous. Biographical notes. Foundations of Science, 7(3):363–364, September 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1019665327653.

Rouchier:2002:BCE

[190] Juliette Rouchier, Martin O’Connor, and Mélanie Requier-Desjardins. Building context in everyday life. Foundations of Science, 7(4):367–392, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1020714428561.

Walczak:2002:CBC

[191] Steven Walczak. A context-based computational model of language acquisition by infants and children. Foundations of Science, 7(4):393–411, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1020721512631.

Ekins:2002:PEC

[192] Sean Ekins and Richard J. McGowan. Postgraduate education and the changing interaction with the pharmaceutical industry: a cross-cultural perspective. Foundations of Science, 7(4):413–424, December 2002. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1020773529470.

Boniolo:2002:NPS

[193] Giovanni Boniolo, Carlo Petrovich, and Gualtiero Pisent. Notes on the philosophical status of nuclear physics. Foundations of Science, 7(4): 425–452, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1020725613540.

Cirkovic:2002:AFV

[194] Milan M. Ćirković. Anthropic fluctuations vs. weak anthropic principle. Foundations of Science, 7(4):453–463, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1020777630378.

Napinen:2002:CCC

[195] Leo Näpinen and Peeter Müürsepp. The concept of chaos in contemporary science: on Jean Bricmont’s critique of Ilya Prigogine’s ideas. Foundations of Science, 7(4):465–479, December 2002. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1020729714449.

Bobryk:2002:SCM

[196] Jerzy Bobryk. The social construction of mind and the future of cognitive science. Foundations of Science, 7(4):481–495, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1020781731287.

Anonymous:2002:BNc

[197] Anonymous. Biographical notes. Foundations of Science, 7(4):497–500, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1020733815357.

Anonymous:2002:CV

[198] Anonymous. Contents of volume 7. Foundations of Science, 7(4):501–512, December 2002. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1020781204665.

Baert:2003:RVP

[199] Patrick Baert. Realism versus pragmatism — an introduction. Foundations of Science, 8(1):1–2, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1022491321867.

Boylan:2003:PEM

[200] Thomas A. Boylan and Paschal F. O’Gorman. Pragmatism in economic methodology: The Duhem–Quine thesis revisited. Foundations of Science, 8(1):3–21, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1022417025502.

Shipman:2003:PRE

[201] Alan Shipman. Pragmatism, realism and the economist/economy divide. Foundations of Science, 8(1):23–50, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/A:1022433410480.

Lewis:2003:RDE

[202] P. A. Lewis. Recent developments in economic methodology: The rhetorical and ontological turns. Foundations of Science, 8(1):51–68, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1022489528227.

Pleasants:2003:PSS

[203] Nigel Pleasants. A philosophy for the social sciences: Realism, pragmatism, or neither? Foundations of Science, 8(1):69–87, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1022441619760.

Baert:2003:PRH

[204] Patrick Baert. Pragmatism, realism and hermeneutics. Foundations of Science, 8(1):89–106, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/A:1022445814115.

Anonymous:2003:BNa

[205] Anonymous. Biographical notes. Foundations of Science, 8(1):107–108, March 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1022406130953.

Belousek:2003:FOM

[206] Darrin W. Belousek. Formalism, ontology and methodology in Bohmian mechanics. Foundations of Science, 8(2):109–172, June 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1023925900377.

Bigaj:2003:IAN

[207] Tomasz Bigaj. The indispensability argument — a new chance for empiricism in mathematics? Foundations of Science, 8(2):173–200, June 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1023/A:1023913316307.

Piron:2003:PQR

[208] C. Piron. Physique quantique revue et completée. (French) [Quantum physics reviewed and completed]. Foundations of Science, 8(2):201– 212, June 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1023978017216.

Anonymous:2003:BNb

[209] Anonymous. Biographical notes. Foundations of Science, 8(2):213– 214, June 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1023930118124.

Bickhard:2003:VVSa

[210] Mark H. Bickhard and Donald T. Campbell. Variations in variation and selection: The ubiquity of the variation-and-selective-retention ratchet in emergent organizational complexity. Foundations of Science, 8(3): 215–282, September 2003. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1025046917589.

Bickhard:2003:VVSb

[211] Mark H. Bickhard. Variations in variation and selection: The ubiquity of the variation-and-selective-retention ratchet in emergent organizational complexity, Part II: Quantum field theory. Foundations of Science, 8(3): 283–293, September 2003. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1025002218498.

Dent:2003:IMA

[212] Eric B. Dent. The interactional model: an alternative to the direct cause and effect construct for mutually causal organizational phenomena. Foundations of Science, 8(3):295–314, September 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1025006302568.

Anonymous:2003:BNc

[213] Anonymous. Biographical notes. Foundations of Science, 8(3):315–316, September 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1025058319407.

Pietarinen:2003:GFT

[214] Ahti-Veikko Pietarinen. Games as formal tools versus games as explanations in logic and science. Foundations of Science, 8(4):317–364, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1026319711838.

Sulcs:2003:NLT

[215] Sue Sulcs. The nature of light and Twentieth Century experimental physics. Foundations of Science, 8(4):365–391, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1023/A:1026323203487.

Neuman:2003:HDS

[216] Yair Neuman. How deep is the surface? A theoretical framework for understanding meaning–making in living systems. Foundations of Science, 8 (4):393–415, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1026371128677.

Cirkovic:2003:NBA

[217] Milan M. Ćirković. Nick Bostrom, anthropic bias: Observation selection effects in science and philosophy. Foundations of Science, 8(4): 417–423, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/A:1026375105749.

Anonymous:2003:BNd

[218] Anonymous. Biographical notes. Foundations of Science, 8(4):425–436, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1026307823716.

Anonymous:2003:CV

[219] Anonymous. Contents of volume 8. Foundations of Science, 8(4):437–438, December 2003. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/A:
1026355814755.

Cirkovic:2004:TCC

[220] Milan M. Ćirković and Vesna Milosević-Zdjelar. Three’s a crowd: On causes, entropy and physical eschatology. Foundations of Science, 9(1): 1–24, March 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/B:
FODA.0000014883.92569.1c.

Hon:2004:GEM

[221] Giora Hon. Gödel, Einstein, Mach: Casting constraints on allembracing concepts. Foundations of Science, 9(1):25–64, March 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000014880.52117.f0.

Kluve:2004:RCI

[222] Jochen Kluve. On the role of counterfactuals in inferring causal effects. Foundations of Science, 9(1):65–101, March 2004. CODEN FOSCFI.

ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/B:FODA.0000014881.82061.7b.

Anonymous:2004:BNa

[223] Anonymous. Biographical notes. Foundations of Science, 9(1):103–104, March 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000014993.56060.1d.

Lombardi:2004:WI

[224] Olimpia Lombardi. What is information? Foundations of Science, 9 (2):105–134, June 2004. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/B:FODA.0000025034.53313.7c.

Lofgren:2004:UFS

[225] Lars Löfgren. Unifying foundations — to be seen in the phenomenon of language. Foundations of Science, 9(2):135–189, June 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000025037.28557.ea.

McAllister:2004:ACN

[226] James W. McAllister. Absence of contingency in the Newtonian Universe. Foundations of Science, 9(2):191–210, June 2004. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/B:FODA.0000025038.33021.0f.

Anonymous:2004:BNb

[227] Anonymous. Biographical notes. Foundations of Science, 9(2):211–212, June 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000025098.97707.26.

Magnani:2004:P

[228] Lorenzo Magnani and Nancy J. Nersessian. Preface. Foundations of Science, 9(3):213–218, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/B:FODA.0000043010.90996.aa.

Magnani:2004:MBM

[229] Lorenzo Magnani. Model-based and manipulative abduction in science. Foundations of Science, 9(3):219–247, September 2004. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/B:FODA.0000042841.18507.22.

Sintonen:2004:RHW

[230] Matti Sintonen. Reasoning to hypotheses: Where do questions come? Foundations of Science, 9(3):249–266, September 2004. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1023/B:FODA.0000042842.55251.c1.

Paavola:2004:ALM

[231] Sami Paavola. Abduction as a logic and methodology of discovery: the importance of strategies. Foundations of Science, 9(3):267–283, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000042843.48932.25.

Hoffmann:2004:HGI

[232] Michael H. G. Hoffmann. How to get it. diagrammatic reasoning as a tool of knowledge development and its pragmatic dimension. Foundations of Science, 9(3):285–305, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1023/B:FODA.0000042844.22765.55.

Roberts:2004:RCE

[233] Lawrence D. Roberts. The relation of children’s early word acquisition to abduction. Foundations of Science, 9(3):307–320, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1023/B:FODA.
0000042845.38424.a3.

Myrstad:2004:UCA

[234] Johan Arnt Myrstad. The use of converse abduction in Kepler. Foundations of Science, 9(3):321–338, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1023/B:FODA.0000042846.93506.b8.

Aliseda:2004:LSD

[235] Atocha Aliseda. Logics in scientific discovery. Foundations of Science, 9 (3):339–363, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1023/B:FODA.0000042847.62285.81.

Anonymous:2004:BNc

[236] Anonymous. Biographical notes. Foundations of Science, 9(3):365–368, September 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1023/B:
FODA.0000043026.38913.ac.

Heller:2004:ASD

[237] Michael Heller. Algebraic self-duality as the “ultimate explanation”. Foundations of Science, 9(4):369–385, December 2004. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-005-4825-0.

DHooghe:2004:QMC

[238] Bart D’Hooghe and Jaroslaw Pykacz. Quantum mechanics and computation. Foundations of Science, 9(4):387–404, December 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-005-4827-y.

Micko:2004:IEC

[239] Hans Christoph Micko. On the impossibility of empirical controls of scientific theories — from the point of view of a psychologist. Foundations of Science, 9(4):405–413, December 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-3062-2.

Appelbaum:2004:TCE

[240] Irene Appelbaum. Two conceptions of the emergence of phonemic structure. Foundations of Science, 9(4):415–435, December 2004. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-004-3064-0.

Magnani:2005:P

[241] Lorenzo Magnani and Nancy J. Nersessian. Preface. Foundations of Science, 10(1):1–6, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-005-3002-9.

Magnani:2005:PII

[242] Lorenzo Magnani and Riccardo Dossena. Perceiving the infinite and the infinitesimal world: Unveiling and optical diagrams in mathematics. Foundations of Science, 10(1):7–23, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-005-3003-8.

deLavalette:2005:TBM

[243] Brigitte Cambon de Lavalette, Charles Tijus, Christine Leproux, and Olivier Bauer. Taxonomy based models for reasoning: Making inferences from electronic road sign information. Foundations of Science, 10(1):25–45, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-005-3004-7.

DeMey:2005:RP

[244] Tim De Mey. Remodeling the past. Foundations of Science, 10(1):47–66, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
005-3005-6.

Portides:2005:TSM

[245] Demetris P. Portides. A theory of scientific model construction: The conceptual process of abstraction and concretisation. Foundations of Science, 10(1):67–88, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-005-3006-5.

Trafton:2005:CIE

[246] J. Gregory Trafton, Susan B. Trickett, and Farilee E. Mintz. Connecting internal and external representations: Spatial transformations of scientific visualizations. Foundations of Science, 10(1):89–106, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-005-3007-4.

Magnani:2005:MAC

[247] Lorenzo Magnani and Matteo Piazza. Morphodynamical abduction. causation by attractors dynamics of explanatory hypotheses in science. Foundations of Science, 10(1):107–132, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-005-3008-3.

Davies:2005:VMA

[248] Jim Davies, Nancy J. Nersessian, and Ashok K. Goel. Visual models in analogical problem solving. Foundations of Science, 10(1):133–152, March 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
005-3009-2.

Burchard:2005:SLN

[249] Hermann G. W. Burchard. Symbolic languages and natural structures a mathematician’s account of empiricism. Foundations of Science, 10 (2):153–245, June 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-3068-9.

Anonymous:2005:BNa

[250] Anonymous. Biographical notes. Foundations of Science, 10(2):247–248, June 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
005-5259-4.

Waaldijk:2005:FCM

[251] Frank Waaldijk. On the foundations of constructive mathematics — especially in relation to the theory of continuous functions. Foundations of Science, 10(3):249–324, September 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-004-3065-z.

Mosselmans:2005:TVH

[252] Bert Mosselmans. Time and value in the history of political economy. Foundations of Science, 10(3):325–345, September 2005. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-004-3069-8.

Anonymous:2005:BNb

[253] Anonymous. Biographical notes. Foundations of Science, 10(3): 347, September 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-3071-1.

Weber:2005:P

[254] Erik Weber and Petri Ylikoski. Preface. Foundations of Science, 10(4): 349–351, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5245-2.

Bird:2005:DCL

[255] Alexander Bird. The dispositionalist conception of laws. Foundations of Science, 10(4):353–370, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-5259-9.

Persson:2005:TM

[256] Johannes Persson. Tropes as mechanisms. Foundations of Science, 10 (4):371–393, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5247-0.

Ylikoski:2005:TDR

[257] Petri Ylikoski. The third dogma revisited. Foundations of Science, 10 (4):395–419, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5249-y.

Schweder:2005:DUT

[258] Rebecca Schweder. A defense of a unificationist theory of explanation. Foundations of Science, 10(4):421–435, December 2005. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-004-5250-5.

Weber:2005:FCE

[259] Erik Weber, Jeroen Van Bouwel, and Robrecht Vanderbeeken. Forms of causal explanation. Foundations of Science, 10(4):437–454, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
005-5357-3.

Anonymous:2005:BNc

[260] Anonymous. Biographical notes. Foundations of Science, 10(4):455– 456, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5252-3.

Anonymous:2005:CV

[261] Anonymous. Contents of volume 10. Foundations of Science, 10(4): 457–458, December 2005. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-005-5375-1.

VanKerkhove:2006:ISI

[262] Bart Van Kerkhove, Jean Paul Van Bendegem, and Sal Restivo. Introduction to the special issue entitled ‘Mathematics: What Does it All Mean?’. Foundations of Science, 11(1–2):1–3, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-004-5910-5.

VanKerkhove:2006:MNO

[263] Bart Van Kerkhove. Mathematical naturalism: Origins, guises, and prospects. Foundations of Science, 11(1–2):5–39, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-004-5911-4.

Oliveri:2006:MQE

[264] Gianluigi Oliveri. Mathematics as a quasi-empirical science. Foundations of Science, 11(1–2):41–79, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-5912-3.

VanMoer:2006:IFS

[265] Ard Van Moer. The intentionality of formal systems. Foundations of Science, 11(1–2):81–119, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-5914-1.

Otte:2006:PAC

[266] Michael Otte. Proof-analysis and continuity. Foundations of Science, 11 (1–2):121–155, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5915-0.

DeCruz:2006:TDA

[267] Helen De Cruz. Towards a Darwinian approach to mathematics. Foundations of Science, 11(1–2):157–196, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-004-5916-z.

Restivo:2006:WMM

[268] Sal Restivo and Wenda K. Bauchspies. The will to mathematics: Minds, morals, and numbers. Foundations of Science, 11(1–2):197–215, March 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
004-5917-y.

Russo:2006:SVF

[269] Federica Russo. Salmon and Van Fraassen on the existence of unobservable entities: a matter of interpretation of probability. Foundations of Science, 11(3):221–247, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-6247-9.

Christiaens:2006:BOO

[270] Wim Christiaens. Basic ontology and the ontology of the phenomenological life world: a proposal. Foundations of Science, 11(3):249– 274, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5909-y.

Hansson:2006:FF

[271] Sven Ove Hansson. Falsificationism falsified. Foundations of Science, 11 (3):275–286, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-5922-1.

Cirkovic:2006:QSP

[272] Milan M. Ćirković. Is quantum suicide painless? On an apparent violation of the principal principle. Foundations of Science, 11(3):287– 296, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-004-6249-7.

Locker:2006:STC

[273] Markus Ekkehard Locker. Systems theory and the conundrum of Ens: Thoughts and aphorisms. Foundations of Science, 11(3):297– 317, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-005-3193-0.

Anonymous:2006:BNa

[274] Anonymous. Biographical notes. Foundations of Science, 11(3):319– 321, September 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-8431-6.

Badino:2006:FRE

[275] Massimiliano Badino. The foundational role of ergodic theory. Foundations of Science, 11(4):323–347, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-005-6283-0.

Parker:2006:UPC

[276] W. S. Parker. Understanding pluralism in climate modeling. Foundations of Science, 11(4):349–368, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-005-3196-x.

Mcarthur:2006:APS

[277] Dan Mcarthur. The anti-philosophical stance, the realism question and scientific practice. Foundations of Science, 11(4):369–397, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
005-3198-8.

Broekaert:2006:GLP

[278] Jan Broekaert, Diederik Aerts, and Bart D’Hooghe. The generalised liar paradox: a quantum model and interpretation. Foundations of Science, 11(4):399–418, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-004-6248-8.

Ducheyne:2006:AUG

[279] Steffen Ducheyne. The argument(s) for universal gravitation. Foundations of Science, 11(4):419–447, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-005-3189-9.

Anonymous:2006:BNb

[280] Anonymous. Biographical notes. Foundations of Science, 11(4):449– 451, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-8459-7.

Anonymous:2006:CV

[281] Anonymous. Contents of volume 11. Foundations of Science, 11(4): 453–454, December 2006. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-8463-y.

Neuman:2007:WTD

[282] Yair Neuman and Itzhak Orion. Why are three-dimensional organisms composed of two-dimensional layers? Foundations of Science, 12(1):1–7, March 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
006-9104-1.

Palmquist:2007:EEG

[283] Stephen Palmquist. Emergence, evolution, and the geometry of logic: Causal leaps and the myth of historical development. Foundations of Science, 12(1):9–37, March 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-0004-1.

Grujic:2007:SEQ

[284] Petar V. Grujić. Some epistemic questions of cosmology. Foundations of Science, 12(1):39–83, March 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-006-0005-0.

Bigaj:2007:CNL

[285] Tomasz Bigaj. Counterfactuals and non-locality of quantum mechanics: The Bedford–Stapp version of the GHZ Theorem. Foundations of Science, 12(1):85–108, March 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-006-0006-z.

Uzan:2007:ATM

[286] Pierre Uzan. The arrow of time and meaning. Foundations of Science, 12(2):109–137, June 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-0007-y.

Hansson:2007:PRS

[287] Sven Ove Hansson. Praxis relevance in science. Foundations of Science, 12(2):139–154, June 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-0003-2.

English:2007:EPC

[288] L. Q. English. On the ‘emptiness’ of particles in condensed-matter physics. Foundations of Science, 12(2):155–171, June 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-006-0001-4.

Kosso:2007:SU

[289] Peter Kosso. Scientific understanding. Foundations of Science, 12 (2):173–188, June 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-006-0002-3.

Bosch:2007:PPR

[290] Reinoud Bosch. Pragmatism and the practical relevance of truth. Foundations of Science, 12(3):189–201, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-006-9101-4.

Klein:2007:ACB

[291] Étienne Klein. About the confusion between the course of time and the arrow of time. Foundations of Science, 12(3):203–221, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-006-9103-2.

Vihalemm:2007:PCI

[292] Rein Vihalemm. Philosophy of chemistry and the image of science. Foundations of Science, 12(3):223–234, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-006-9105-0.

Pratten:2007:SOT

[293] Stephen Pratten. The scope of ontological theorising. Foundations of Science, 12(3):235–256, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-007-9106-7.

Hansson:2007:VPA

[294] Sven Ove Hansson. Values in pure and applied science. Foundations of Science, 12(3):257–268, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-007-9107-6.

Ezell:2007:PIT

[295] Barry Charles Ezell and Kenneth G. Crowther. Philosophical issues and their implications for the systems architect. Foundations of Science, 12 (3):269–276, September 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-007-9108-5.

Meschini:2007:PSP

[296] Diego Meschini. Planck-scale physics: Facts and beliefs. Foundations of Science, 12(4):277–294, December 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-006-9102-3.

MacKinnon:2007:SOQ

[297] Edward MacKinnon. Schwinger and the ontology of quantum field theory. Foundations of Science, 12(4):295–323, December 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-007-9109-4.

Bouchard:2007:FCO

[298] Yves Bouchard. The foundationalism–coherentism opposition revisited: The case for complementarism. Foundations of Science, 12(4): 325–336, December 2007. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-007-9110-y.

Longo:2007:DMC

[299] Giuseppe Longo and Pierre-Emmanuel Tendero. The differential method and the causal incompleteness of programming theory in molecular biology. Foundations of Science, 12(4):337–366, December 2007. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-007-9111-x.

Magnani:2008:Pa

[300] Lorenzo Magnani. Preface. Foundations of Science, 13(1):1–4, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
007-9112-9.

Addis:2008:ALT

[301] Tom Addis, Jan Townsend Addis, Dave Billinge, David Gooding, and Bart-Floris Visscher. The abductive loop: Tracking irrational sets. Foundations of Science, 13(1):5–16, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-007-9115-6.

Gooding:2008:MEM

[302] D. C. Gooding and T. R. Addis. Modelling experiments as mediating models. Foundations of Science, 13(1):17–35, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-007-9114-7.

Addis:2008:SMA

[303] T. R. Addis and D. C. Gooding. Simulation methods for an abductive system in science. Foundations of Science, 13(1):37–52, March 2008.

CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-007-9113-8.

Billinge:2008:SAM

[304] Dave Billinge and Tom Addis. Seeking allies: Modelling how listeners choose their musical friends. Foundations of Science, 13(1):53–66, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
007-9120-9.

Vadera:2008:UWF

[305] Sunil Vadera, Andres Rodriguez, Enrique Succar, and Jia Wu. Using Wittgenstein’s family resemblance principle to learn exemplars. Foundations of Science, 13(1):67–74, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-007-9119-2.

Arrighi:2008:ARI

[306] Claudia Arrighi and Roberta Ferrario. Abductive reasoning, interpretation and collaborative processes. Foundations of Science, 13(1):75–87, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
007-9117-4.

Gorman:2008:TZM

[307] Michael E. Gorman. Trading zones, moral imagination and socially sensitive computing. Foundations of Science, 13(1):89–97, March 2008. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-007-9118-3.

Magnani:2008:DME

[308] Lorenzo Magnani and Emanuele Bardone. Distributed morality: Externalizing ethical knowledge in technological artifacts. Foundations of Science, 13(1):99–108, March 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-007-9116-5.

Magnani:2008:Pb

[309] Lorenzo Magnani. Preface. Foundations of Science, 13(2):109–111, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
008-9128-9.

Amigoni:2008:MAM

[310] Francesco Amigoni and Viola Schiaffonati. A multiagent approach to modelling complex phenomena. Foundations of Science, 13(2):113–125, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
008-9121-3.

Boon:2008:DME

[311] Mieke Boon. Diagrammatic models in the engineering sciences. Foundations of Science, 13(2):127–142, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-008-9122-2.

Finkeissen:2008:SEP

[312] Ekkehard Finkeissen. Steps from erratic projects towards structured programs in research. Foundations of Science, 13(2):143–148, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9123-1.

Heeffer:2008:ESA

[313] Albrecht Heeffer. The emergence of symbolic algebra as a shift in predominant models. Foundations of Science, 13(2):149–161, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9124-0.

Steinberg:2008:IEP

[314] Melvin S. Steinberg. Inventing electric potential. Foundations of Science, 13(2):163–175, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9126-y.

Sung:2008:EAR

[315] John J. Sung. Embodied anomaly resolution in molecular genetics: a case study of RNAi. Foundations of Science, 13(2):177–193, June 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9127-x.

Mosselmans:2008:ALQ

[316] Bert Mosselmans. Aristotle’s logic and the quest for the quantification of the predicate. Foundations of Science, 13(3–4):195–198, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9141-z.

Tighe:2008:GCA

[317] Joseph A. Tighe. The God concept: Aristotle and the philosophical tradition. Foundations of Science, 13(3–4):217–228, November 2008. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9139-6.

deGaray:2008:AD

[318] Jesús de Garay. Aristotelism of difference. Foundations of Science, 13(3– 4):229–237, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9140-0.

Pagan:2008:CMS

[319] Nicholas O. Pagan. Configuring the moral self: Aristotle and Dewey. Foundations of Science, 13(3–4):239–250, November 2008. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-008-9137-8.

Scott-Smith:2008:AUP

[320] Giles Scott-Smith. Aristotle, US public diplomacy, and the Cold War: The work of Carnes Lord. Foundations of Science, 13(3–4): 251–264, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9138-7.

Pack:2008:ADR

[321] Spencer J. Pack. Aristotle’s difficult relationship with modern economic theory. Foundations of Science, 13(3–4):265–280, November 2008. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9135-x.

Burke:2008:AAP

[322] M. Burke. Advertising Aristotle: a preliminary investigation into the contemporary relevance of Aristotle’s art of rhetoric. Foundations of Science, 13(3–4):295–305, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-008-9133-z.

DeCuypere:2008:MRA

[323] Ludovic De Cuypere and Klaas Willems. Meaning and reference in Aristotle’s concept of the linguistic sign. Foundations of Science, 13(3–4): 307–324, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9134-y.

Krijanskaia:2008:NAM

[324] Dasha Krijanskaia. A non–Aristotelian model: Time as space and landscape in postmodern theatre. Foundations of Science, 13(3–4): 337–345, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9132-0.

Crespo:2008:EAA

[325] Ricardo Crespo. ‘the economic’ according to Aristotle: Ethical, political and epistemological implications. Foundations of Science, 13(3– 4):??, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9136-9.

Hennig:2008:MZ

[326] Boris Hennig. Matter in Z3. Foundations of Science, 13(3–4):??, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
008-9142-y.

Tindemans:2008:PPA

[327] Klaas Tindemans. The politics of the poetics: Aristotle and drama theory in 17th Century France. Foundations of Science, 13(3–4):??, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
008-9131-1.

Verelst:2008:WO

[328] Karin Verelst. On what ontology is and not-is. Foundations of Science, 13(3–4):??, November 2008. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9130-2.

Azzouni:2009:WDI

[329] Jody Azzouni. Why do informal proofs conform to formal norms? Foundations of Science, 14(1–2):9–26, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-008-9144-9.

Coleman:2009:SLP

[330] Edwin Coleman. The surveyability of long proofs. Foundations of Science, 14(1–2):27–43, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-008-9145-8.

VanBendegem:2009:MAC

[331] Jean Paul Van Bendegem and Bart Van Kerkhove. Mathematical arguments in context. Foundations of Science, 14(1–2):45–57, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9146-7.

Kulpa:2009:MPD

[332] Zenon Kulpa. Main problems of diagrammatic reasoning. Part I: The generalization problem. Foundations of Science, 14(1–2):75–96, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9148-5.

Inglis:2009:PVA

[333] Matthew Inglis and Juan Pablo Mejía-Ramos. On the persuasiveness of visual arguments in mathematics. Foundations of Science, 14(1– 2):97–110, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9149-4.

Pease:2009:BGB

[334] Alison Pease, Alan Smaill, Simon Colton, and John Lee. Bridging the gap between argumentation theory and the philosophy of mathematics. Foundations of Science, 14(1–2):111–135, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-008-9150-y.

Dove:2009:TTM

[335] Ian J. Dove. Towards a theory of mathematical argument. Foundations of Science, 14(1–2):137–152, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-008-9156-5.

Aberdein:2009:MA

[336] Andrew Aberdein. Mathematics and argumentation. Foundations of Science, 14(1–2):??, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9158-3.

Sherry:2009:RDM

[337] David Sherry. The role of diagrams in mathematical arguments. Foundations of Science, 14(1–2):??, March 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-008-9147-6.

Gauthier:2009:CCT

[338] Yvon Gauthier. The construction of chaos theory. Foundations of Science, 14(3):153–165, August 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-008-9129-8.

Knyazeva:2009:NCC

[339] Helena Knyazeva. Nonlinear cobweb of cognition. Foundations of Science, 14(3):167–179, August 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-008-9143-x.

Boniolo:2009:STE

[340] Giovanni Boniolo, Rossella Faraldo, and Antonio Saggion. On spatial and temporal ex mensura boundaries. Foundations of Science, 14(3):181–193, August 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
008-9151-x.

Laland:2009:CBP

[341] Kevin N. Laland, John Odling-Smee, Marcus W. Feldman, and Jeremy Kendal. Conceptual barriers to progress within evolutionary biology. Foundations of Science, 14(3):195–216, August 2009. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-008-9153-8.

Dennis:2009:FLN

[342] L. Dennis, R. W. Gray, L. H. Kauffman, J. Brender McNair, and N. J. Woolf. A framework linking non-living and living systems: Classification of persistence, survival and evolution transitions. Foundations of Science, 14(3):217–238, August 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-008-9154-7.

Mets:2009:CRA

[343] Ave Mets and Piret Kuusk. The constructive realist account of science and its application to Ilya Prigogine’s conception of laws of nature. Foundations of Science, 14(3):239–248, August 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-008-9157-4.

Darvas:2009:CCP

[344] György Darvas. Can the causal paradoxes of QM be explained in the framework of QED? Foundations of Science, 14(4):273–280, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
009-9159-x.

Isenman:2009:DES

[345] Lois Isenman. Digestive enzyme secretion, intuition, and the history of science: Part II. Foundations of Science, 14(4):331–349, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-009-9163-1.

Obradovic:2009:HFM

[346] Stojan Obradović and Slobodan Ninković. The heuristic function of mathematics in physics and astronomy. Foundations of Science, 14(4): 351–360, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-009-9162-2.

Aerts:2009:QPC

[347] Diederik Aerts. Quantum particles as conceptual entities: a possible explanatory framework for quantum theory. Foundations of Science, 14 (4):361–411, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-009-9166-y.

Abrantes:2009:GHI

[348] Paulo Abrantes and Charbel Niño El-Hani. Gould, Hull, and the individuation of scientific theories. Foundations of Science, 14(4):??, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
009-9161-3.

Farrell:2009:EES

[349] R. P. Farrell and C. A. Hooker. Error, error-statistics and self-directed anticipative learning. Foundations of Science, 14(4):??, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-008-9155-6.

Isenman:2009:TYG

[350] Lois Isenman. Trusting your gut, among other things: Digestive enzyme secretion, intuition, and the history of science. Foundations of Science, 14(4):??, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-009-9164-0.

Rowbottom:2009:MBP

[351] Darrell P. Rowbottom. Models in biology and physics: What’s the difference? Foundations of Science, 14(4):??, November 2009. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-009-9160-4.

Uzan:2010:IBU

[352] Pierre Uzan. Informational branching universe. Foundations of Science, 15(1):1–28, February 2010. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-009-9165-z.

Hennig:2010:MMR

[353] Christian Hennig. Mathematical models and reality: a constructivist perspective. Foundations of Science, 15(1):29–48, February 2010. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-009-9167-x.

Luk:2010:USS

[354] Robert W. P. Luk. Understanding scientific study via process modeling. Foundations of Science, 15(1):49–78, February 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-009-9168-9.

Krause:2010:LAQ

[355] Décio Krause. Logical aspects of quantum (non-)individuality. Foundations of Science, 15(1):79–94, February 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-009-9169-8.

Vidal:2010:ISI

[356] Clément Vidal. Introduction to the special issue on the evolution and development of the Universe. Foundations of Science, 15(2):95–99, June 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

010-9176-9;   http://link.springer.com/content/pdf/10.1007/
s10699-010-9176-9.pdf.

Nottale:2010:SRF

[357] Laurent Nottale. Scale relativity and fractal space–time: Theory and applications. Foundations of Science, 15(2):101–152, June 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9170-2.

Greben:2010:REC

[358] Jan M. Greben. The role of energy conservation and vacuum energy in the evolution of the Universe. Foundations of Science, 15(2):153–176, June 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9172-0.

Durt:2010:AQD

[359] Thomas Durt. Anthropomorphic quantum Darwinism as an explanation for classicality. Foundations of Science, 15(2):177–197, June 2010. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9173-z.

Lori:2010:AQD

[360] Nicolás F. Lori and Alex H. Blin. Application of quantum Darwinism to cosmic inflation: an example of the limits imposed in Aristotelian logic by information-based approach to Gödel’s incompleteness. Foundations of Science, 15(2):199–211, June 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9174-y.

Kosso:2010:YIM

[361] Peter Kosso. And yet it moves: The observability of the rotation of the Earth. Foundations of Science, 15(3):213–225, August 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9175-x.

Boudry:2010:HAI

[362] Maarten Boudry, Stefaan Blancke, and Johan Braeckman. How not to attack intelligent design Creationism: Philosophical misconceptions about methodological naturalism. Foundations of Science, 15(3):227–244, August 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9178-7.

opAkkerhuis:2010:THD

[363] Gerard A. J. M. Jagers op Akkerhuis. Towards a hierarchical definition of life, the organism, and death. Foundations of Science, 15(3):245–262, August 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9177-8.

Ekstig:2010:CES

[364] Börje Ekstig. Complexity and evolution: a study of the growth of complexity in organic and cultural evolution. Foundations of Science, 15 (3):263–278, August 2010. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9179-6.

Chaline:2010:DSE

[365] Jean Chaline. Does species evolution follow scale laws? First applications of the scale relativity theory to fossil and living-beings. Foundations of Science, 15(3):279–302, August 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9180-0.

Uskokovic:2010:MCM

[366] Vuk Uskoković. Major challenges for the modern chemistry in particular and science in general. Foundations of Science, 15(4):303–344, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9185-8.

Heylighen:2010:SOT

[367] Francis Heylighen. The self-organization of time and causality: Steps towards understanding the ultimate origin. Foundations of Science, 15 (4):345–356, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9171-1.

Salthe:2010:DEU

[368] Stanley N. Salthe. Development (and evolution) of the Universe. Foundations of Science, 15(4):357–367, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9181-z.

Crane:2010:PIQ

[369] Louis Crane. Possible implications of the quantum theory of gravity: an introduction to the Meduso–anthropic principle. Foundations of Science, 15(4):369–373, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9182-y.

Vidal:2010:CBA

[370] Clément Vidal. Computational and biological analogies for understanding fine-tuned parameters in physics. Foundations of Science, 15(4): 375–393, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9183-x.

Stewart:2010:MLD

[371] John E. Stewart. The meaning of life in a developing Universe. Foundations of Science, 15(4):395–409, November 2010. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9184-9.

Napoletani:2011:AST

[372] D. Napoletani, M. Panza, and D. C. Struppa. Agnostic science. towards a philosophy of data analysis. Foundations of Science, 16(1):1–20, February 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9186-7.

Park:2011:COB

[373] Seungbae Park. Coherence of our best scientific theories. Foundations of Science, 16(1):21–30, February 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9198-3.

Perovic:2011:GAR

[374] Slobodan Perovic and Paul-Antoine Miquel. On gene’s action and reciprocal causation. Foundations of Science, 16(1):31–46, February 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9199-2.

Tohme:2011:DRP

[375] Fernando Tohmé, Claudio Delrieux, and Otávio Bueno. Defeasible reasoning + partial models: a formal framework for the methodology of research programs. Foundations of Science, 16(1):47–65, February 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9200-0.

Burchard:2011:RCA

[376] Hermann G. W. Burchard. The role of conscious attention in perception. Foundations of Science, 16(1):67–99, February 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9201-z.

Goeminne:2011:OBI

[377] Gert Goeminne and Erik Paredis. Opening up the in-between: Ihde’s postphenomenology and beyond. Foundations of Science, 16(2–3): 101–107, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9229-8;   http://link.springer.com/content/
pdf/10.1007/s10699-011-9229-8.pdf.

Ihde:2011:SBE

[378] Don Ihde. Stretching the in-between: Embodiment and beyond. Foundations of Science, 16(2–3):109–118, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9187-6.

DePreester:2011:TBI

[379] Helena De Preester. Technology and the body: the (im)possibilities of re-embodiment. Foundations of Science, 16(2–3):119–137, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9188-5.

VanDenEede:2011:BUT

[380] Yoni Van Den Eede. In between us: On the transparency and opacity of technological mediation. Foundations of Science, 16(2–3):139–159, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9190-y.

Francois:2011:BSP

[381] Karen François. In-between science and politics. Foundations of Science, 16(2–3):161–171, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9195-6.

Goeminne:2011:PPS

[382] Gert Goeminne. Postphenomenology and the politics of sustainable technology. Foundations of Science, 16(2–3):173–194, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9196-5.

Paredis:2011:STN

[383] Erik Paredis. Sustainability transitions and the nature of technology. Foundations of Science, 16(2–3):195–225, May 2011. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9197-4.

Scharff:2011:DEB

[384] Robert C. Scharff. Displacing epistemology: Being in the midst of technoscientific practice. Foundations of Science, 16(2–3):227–243, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9189-4.

Vanderbeeken:2011:SB

[385] Robrecht Vanderbeeken. The screen as an in-between. Foundations of Science, 16(2–3):245–257, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9191-x.

Kamphof:2011:WSN

[386] Ike Kamphof. Webcams to save nature: Online space as affective and ethical space. Foundations of Science, 16(2–3):259–274, May 2011. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9194-7.

Vanhoutte:2011:PPN

[387] Kurt Vanhoutte and Nele Wynants. Performing phenomenology: Negotiating presence in intermedial theatre. Foundations of Science, 16 (2–3):275–284, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9193-8.

Riis:2011:DBW

[388] Søren Riis. Dwelling in-between walls: The architectural surround. Foundations of Science, 16(2–3):285–301, May 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9192-9.

Auffray:2011:SRE

[389] Charles Auffray and Denis Noble. Scale relativity: an extended paradigm for physics and biology? Foundations of Science, 16(4):303–305, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9203-x.

Nottale:2011:MIS

[390] Laurent Nottale. Multiscale integration in scale relativity theory. Foundations of Science, 16(4):307–309, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-010-9204-9.

Lori:2011:DIP

[391] Nicolás F. Lori. On definitions of information in physics. Foundations of Science, 16(4):311–314, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9205-8.

Durt:2011:CDI

[392] Thomas Durt. Competing definitions of information versus entropy in physics. Foundations of Science, 16(4):315–318, November 2011. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9206-7.

vanStraalen:2011:ICJ

[393] Nico M. van Straalen. The issue of “closure” in Jagers op Akkerhuis’s operator theory. Foundations of Science, 16(4):319–321, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-010-9207-6.

Hengeveld:2011:DLO

[394] Rob Hengeveld. Definitions of life are not only unnecessary, but they can do harm to understanding. Foundations of Science, 16(4):323– 325, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9208-5.

opAkkerhuis:2011:EOL

[395] Gerard Jagers op Akkerhuis. Explaining the origin of life is not enough for a definition of life. Foundations of Science, 16(4):327–329, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9209-4.

Longo:2011:SSB

[396] Giuseppe Longo. Symmetries and symmetry-breakings: The fabric of physical interactions and the flow of time. Foundations of Science, 16 (4):331–333, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9210-y.

Heylighen:2011:SPR

[397] Francis Heylighen. Symmetry, potentiality and reversibility. Foundations of Science, 16(4):335–336, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9211-x.

Riss:2011:OPM

[398] Uwe V. Riss. Objects and processes in mathematical practice. Foundations of Science, 16(4):337–351, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9202-y.

Iftime:2011:NMM

[399] Mihaela D. Iftime. A nonlinear method for measuring the effects of environmental variations. Foundations of Science, 16(4):353–361, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
011-9225-z.

Krause:2011:AMS

[400] Décio Krause, Jonas R. B. Arenhart, and Fernando T. F. Moraes. Axiomatization and models of scientific theories. Foundations of Science, 16 (4):363–382, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9226-y.

Vazquez:2011:MPV

[401] Margarita Vázquez and Manuel Liz. Models as points of view: The case of system dynamics. Foundations of Science, 16(4):383–391, November 2011. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
011-9224-0.

deBianchi:2011:EPI

[402] Massimiliano Sassoli de Bianchi. Ephemeral properties and the illusion of microscopic particles. Foundations of Science, 16(4):??, November 2011.

CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9227-x.

Fairlamb:2012:MCS

[403] Horace Fairlamb. Must complex systems theory be materialistic? Foundations of Science, 17(1):1–3, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9212-9.

VandeVijver:2012:FW

[404] Gertrudis Van de Vijver. Friends of wisdom? Foundations of Science, 17(1):5–7, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9213-8.

Salthe:2012:MRC

[405] Stanley Salthe. Materialism: Replies to comments from readers. Foundations of Science, 17(1):9–11, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9214-7.

Vidal:2012:TPB

[406] Clément Vidal. Two purposes of black hole production. Foundations of Science, 17(1):13–15, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-010-9215-6.

Crane:2012:PE

[407] Louis Crane. From philosophy to engineering. Foundations of Science, 17(1):17–19, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-010-9216-5.

Greben:2012:NIC

[408] Jan M. Greben. On the nature of initial conditions and fundamental parameters in physics and cosmology. Foundations of Science, 17(1):21–23, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9217-4.

Vaas:2012:CAS

[409] Rüdiger Vaas. Cosmological artificial selection: Creation out of something? Foundations of Science, 17(1):25–28, March 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-010-9218-3.

Vidal:2012:FTQ

[410] Clément Vidal. Fine-tuning, quantum mechanics and cosmological artificial selection. Foundations of Science, 17(1):29–38, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-010-9219-2.

Rottiers:2012:PML

[411] Franc Rottiers. Participating in the meaning of life, a contributor’s critique. Foundations of Science, 17(1):39–41, March 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-010-9220-9.

Vidal:2012:ASS

[412] Clément Vidal. Analysis of some speculations concerning the far future of intelligent civilizations. Foundations of Science, 17(1):43–46, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
010-9221-8.

Stewart:2012:FLW

[413] John E. Stewart. The future of life and what it means for humanity. Foundations of Science, 17(1):47–50, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-010-9222-7.

Katz:2012:BCN

[414] Karin Usadi Katz and Mikhail G. Katz. A Burgessian critique of nominalistic tendencies in contemporary mathematics and its historiography. Foundations of Science, 17(1):51–89, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9223-1.

Kertesz:2012:GSS

[415] András Kertész. The ‘Galilean style in science’ and the inconsistency of linguistic theorising. Foundations of Science, 17(1):91–108, March 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9234-y.

Katz:2012:SNR

[416] Karin Usadi Katz and Mikhail G. Katz. Stevin numbers and reality. Foundations of Science, 17(2):109–123, June 2012. CODEN FOSCFI.

ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9228-9.

Posner:2012:SCEa

[417] Zeev Posner. Spatio-cultural evolution as information dynamics: Part I. Foundations of Science, 17(2):125–162, June 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9230-2.

Posner:2012:SCEb

[418] Zeev Posner. Spatio-cultural evolution as information dynamics — Part II. Foundations of Science, 17(2):163–203, June 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9231-1.

Bigaj:2012:UDQ

[419] Tomasz Bigaj. Ungrounded dispositions in quantum mechanics. Foundations of Science, 17(3):205–221, August 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-011-9232-0.

deBianchi:2012:PTA

[420] Massimiliano Sassoli de Bianchi. From permanence to total availability: a quantum conceptual upgrade. Foundations of Science, 17(3):223–244, August 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
011-9233-z.

Borovik:2012:WGY

[421] Alexandre Borovik and Mikhail G. Katz. Who gave you the Cauchy– Weierstrass tale? The dual history of rigorous calculus. Foundations of Science, 17(3):245–276, August 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9235-x.

Riofrio:2012:SMM

[422] Walter Riofrio. Studies on molecular mechanisms of prebiotic systems. Foundations of Science, 17(3):277–289, August 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9236-9.

Romero:2012:PR

[423] Gustavo E. Romero. Parmenides reloaded. Foundations of Science, 17 (3):291–299, August 2012. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9272-5.

Braating:2012:AFC

[424] Kajsa Bråting. Ambiguities of fundamental concepts in mathematical analysis during the mid-nineteenth century. Foundations of Science, 17 (4):301–320, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9274-3.

Park:2012:AEA

[425] Woosuk Park. Abduction and estimation in animals. Foundations of Science, 17(4):321–337, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9275-2.

Arenhart:2012:OFS

[426] Jonas R. B. Arenhart. Ontological frameworks for scientific theories. Foundations of Science, 17(4):339–356, November 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9288-5.

Olesen:2012:SOP

[427] Finn Olesen. Scientific objectivity and postphenomenological perception. Foundations of Science, 17(4):357–362, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9241-z.

BergOlsenFriis:2012:PEB

[428] Jan Kyrre Berg Olsen Friis. Perception: Embodiment and beyond. Foundations of Science, 17(4):363–367, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9242-y.

Ihde:2012:CRS

[429] Don Ihde. ‘Cartesianism’ redux or situated knowledges. Foundations of Science, 17(4):369–372, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9243-x.

Ihde:2012:PRE

[430] Don Ihde. Postphenomenological re-embodiment. Foundations of Science, 17(4):373–377, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9244-9.

Lenay:2012:STC

[431] Charles Lenay. Separability and technical constitution. Foundations of Science, 17(4):379–384, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9245-8.

DePreester:2012:TMN

[432] Helena De Preester. Technology and the myth of ‘natural man’. Foundations of Science, 17(4):385–390, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-011-9246-7.

Verbeek:2012:EMT

[433] Peter-Paul Verbeek. Expanding mediation theory. Foundations of Science, 17(4):391–395, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9253-8.

Levinson:2012:SRR

[434] Paul Levinson. Skype and the reality of remedial media. Foundations of Science, 17(4):397–399, November 2012. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9254-7.

VanDenEede:2012:HCC

[435] Yoni Van Den Eede. Of humans & cyborgs, caterpillars & butterflies. Foundations of Science, 17(4):401–405, November 2012. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9252-9.

Kastner:2013:BWB

[436] Ruth E. Kastner. de Broglie waves as the “bridge of becoming” between quantum theory and relativity. Foundations of Science, 18(1):1–9, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
011-9273-4.

deBianchi:2013:QMM

[437] Massimiliano Sassoli de Bianchi. The δ-quantum machine, the k-model, and the non-ordinary spatiality of quantum entities. Foundations of Science, 18(1):11–41, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9284-1.

Blaszczyk:2013:TMH

[438] Piotr Blaszczyk, Mikhail G. Katz, and David Sherry. Ten misconceptions from the history of analysis and their debunking. Foundations of Science, 18(1):43–74, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9285-8.

Rosenberger:2013:MMP

[439] Robert Rosenberger. Mediating Mars: Perceptual experience and scientific imaging technologies. Foundations of Science, 18(1):75–91, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
012-9286-7.

Cellucci:2013:TBP

[440] Carlo Cellucci. Top-down and bottom-up philosophy of mathematics. Foundations of Science, 18(1):93–106, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9287-6.

Katz:2013:CDD

[441] Mikhail G. Katz and David Tall. A Cauchy–Dirac delta function. Foundations of Science, 18(1):107–123, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-012-9289-4.

Ferrero:2013:FRI

[442] M. Ferrero, V. Gómez Pin, D. Salgado, and J. L. Sánchez-Gómez. A further review of the incompatibility between classical principles and quantum postulates. Foundations of Science, 18(1):125–138, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9290-y.

Romero:2013:CSE

[443] Gustavo E. Romero. From change to spacetime: an eleatic journey. Foundations of Science, 18(1):139–148, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-012-9297-4.

Vanhoutte:2013:LIP

[444] Kurt Vanhoutte. Luddite interventions: on the poetics of catastrophe and the art of criticism. Foundations of Science, 18(1):149–153, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9257-4.

Hoens:2013:WNA

[445] Dominiek Hoens. What is new about new media? Foundations of Science, 18(1):155–158, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9256-5.

Vanderbeeken:2013:VI

[446] Robrecht Vanderbeeken. Virtual invaders. Foundations of Science, 18 (1):159–164, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9255-6.

Wise:2013:PC

[447] J. Macgregor Wise. The politics of care. Foundations of Science, 18 (1):165–168, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9266-3.

vandeVall:2013:PP

[448] Renée van de Vall. Promises of presence. Foundations of Science, 18 (1):169–172, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9265-4.

Kamphof:2013:TAT

[449] Ike Kamphof. Thinking (-animal-technology-human-) touch. Foundations of Science, 18(1):173–178, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-011-9264-5.

Paulus:2013:MDP

[450] Tom Paulus. Movements of discovery: the pragmatics of practice-based research. Foundations of Science, 18(1):179–183, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9263-6.

Merx:2013:DPP

[451] Sigrid Merx. From doing to performing phenomenology: Implications and possibilities. Foundations of Science, 18(1):185–190, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9262-7.

Vanhoutte:2013:DGP

[452] Kurt Vanhoutte and Nele Wynants. Dealing with the ghost: Phantasmagorical apparitions of Bertolt Brecht. Foundations of Science, 18 (1):191–194, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9261-8.

Read:2013:TB

[453] Stephen Read. Technology as in-between. Foundations of Science, 18 (1):195–200, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9260-9.

Benjamin:2013:ATD

[454] Andrew Benjamin. Architecture and technology: a discontinuous relation. Foundations of Science, 18(1):201–204, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9259-2.

Riis:2013:RA

[455] Søren Riis. Reframing architecture. Foundations of Science, 18(1): 205–211, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9258-3.

deBianchi:2013:OE

[456] Massimiliano Sassoli de Bianchi. The observer effect. Foundations of Science, 18(2):213–243, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-012-9298-3.

Marriner:2013:DMI

[457] Nick Marriner and Christophe Morhange. Data mining the intellectual revival of ‘catastrophic’ mother nature. Foundations of Science, 18(2):245–257, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9299-2.

Kanovei:2013:TOC

[458] Vladimir Kanovei, Mikhail G. Katz, and Thomas Mormann. Tools, objects, and chimeras: Connes on the role of hyperreals in mathematics. Foundations of Science, 18(2):259–296, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9316-5.

Romero:2013:ASO

[459] Gustavo E. Romero. Adversus singularitates: The ontology of space– time singularities. Foundations of Science, 18(2):297–306, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9309-4.

Obradovic:2013:EES

[460] Stojan Obradović. Empirical evidence in the structure of physical theories. Foundations of Science, 18(2):307–318, June 2013. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9319-2.

Bair:2013:OCM

[461] Jacques Bair and Valérie Henry. Osculating circle with microscopes within microscopes. Foundations of Science, 18(2):319–325, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9320-9.

deSutter:2013:RRW

[462] Laurent de Sutter. Representation reloaded: Why Socrates should die again. Foundations of Science, 18(2):327–329, June 2013. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9269-0.

Jensen:2013:WIW

[463] Casper Bruun Jensen. What if we were already in the in-between? Further ventures into the ontologies of science and politics. Foundations of Science, 18(2):331–336, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9268-1.

Francois:2013:PL

[464] Karen François. Pro-Latour. Foundations of Science, 18(2):337–342, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
011-9267-2.

Scharff:2013:WTM

[465] Robert C. Scharff. “Who” is a “topical measuring” postphenomenologist and how does one get that way? Foundations of Science, 18(2):343–350, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

011-9251-x. See response [467].

Puech:2013:WPP

[466] Michel Puech. Why not post-political? Foundations of Science, 18 (2):351–353, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9250-y. See response [467].

Goeminne:2013:WAP

[467] Gert Goeminne. Who is afraid of the political? A response to Robert Scharff and Michel Puech. Foundations of Science, 18(2):355–360, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

013-9331-1. See [465, 466].

Vergragt:2013:BPT

[468] Philip J. Vergragt. Beyond politization of technology and sustainability: a plea for visioning. Foundations of Science, 18(2):361–365, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

011-9270-7. See response [470].

Bos:2013:NTS

[469] A. P. Bos. Nurturing technologies for sustainability transitions. Foundations of Science, 18(2):367–372, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-011-9271-6. See response [470].

Paredis:2013:EPT

[470] Erik Paredis. Embracing the political in technology and transition studies: a response to Philip Vergragt and Bram Bos. Foundations of Science, 18(2):373–377, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9308-5. See [469, 468].

Heelan:2013:POQ

[471] Patrick A. Heelan. Phenomenology, ontology, and quantum physics. Foundations of Science, 18(2):379–385, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9247-6.

Ruin:2013:TTP

[472] Hans Ruin. Thinking through the prism of life. Foundations of Science, 18(2):387–392, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9248-5.

Scharff:2013:BPP

[473] Robert C. Scharff. Being post-positivist … or just talking about it? Foundations of Science, 18(2):393–397, June 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-011-9249-4.

VanDyck:2013:WNE

[474] Maarten Van Dyck and Karin Verelst. “Whatever is neither everywhere nor anywhere does not exist”: The concepts of space and time in Newton and Leibniz. Foundations of Science, 18(3):399–402, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9276-1.

Janiak:2013:MNP

[475] Andrew Janiak. Metaphysics and natural philosophy in Descartes and Newton. Foundations of Science, 18(3):403–417, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-011-9277-0.

Slowik:2013:NNP

[476] Edward Slowik. Newton’s neo-Platonic ontology of space. Foundations of Science, 18(3):419–448, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-011-9278-z.

Schliesser:2013:NES

[477] Eric Schliesser. Newtonian emanation, Spinozism, measurement and the Baconian origins of the laws of nature. Foundations of Science, 18 (3):449–466, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9279-y.

Belkind:2013:LNS

[478] Ori Belkind. Leibniz and Newton on space. Foundations of Science, 18 (3):467–497, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9280-5.

Arthur:2013:LTS

[479] Richard T. W. Arthur. Leibniz’s theory of space. Foundations of Science, 18(3):499–528, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-011-9281-4.

Kerszberg:2013:DVD

[480] Pierre Kerszberg. Deduction versus discourse: Newton and the cosmic phenomena. Foundations of Science, 18(3):529–544, August 2013. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-011-9283-2.

deSojo:2013:TSD

[481] Aurea Anguera de Sojo, Juan Ares, Juan A. Lara, David Lizcano, María A. Martínez, and Juan Pazos. Turing and the serendipitous discovery of the modern computer. Foundations of Science, 18(3):545–557, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9327-x.

deAndrade:2013:SPP

[482] Elaine Maria Paiva de Andrade, Jean Faber, and Luiz Pinguelli Rosa. A spontaneous physics philosophy on the concept of ether throughout the history of science: Birth, death and revival. Foundations of Science, 18 (3):559–577, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-013-9336-9.

Humphreys:2013:DAM

[483] Paul Humphreys. Data analysis: Models or techniques? Foundations of Science, 18(3):579–581, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-012-9317-4.

Lenhard:2013:CD

[484] Johannes Lenhard. Coal to diamonds. Foundations of Science, 18(3): 583–586, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9318-3. See reply [486].

Napoletani:2013:PRT

[485] Domenico Napoletani, Marco Panza, and Daniele C. Struppa. Processes rather than descriptions? Foundations of Science, 18(3):587–590, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9332-0.

Napoletani:2013:ADS

[486] Domenico Napoletani, Marco Panza, and Daniele C. Struppa. Artificial diamonds are still diamonds. Foundations of Science, 18(3):591–594, August 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9333-z. See [484].

Weber:2013:I

[487] Erik Weber, Dietlinde Wouters, and Joke Meheus. Introduction. Foundations of Science, 18(4):595–597, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9291-x;
 http://link.springer.com/content/pdf/10.1007/s10699-012-9291-
x.pdf.

Demey:2013:CEL

[488] Lorenz Demey. Contemporary epistemic logic and the Lockean thesis. Foundations of Science, 18(4):599–610, November 2013. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9292-9.

Gauderis:2013:MAS

[489] Tjerk Gauderis. Modelling abduction in science by means of a modal adaptive logic. Foundations of Science, 18(4):611–624, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9293-8.

McGinnis:2013:UAP

[490] Nicholas D. McGinnis. The unexpected applicability of paraconsistent logic: a Chomskyan route to dialetheism. Foundations of Science, 18(4): 625–640, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9294-7.

Urbaniak:2013:ISI

[491] Rafal Urbaniak and Frederik Van De Putte. Induction from a single instance: Incomplete frames. Foundations of Science, 18(4):641– 653, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9295-6;   http://link.springer.com/content/
pdf/10.1007/s10699-012-9295-6.pdf.

Verdee:2013:NMS

[492] Peter Verdée. Non-monotonic set theory as a pragmatic foundation of mathematics. Foundations of Science, 18(4):655–680, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9296-5.

Santos:2013:PC

[493] Gil C. Santos. Philosophy and complexity. Foundations of Science, 18 (4):681–686, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9321-8.

Baker:2013:CNN

[494] Alan Baker. Complexity, networks, and non-uniqueness. Foundations of Science, 18(4):687–705, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-012-9300-0.

Bechtel:2013:TDA

[495] William Bechtel and Adele A. Abrahamsen. Thinking dynamically about biological mechanisms: Networks of coupled oscillators. Foundations of Science, 18(4):707–723, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-012-9301-z.

Boogerd:2013:MEM

[496] Fred C. Boogerd, Frank J. Bruggeman, and Robert C. Richardson. Mechanistic explanations and models in molecular systems biology. Foundations of Science, 18(4):725–744, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-012-9302-y.

Edmonds:2013:CCD

[497] Bruce Edmonds. Complexity and context-dependency. Foundations of Science, 18(4):745–755, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-012-9303-x.

Hooker:2013:ICC

[498] Cliff Hooker. On the import of constraints in complex dynamical systems. Foundations of Science, 18(4):757–780, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-012-9304-9.

Gershenson:2013:IIS

[499] Carlos Gershenson. The implications of interactions for science and philosophy. Foundations of Science, 18(4):781–790, November 2013. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9305-8.

Bedau:2013:MME

[500] Mark A. Bedau. Minimal memetics and the evolution of patented technology. Foundations of Science, 18(4):791–807, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-012-9306-7.

Symons:2013:HCM

[501] John Symons and Fabio Boschetti. How computational models predict the behavior of complex systems. Foundations of Science, 18(4): 809–821, November 2013. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9307-6.

VanDyck:2014:SSW

[502] Maarten Van Dyck and Albrecht Heeffer. Script and symbolic writing in mathematics and natural philosophy. Foundations of Science, 19(1):1–10, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
012-9310-y.

Schemmel:2014:MRC

[503] Matthias Schemmel. Medieval representations of change and their early modern application. Foundations of Science, 19(1):11–34, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-012-9312-9.

Hyslop:2014:AC

[504] Scott J. Hyslop. Algebraic collisions. Foundations of Science, 19 (1):35–51, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-012-9313-8.

Leduc:2014:EFS

[505] Christian Leduc. The epistemological functions of symbolization in Leibniz’s universal characteristic. Foundations of Science, 19(1):53–68, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
012-9315-6.

Mertens:2014:MGB

[506] Manuel Mertens. Memory and geometry in Bruno: Some analogies. Foundations of Science, 19(1):69–88, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9314-7.

Heeffer:2014:EJO

[507] Albrecht Heeffer. Epistemic justification and operational symbolism. Foundations of Science, 19(1):89–113, March 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-012-9311-x.

Serrano:2014:CS

[508] J. Ignacio Serrano, M. Dolores del Castillo, and Manuel Carretero. Cognitive? Science? Foundations of Science, 19(2):115–131, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-013-9323-1.

Adams:2014:WCP

[509] Fred Adams. What is a cognitive process? Foundations of Science, 19(2):133–135, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-013-9324-0.

Charro:2014:PVB

[510] Fernando Charro and Juan J. Colomina. Points of view beyond models: Towards a formal approach to points of view as access to the world. Foundations of Science, 19(2):137–151, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-013-9325-z.

Oh:2014:UNS

[511] Jun-Young Oh. Understanding natural science based on abductive inference: Continental drift. Foundations of Science, 19(2):153–174, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9322-2.

Leal:2014:FNC

[512] João P. Leal. The forgotten names of chemical elements. Foundations of Science, 19(2):175–183, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-013-9326-y.

Franco-Mariscal:2014:HCW

[513] Antonio Joaquín Franco-Mariscal. How can we teach the chemical elements to make the memorization task more enjoyable? Foundations of Science, 19(2):185–188, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9355-1.

Zolfagharian:2014:TKS

[514] Mohammadreza Zolfagharian, Reza Akbari, and Hamidreza Fartookzadeh. Theory of knowledge in system dynamics models. Foundations of Science, 19(2):189–207, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-013-9328-9.

Romero:2014:CS

[515] Gustavo E. Romero. The collapse of supertasks. Foundations of Science, 19(2):209–216, June 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-013-9338-7.

Garola:2014:EQM

[516] Claudio Garola and Marco Persano. Embedding quantum mechanics into a broader noncontextual theory. Foundations of Science, 19(3):217–239, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9341-z.

Khatam:2014:OIE

[517] Iman Khatam and Afshin Shafiee. Objective information in the empiricist view of von Weizsäcker. Foundations of Science, 19(3):241–255, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9335-x.

Gornitz:2014:BUM

[518] Thomas Görnitz. The basis for an understanding of matter and mind. Foundations of Science, 19(3):257–262, August 2014. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9343-5.

Nescolarde-Selva:2014:SVI

[519] Josué Antonio Nescolarde-Selva and Josep-Lluis Usó-Doménech. Semiotic vision of ideologies. Foundations of Science, 19(3):263–282, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
013-9329-8.

Gash:2014:FPI

[520] Hugh Gash. Fixed or probable ideas? Foundations of Science, 19 (3):283–284, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-013-9330-2.

Nescolarde-Selva:2014:MMT

[521] J. Nescolarde-Selva and J. L. Usó-Doménech. Model, metamodel and topology. Foundations of Science, 19(3):285–288, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-013-9334-y.

Nescolarde-Selva:2014:RSI

[522] J. Nescolarde-Selva and J. L. Usó-Doménech. Reality, systems and impure systems. Foundations of Science, 19(3):289–306, August 2014. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-013-9337-8.

Gash:2014:CTR

[523] Hugh Gash. Constructivism, truth and reality. Foundations of Science, 19(3):307–309, August 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9351-5.

Uso-Domenech:2014:RFI

[524] J. L. Usó-Doménech and J. Nescolarde-Selva. Revealing the face of Isis. Foundations of Science, 19(3):311–318, August 2014. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9361-3.

Mamedov:2014:PNE

[525] B. A. Mamedov and M. Y. Esmer. On the philosophical nature of Einstein’s mass–energy equivalence formula E = mc^{2}. Foundations of Science, 19(4):319–329, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-013-9339-6.

Weinert:2014:LDK

[526] Friedel Weinert. Lines of descent: Kuhn and beyond. Foundations of Science, 19(4):331–352, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-013-9342-y.

Wallis:2014:AIB

[527] Steven E. Wallis. Abstraction and insight: Building better conceptual systems to support more effective social change. Foundations of Science, 19(4):353–362, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9344-4.

Park:2014:MC

[528] Woosuk Park. Misrepresentation in context. Foundations of Science, 19 (4):363–374, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9345-3.

Niiniluoto:2014:RT

[529] Ilkka Niiniluoto. Representation and truthlikeness. Foundations of Science, 19(4):375–379, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9346-2.

Pietarinen:2014:MMC

[530] Ahti-Veikko Pietarinen. Misrepresentation in “misrepresentation in context” in context. Foundations of Science, 19(4):381–386, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9379-6.

deSojo:2014:SDD

[531] Áurea Anguera de Sojo, Juan Ares, María Aurora Martínez, Juan Pazos, Santiago Rodríguez, and José Gabriel Zato. Serendipity and the discovery of DNA. Foundations of Science, 19(4):387–401, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9348-0.

Edis:2014:BPP

[532] Taner Edis and Maarten Boudry. Beyond physics? On the prospects of finding a meaningful oracle. Foundations of Science, 19(4):403– 422, November 2014. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9349-z.

Kanovei:2015:PRW

[533] Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, and Mary Schaps. Proofs and retributions, or: Why Sarah can’t take limits. Foundations of Science, 20(1):1–25, March 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-013-9340-0.

Nescolarde-Selva:2015:LKR

[534] J. Nescolarde-Selva, J. L. Usó-Doménech, and M. J. Sabán. Linguistic knowledge of reality: a metaphysical impossibility? Foundations of Science, 20(1):27–58, March 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9347-1.

Krysztofiak:2015:HSF

[535] Wojciech Krysztofiak. Hyper-slingshot. is fact-arithmetic possible? Foundations of Science, 20(1):59–76, March 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-014-9350-6.

deBianchi:2015:GMP

[536] Massimiliano Sassoli de Bianchi. God may not play dice, but human observers surely do. Foundations of Science, 20(1):77–105, March 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9352-4.

Bandyopadhyay:2015:HUU

[537] Prasanta S. Bandyopadhyay, John G. Bennett, and Megan D. Higgs. How to undermine underdetermination? Foundations of Science, 20 (2):107–127, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9353-3.

Bunge:2015:DAB

[538] Mario Bunge. Does the Aharonov–Bohm effect occur? Foundations of Science, 20(2):129–133, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9354-2.

Romero:2015:PT

[539] Gustavo E. Romero. Present time. Foundations of Science, 20(2): 135–145, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9356-0.

Kertesz:2015:PTE

[540] András Kertész. The puzzle of thought experiments in conceptual metaphor research. Foundations of Science, 20(2):147–174, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9357-z.

Ekstig:2015:CNS

[541] Börje Ekstig. Complexity, natural selection and the evolution of life and humans. Foundations of Science, 20(2):175–187, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9358-y;   http://link.springer.com/content/pdf/10.1007/
s10699-014-9358-y.pdf.

Wallis:2015:AIB

[542] Steven E. Wallis. Abstraction and insight: Building better conceptual systems to support more effective social change. Foundations of Science, 20(2):189–198, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9359-x. See erratum [647].

Ferrero:2015:CMR

[543] Miguel Ferrero and J. L. Sánchez-Gómez. Coming from material reality. Foundations of Science, 20(2):199–212, June 2015. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-014-9362-2.

Wallis:2015:SLP

[544] Steven E. Wallis. Structures of logic in policy and theory: Identifying sub-systemic bricks for investigating, building, and understanding conceptual systems. Foundations of Science, 20(3):213–231, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9360-4.

Macchia:2015:PSE

[545] Giovanni Macchia. Philosophy of space and expanding Universe in G. J. Whitrow. Foundations of Science, 20(3):233–247, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9363-1.

Kun:2015:IOE

[546] Wu Kun and Joseph E. Brenner. An informational ontology and epistemology of cognition. Foundations of Science, 20(3):249–279, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

014-9364-0. See erratum [547].

Wu:2015:EIO

[547] Kun Wu and Joseph E. Brenner. Erratum to: An Informational Ontology and Epistemology of Cognition. Foundations of Science, 20(3):281, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9380-0;   http://link.springer.com/content/pdf/10.1007/
s10699-014-9380-0.pdf. See [546].

Weinbaum:2015:CPB

[548] David R. Weinbaum. Complexity and the philosophy of becoming. Foundations of Science, 20(3):283–322, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-014-9370-2.

Harbecke:2015:RCL

[549] Jens Harbecke. Regularity constitution and the location of mechanistic levels. Foundations of Science, 20(3):323–338, August 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9371-1.

Cellucci:2015:MBU

[550] Carlo Cellucci. Mathematical beauty, understanding, and discovery. Foundations of Science, 20(4):339–355, November 2015. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-014-9378-7.

Santos:2015:TRI

[551] Emilio Santos. Towards a realistic interpretation of quantum mechanics providing a model of the physical world. Foundations of Science, 20(4): 357–386, November 2015. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9366-y.

Fiscaletti:2015:BES

[552] Davide Fiscaletti and Amrit Sorli. Bijective epistemology and space– time. Foundations of Science, 20(4):387–398, November 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9381-z.

Aerts:2015:MMM

[553] Diederik Aerts and Massimiliano Sassoli de Bianchi. Many-measurements or many-worlds? A dialogue. Foundations of Science, 20(4):399– 427, November 2015. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9382-y.

Santos:2015:OEH

[554] Gil C. Santos. Ontological emergence: How is that possible? Towards a new relational ontology. Foundations of Science, 20(4):429–446, November 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
015-9419-x.

Nugayev:2015:CRM

[555] Rinat M. Nugayev. Communicative rationality of the Maxwellian revolution. Foundations of Science, 20(4):447–478, November 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9417-z.

Froeyman:2016:ISE

[556] Anton Froeyman, Laszlo Kosolosky, and Jeroen Van Bouwel. Introduction: Social epistemology meets the philosophy of the humanities. Foundations of Science, 21(1):1–13, March 2016. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9372-0;

 http://link.springer.com/content/pdf/10.1007/s10699-014-9372-
0.pdf.

Miller:2016:SCE

[557] Boaz Miller. Scientific consensus and expert testimony in courts: Lessons from the Bendectin litigation. Foundations of Science, 21(1):15–33, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9373-z.

Stegenga:2016:TCC

[558] Jacob Stegenga. Three criteria for consensus conferences. Foundations of Science, 21(1):35–49, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9374-y.

Lacey:2016:SRN

[559] Hugh Lacey. Science, respect for nature, and human well-being: Democratic values and the responsibilities of scientists today. Foundations of Science, 21(1):51–67, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9376-9.

King:2016:IDE

[560] L. King, B. Morgan-Olsen, and J. Wong. Identifying difference, engaging dissent: What is at stake in democratizing knowledge? Foundations of Science, 21(1):69–88, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9375-x.

Jukola:2016:CRQ

[561] S. Jukola. The commercialization of research and the quest for the objectivity of science. Foundations of Science, 21(1):89–103, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9377-8.

Stearns:2016:WHI

[562] Peter N. Stearns. World history, identity and political change. Foundations of Science, 21(1):105–115, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-014-9367-x.

Seidel:2016:CSS

[563] M. Seidel. Changing society by scientific investigations? The unexpected shared ground between early sociology of knowledge and the Vienna Circle. Foundations of Science, 21(1):117–128, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9368-9.

DeVos:2016:DRP

[564] J. De Vos. The death and the resurrection of (psy)critique: The case of neuroeducation. Foundations of Science, 21(1):129–145, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9369-8.

Uso-Domenech:2016:WBS

[565] J. L. Usó-Doménech and J. Nescolarde-Selva. What are belief systems? Foundations of Science, 21(1):147–152, March 2016. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-015-9409-z;
 http://link.springer.com/content/pdf/10.1007/s10699-015-9409-

z.pdf. See correction [860].

Nescolarde-Selva:2016:TTC

[566] J. Nescolarde-Selva and J. L. Usó-Doménech. Textual theory and complex belief systems: Topological theory. Foundations of Science, 21 (1):153–175, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9410-6.

Gash:2016:SB

[567] Hugh Gash. Systems and beliefs. Foundations of Science, 21(1):177–187, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
015-9411-5.

Segura:2016:RBM

[568] Lorena Segura and Juan Matías Sepulcre. A rational belief: The method of discovery in the complex variable. Foundations of Science, 21(1): 189–194, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9412-4.

Poli:2016:BSM

[569] Roberto Poli. Belief systems and the modeling relation. Foundations of Science, 21(1):195–206, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9413-3.

Segura:2016:ARB

[570] Lorena Segura and Juan Matías Sepulcre. Arithmetization and rigor as beliefs in the development of mathematics. Foundations of Science, 21(1):207–214, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9414-2.

SignesPont:2016:CMB

[571] María Teresa Signes Pont, Higinio Mora Mora, Gregorio De Miguel Casado, and David Gil Méndez. A computational model of the belief system under the scope of social communication. Foundations of Science, 21(1):215–223, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9415-1.

Bueno:2016:BSP

[572] Otávio Bueno. Belief systems and partial spaces. Foundations of Science, 21(1):225–236, March 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9416-0.

Claes:2016:IMV

[573] Erik Claes and Nicole Note. Introduction. Meaningfulness, volunteers, citizenship. Foundations of Science, 21(2):237–251, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9383-x.

Wolf:2016:MTD

[574] Susan Wolf. Meaningfulness: a third dimension of the good life. Foundations of Science, 21(2):253–269, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-014-9384-9. See response [576].

Burms:2016:FFF

[575] Arnold Burms. Fulfillment and fitting fulfillment. Foundations of Science, 21(2):271–273, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9385-8.

Calhoun:2016:RLR

[576] Cheshire Calhoun. Reasons of love: Response to Wolf. Foundations of Science, 21(2):275–277, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9386-7. See [574].

Wolf:2016:MLM

[577] Susan Wolf. Meaning in life: Meeting the challenges. Foundations of Science, 21(2):279–282, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9387-6.

Note:2016:MVB

[578] Nicole Note and Emilie Van Daele. Meaningfulness, volunteering and being moved: The event of wit(h)nessing. Foundations of Science, 21 (2):283–300, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9388-5.

Burggraeve:2016:VEM

[579] Roger Burggraeve. Volunteering and ethical meaningfulness. Foundations of Science, 21(2):301–304, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-014-9389-4. See response [581].

Lindfelt:2016:MLF

[580] Mikael Lindfelt. Meaning of life in fragile wit(h)nessing: On experiencing radical uniqueness as gift and grace. Foundations of Science, 21(2):305–309, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9390-y. See response [581].

Note:2016:BRM

[581] Nicole Note. Being-with: Response to Mikael Lindfelt and Roger Burggraeve. Foundations of Science, 21(2):311–314, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9391-x. See [579,

vonEssen:2016:MVS

[582] Johan von Essen. On the meaning of volunteering: a study of worldviews in everyday life. Foundations of Science, 21(2):315–333, June 2016. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-014-9392-9.

VanDaele:2016:DPM

[583] Emilie Van Daele. Different perspectives on meaning and meaningfulness. Foundations of Science, 21(2):335–338, June 2016. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-014-9393-8.

Evers:2016:MVG

[584] Adalbert Evers. The meaning of volunteering: The general and constant versus the differentiating and shifting. Foundations of Science, 21(2):339–342, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9394-7.

vonEssen:2016:DCE

[585] Johan von Essen. The dynamics of change in everyday life: Final response. Foundations of Science, 21(2):343–346, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9395-6.

Claes:2016:CM

[586] Erik Claes. Civic meaningfulness. Foundations of Science, 21(2):347–372, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

014-9396-5. See responses [588, 587].

Honohan:2016:MCP

[587] Iseult Honohan. Meaningfulness: Civic or political? A response to Erik Claes’s “Civic Meaningfulness”. Foundations of Science, 21(2):373–375, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9397-4. See [586].

Dekker:2016:MCM

[588] Paul Dekker. Meaningful civicness for the many: a comment on Erik Claes. Foundations of Science, 21(2):377–380, June 2016. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-014-9408-5. See [586].

Claes:2016:CMR

[589] Erik Claes. Civic meaningfulness ‘revisited’: a final reply to Honohan and Dekker. Foundations of Science, 21(2):381–384, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9398-3.

Dominelli:2016:CVM

[590] Lena Dominelli. Citizenship and voluntarism: a meaningful combination or the basis for exploitative relationships? Foundations of Science, 21 (2):385–397, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-014-9399-2. See replies [592, 591].

DeBie:2016:VCR

[591] Maria De Bie and Rudi Roose. Voluntarism and citizenship: a response to Lena Dominelli. Foundations of Science, 21(2):399–403, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9400-0. See [590].

Claes:2016:CMP

[592] Erik Claes. Civic meaningfulness: The political dimension — a reply to Lena Dominelli. Foundations of Science, 21(2):405–408, June 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
014-9401-z. See [590].

Romero:2016:SRR

[593] Gustavo E. Romero. Sufficient reason and reason enough. Foundations of Science, 21(3):455–460, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9418-y.

Persson:2016:FF

[594] Ulf Persson. Is falsification falsifiable? Foundations of Science, 21 (3):461–475, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9420-4.

Cofnas:2016:SAS

[595] Nathan Cofnas. Science is not always “self-correcting”. Foundations of Science, 21(3):477–492, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9421-3.

Hautamaki:2016:PVC

[596] Antti Hautamäki. Points of view: a conceptual space approach. Foundations of Science, 21(3):493–510, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9422-2.

Park:2016:WIH

[597] Woosuk Park. What if haecceity is not a property? Foundations of Science, 21(3):511–526, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9429-8.

Schubring:2016:CPA

[598] Gert Schubring. Comments on a paper on alleged misconceptions regarding the history of analysis: Who has misconceptions? Foundations of Science, 21(3):527–532, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9424-0.

Viaro:2016:FMT

[599] Umberto Viaro. Feedback models of two classical philosophical positions and a semantic problem. Foundations of Science, 21(3):533–542, August 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
015-9430-2.

Ulazia:2016:MRA

[600] Alain Ulazia. Multiple roles for analogies in the genesis of fluid mechanics: How analogies can cooperate with other heuristic strategies. Foundations of Science, 21(4):543–565, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9423-1.

Oh:2016:USI

[601] Jun-Young Oh. Understanding scientific inquiries of Galileo’s formulation for the law of free falling motion. Foundations of Science, 21(4): 567–578, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9426-y.

Wallis:2016:SCS

[602] Steven E. Wallis. The science of conceptual systems: a progress report. Foundations of Science, 21(4):579–602, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9425-z.

Park:2016:RVS

[603] Seungbae Park. Realism versus surrealism. Foundations of Science, 21 (4):603–614, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9427-x.

Uso-Domenech:2016:MPS

[604] José-Luis Usó-Doménech, Josué Antonio Nescolarde Selva, and Mónica Belmonte Requena. Mathematical, philosophical and semantic considerations on infinity (I): General concepts. Foundations of Science, 21(4): 615–630, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9428-9.

R:2016:CBS

[605] Humberto Maturana R., Ximena Dávila Yáñez, and Simón Ramírez Muñoz. Cultural-biology: Systemic consequences of our evolutionary natural drift as molecular autopoietic systems. Foundations of Science, 21(4):631–678, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9431-1.

Herce:2016:PWS

[606] Rubén Herce. Penrose on what scientists know. Foundations of Science, 21(4):679–694, November 2016. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9432-0.

Zwick:2017:FNP

[607] Martin Zwick. Freedom as a natural phenomenon. Foundations of Science, 22(1):1–10, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9433-z.

Luk:2017:TSS

[608] Robert W. P. Luk. A theory of scientific study. Foundations of Science, 22(1):11–38, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9435-x.

Last:2017:BHF

[609] Cadell Last. Big historical foundations for deep future speculations: Cosmic evolution, atechnogenesis, and technocultural civilization. Foundations of Science, 22(1):39–124, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9434-y.

Blaszczyk:2017:CFA

[610] Piotr Blaszczyk, Vladimir Kanovei, Mikhail G. Katz, and David Sherry. Controversies in the foundations of analysis: Comments on Schubring’s conflicts. Foundations of Science, 22(1):125–140, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9473-4.

Romero:2017:OSS

[611] Gustavo E. Romero. On the ontology of spacetime: Substantivalism, relationism, eternalism, and emergence. Foundations of Science, 22 (1):141–159, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9476-1.

Park:2017:PDS

[612] Woosuk Park. A possible dilemma for situation semanticists. Foundations of Science, 22(1):161–182, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9477-0.

Cellucci:2017:MPS

[613] Carlo Cellucci. Is mathematics problem solving or theorem proving? Foundations of Science, 22(1):183–199, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-015-9475-2.

Sankey:2017:RPH

[614] Howard Sankey. Realism, progress and the historical turn. Foundations of Science, 22(1):201–214, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9481-4.

Todt:2017:NCV

[615] Oliver Todt and José Luis Luján. Non-cognitive values and methodological learning in the decision-oriented sciences. Foundations of Science, 22(1):215–234, March 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9482-3.

VanDenEede:2017:ALT

[616] Yoni Van Den Eede, Gert Goeminne, and Marc Van den Bossche. The art of living with technology: Turning over philosophy of Technology’s empirical turn. Foundations of Science, 22(2):235–246, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
015-9472-5;   http://link.springer.com/content/pdf/10.1007/
s10699-015-9472-5.pdf.

VandenBossche:2017:RNA

[617] Marc Van den Bossche. Releasement and nihilism in the art of living with technology. Foundations of Science, 22(2):247–253, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9460-9.

Scharff:2017:LTT

[618] Robert C. Scharff. On living with technology through renunciation and releasement. Foundations of Science, 22(2):255–260, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9462-7.

Babich:2017:HV

[619] Babette Babich. Heidegger on Verfallenheit. Foundations of Science, 22(2):261–264, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9461-8.

VandenBossche:2017:HA

[620] Marc Van den Bossche. On Heidegger’s actuality. Foundations of Science, 22(2):265–267, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9463-6.

Puech:2017:NCA

[621] Michel Puech. A non-confrontational art of living in the technosphere and infosphere. Foundations of Science, 22(2):269–274, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9452-9.

Swierstra:2017:DOA

[622] Tsjalling Swierstra. Does an old art suffice for new problems? Foundations of Science, 22(2):275–278, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9454-7.

Paredis:2017:WWC

[623] Erik Paredis. Which wisdom can change the world? Foundations of Science, 22(2):279–282, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9453-8.

Puech:2017:WAN

[624] Michel Puech. Whose agency now? Foundations of Science, 22 (2):283–286, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9455-6.

Lemmens:2017:SAH

[625] Pieter Lemmens. Social autonomy and heteronomy in the age of ICT: The digital pharmakon and the (dis)empowerment of the general intellect. Foundations of Science, 22(2):287–296, June 2017. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-015-9468-1.

Milberry:2017:ACL

[626] Kate Milberry. The art of collectively loving well in the Digital Age. Foundations of Science, 22(2):297–300, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-015-9469-0.

Verbeek:2017:STT

[627] Peter-Paul Verbeek. The struggle for technology: Towards a realistic political theory of technology. Foundations of Science, 22(2):301–304, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
015-9470-7.

Lemmens:2017:LR

[628] Pieter Lemmens. Love and realism. Foundations of Science, 22(2): 305–310, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9471-6.

Dorrestijn:2017:COH

[629] Steven Dorrestijn. The care of our hybrid selves: Ethics in times of technical mediation. Foundations of Science, 22(2):311–321, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-015-9440-0.

Huijer:2017:CUF

[630] Marli Huijer. A critical use of Foucault’s art of living. Foundations of Science, 22(2):323–327, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9441-z.

Wellner:2017:ETP

[631] Galit Wellner. Ethics in times of posthumanism. Foundations of Science, 22(2):329–332, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-015-9442-y.

Dorrestijn:2017:URT

[632] Steven Dorrestijn. The uses of reason in times of technical mediation. Foundations of Science, 22(2):333–337, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-015-9443-x.

Coeckelbergh:2017:ALI

[633] Mark Coeckelbergh. The art of living with ICTs: The ethics–aesthetics of vulnerability coping and its implications for understanding and evaluating ICT cultures. Foundations of Science, 22(2):339–348, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-015-9436-9.

Gunkel:2017:BLT

[634] David J. Gunkel. Better living through technology. Foundations of Science, 22(2):349–352, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-015-9437-8.

Zwart:2017:ALN

[635] Hub Zwart. The art of living with NZT and ICT: Dialectics of an artistic case study. Foundations of Science, 22(2):353–356, June 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9438-7.

Bunge:2017:ESR

[636] Mario Bunge. Evaluating scientific research projects: The units of science in the making. Foundations of Science, 22(3):455–469, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-015-9474-3.

Rosenberger:2017:NNP

[637] Robert Rosenberger. Notes on a nonfoundational phenomenology of technology. Foundations of Science, 22(3):471–494, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-015-9480-5.

Yanofsky:2017:RSM

[638] Noson S. Yanofsky and Mark Zelcer. The role of symmetry in mathematics. Foundations of Science, 22(3):495–515, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-016-9486-7.

Mohrhoff:2017:QMN

[639] Ulrich J. Mohrhoff. Quantum mechanics in a new light. Foundations of Science, 22(3):517–537, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-016-9487-6.

Gutman:2017:MIF

[640] Alexander E. Gutman, Mikhail G. Katz, Taras S. Kudryk, and Semen S. Kutateladze. The mathematical intelligencer flunks the Olympics. Foundations of Science, 22(3):539–555, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9485-8.

Lissack:2017:SOS

[641] Michael Lissack. Second order science: Examining hidden presuppositions in the practice of science. Foundations of Science, 22(3):557– 573, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9483-x.

Lissack:2017:WSO

[642] Michael Lissack. What second order science reveals about scientific claims: Incommensurability, doubt, and a lack of explication. Foundations of Science, 22(3):575–593, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9484-9.

Calude:2017:DSC

[643] Cristian S. Calude and Giuseppe Longo. The deluge of spurious correlations in big data. Foundations of Science, 22(3):595–612, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9489-4.

Park:2017:WSW

[644] Seungbae Park. Why should we be pessimistic about antirealists and pessimists? Foundations of Science, 22(3):613–625, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-016-9490-y.

deBianchi:2017:TCA

[645] Massimiliano Sassoli de Bianchi. Theoretical and conceptual analysis of the celebrated 4π-symmetry neutron interferometry experiments. Foundations of Science, 22(3):627–653, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9491-x.

Uso-Domenech:2017:MPS

[646] José-Luis Usó-Doménech, Josué Antonio Nescolarde-Selva, Mónica Belmonte-Requena, and L. Segura-Abad. Mathematics, philosophical and semantic considerations on infinity (II): Dialectical vision. Foundations of Science, 22(3):655–674, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9488-5.

Wallis:2017:EAI

[647] Steven E. Wallis. Erratum to: Abstraction and insight: Building better conceptual systems to support more effective social change. Foundations of Science, 22(3):675, September 2017. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-017-9537-8;
 http://link.springer.com/content/pdf/10.1007/s10699-017-9537-

8.pdf. See [542].

Crespo:2017:FME

[648] Ricardo Crespo and Fernando Tohmé. The future of mathematics in economics: a philosophically grounded proposal. Foundations of Science, 22 (4):677–693, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9492-9.

Bunge:2017:WA

[649] Mario Bunge. Why axiomatize? Foundations of Science, 22(4):695– 707, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9493-8.

Schinckus:2017:CSQ

[650] Christophe Schinckus. From cubist simultaneity to quantum complementarity. Foundations of Science, 22(4):709–716, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9494-7;   http://link.springer.com/content/pdf/10.1007/
s10699-016-9494-7.pdf.

Blaszczyk:2017:LCE

[651] Piotr Blaszczyk, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Taras Kudryk, Thomas Mormann, and David Sherry. Is Leibnizian calculus embeddable in first order logic? Foundations of Science, 22(4): 717–731, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9495-6.

Wallis:2017:WST

[652] Steven E. Wallis and Vladislav Valentinov. What is sustainable theory? A Luhmannian perspective on the science of conceptual systems. Foundations of Science, 22(4):733–747, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-016-9496-5.

Gornitz:2017:CQT

[653] Thomas Görnitz. A century of quantum theory: Time for a change in thinking. Foundations of Science, 22(4):749–762, December 2017. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-016-9497-4.

Blaszczyk:2017:THM

[654] Piotr Blaszczyk, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Semen S. Kutateladze, and David Sherry. Toward a history of mathematics focused on procedures. Foundations of Science, 22(4):763–783, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9498-3.

Tavakol:2017:ITS

[655] Reza Tavakol and Fabio Gironi. The infinite turn and speculative explanations in cosmology. Foundations of Science, 22(4):785–798, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9499-2.

Morelato:2017:FSC

[656] Marie Morelato, Mark Barash, Lucas Blanes, Scott Chadwick, Jessirie Dilag, Unnikrishnan Kuzhiumparambil, Katie D. Nizio, Xanthe Spindler, and Sebastien Moret. Forensic science: Current state and perspective by a group of early career researchers. Foundations of Science, 22(4): 799–825, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9500-0.

Oh:2017:GEG

[657] Jun-Young Oh and Eui Chan Jeon. Greenhouse effects in global warming based on analogical reasoning. Foundations of Science, 22(4):827– 847, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9501-z.

Locker:2017:BSS

[658] Markus Locker. Blindness and seeing in systems epistemology: Alfred Locker’s trans-classical systems theory. Foundations of Science, 22(4): 849–862, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9502-y.

Licata:2017:ECF

[659] Ignazio Licata and Gianfranco Minati. Emergence, computation and the freedom degree loss information principle in complex systems. Foundations of Science, 22(4):863–881, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-016-9503-x.

Borgatti:2018:SHC

[660] Elena Borgatti, Daniele Casagrande, Wiesław Krajewski, and Umberto Viaro. A study of history from a control-theory perspective. Foundations of Science, 23(1):1–16, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-016-9504-9.

Caamano-Alegre:2018:DTP

[661] María Caamaño-Alegre. Drift theory and plate tectonics: a case of embedding in geology. Foundations of Science, 23(1):17–35, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-016-9505-8.

Protzko:2018:DMC

[662] John Protzko. Disentangling mechanisms from causes: And the effects on science. Foundations of Science, 23(1):37–50, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-016-9511-x.

Gornitz:2018:SIH

[663] Thomas Görnitz and Uwe Schomäcker. The structures of interactions: How to explain the gauge groups U(1), SU(2) and SU(3). Foundations of Science, 23(1):51–73, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-016-9507-6.

Lopez-Corona:2018:HFC

[664] O. López-Corona, P. Padilla, O. Escolero, and E. Morales-Casique. Heuristic formulation of a contextual statistic theory for groundwater. Foundations of Science, 23(1):75–83, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9508-5.

VanDaele:2018:VES

[665] Wim Van Daele. Vibrantly entangled in Sri Lanka: Food as the polyrhythmic and polyphonic assemblage of life. Foundations of Science, 23(1):85–102, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9509-4.

Luk:2018:IEL

[666] Robert Luk. On the implications and extensions of Luk’s theory and model of scientific study. Foundations of Science, 23(1):103–118, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9510-y.

Holik:2018:PRN

[667] Federico Holik, Giuseppe Sergioli, Hector Freytes, and Angelo Plastino. Pattern recognition in non–Kolmogorovian structures. Foundations of Science, 23(1):119–132, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9520-4.

Bascelli:2018:GSO

[668] Tiziana Bascelli, Piotr Blaszczyk, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Semen S. Kutateladze, Tahl Nowik, David M. Schaps, and David Sherry. Gregory’s sixth operation. Foundations of Science, 23(1):133–144, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9512-9.

Gao:2018:ECC

[669] Shan Gao. Is an electron a charge cloud? A reexamination of Schrödinger’s charge density hypothesis. Foundations of Science, 23 (1):145–157, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-017-9521-3.

Tylman:2018:CSP

[670] Wojciech Tylman. Computer science and philosophy: Did Plato foresee object-oriented programming? Foundations of Science, 23(1):159–172, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-

016-9506-7. See response [671].

Demey:2018:PTM

[671] Lorenz Demey. The porphyrian tree and multiple inheritance: a rejoinder to Tylman on computer science and philosophy. Foundations of Science, 23(1):173–180, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-017-9531-1. See [670].

Khrennikov:2018:TBU

[672] Andrei Khrennikov. Towards better understanding QBism. Foundations of Science, 23(1):181–195, March 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9524-0;   http://link.springer.com/
content/pdf/10.1007/s10699-017-9524-0.pdf.

LeRoy:2018:DRT

[673] Frederik Le Roy and Robrecht Vanderbeeken. The documentary real: Thinking documentary aesthetics. Foundations of Science, 23 (2):197–205, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-016-9513-8.

Jacobs:2018:DSF

[674] Steven Jacobs. The documentary surreal: Film and painting in Luciano Emmer’s La Leggenda di Sant’Orsola (1948) and Henri Storck’s Le Monde de Paul Delvaux (1946). Foundations of Science, 23(2):207–215, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9514-7.

Carels:2018:RTT

[675] Edwin Carels. Revisiting Tom Tom: Performative anamnesis and autonomous vision in Ken Jacobs’ appropriations of Tom Tom the Piper’s Son. Foundations of Science, 23(2):217–230, June 2018. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-016-9515-6.

Strauven:2018:UDV

[676] Wanda Strauven. The unseen déjà-vu: From Erkki Huhtamo’s topoi to Ken Jacobs’ remakes. Foundations of Science, 23(2):231–236, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
016-9516-5;   http://link.springer.com/content/pdf/10.1007/
s10699-016-9516-5.pdf.

Bruzzi:2018:AMM

[677] Stella Bruzzi. Approximation, mad men and the death of JFK. Foundations of Science, 23(2):237–244, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-016-9517-4;   http://link.springer.
com/content/pdf/10.1007/s10699-016-9517-4.pdf.

DeKesel:2018:DRS

[678] Marc De Kesel. The documentary real and the Shoah. Foundations of Science, 23(2):245–254, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-016-9518-3.

Broekaert:2018:TQM

[679] Jan Broekaert. The tacit ‘quantum’ of meeting the aesthetic sign; contextualize, entangle, superpose, collapse or decohere. Foundations of Science, 23(2):255–266, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-016-9519-2;   http://link.springer.com/
content/pdf/10.1007/s10699-016-9519-2.pdf.

Bascelli:2018:CIH

[680] Tiziana Bascelli, Piotr Blaszczyk, Alexandre Borovik, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, Semen S. Kutateladze, Thomas Mc-Gaffey, David M. Schaps, and David Sherry. Cauchy’s infinitesimals, his sum theorem, and foundational paradigms. Foundations of Science, 23(2):267–296, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-017-9534-y.

Uso-Domenech:2018:ISE

[681] José-Luis Usó-Doménech, Josué-Antonio Nescolarde-Selva, and Miguel Lloret-Climent. Impure systems and ecological models (I): Axiomatization. Foundations of Science, 23(2):297–321, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-017-9522-2.

Aerts:2018:SWDa

[682] Diederik Aerts, Jonito Aerts Arguëlles, Lester Beltran, Suzette Geriente, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz. Spin and wind directions I: Identifying entanglement in nature and cognition. Foundations of Science, 23(2):323–335, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
017-9528-9;   http://link.springer.com/content/pdf/10.1007/
s10699-017-9528-9.pdf.

Aerts:2018:SWDb

[683] Diederik Aerts, Jonito Aerts Arguëlles, Lester Beltran, Suzette Geriente, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz.

Spin and wind directions II: a Bell state quantum model. Foundations of Science, 23(2):337–365, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9530-2;   http://link.springer.com/
content/pdf/10.1007/s10699-017-9530-2.pdf.

Sherry:2018:JMI

[684] David Sherry. The Jesuits and the method of indivisibles. Foundations of Science, 23(2):367–392, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9525-z. See responses [685, 696].

Alexander:2018:IIR

[685] Amir Alexander. On indivisibles and infinitesimals: a response to David Sherry, “The Jesuits and the Method of Indivisibles”. Foundations of Science, 23(2):393–398, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9523-1. See [684, 696].

Bunge:2018:GWS

[686] Mario Bunge. Gravitational waves and spacetime. Foundations of Science, 23(2):399–403, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9526-y. See comments [687].

Romero:2018:MBG

[687] Gustavo E. Romero. Mario Bunge on gravitational waves and the reality of spacetime. Foundations of Science, 23(2):405–409, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-017-9529-8. See [686].

Luk:2018:EPW

[688] Robert W. P. Luk. To explain or to predict: Which one is mandatory? Foundations of Science, 23(2):411–414, June 2018. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-017-9533-z.

Romero:2018:CBH

[689] Gustavo E. Romero, Daniela Pérez, and Federico G. López Armengol. Cosmological black holes and the direction of time. Foundations of Science, 23(2):415–426, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-017-9527-x.

Cirkovic:2018:PCB

[690] Milan M. Ćirković and Jelena Dimitrijević. Putting the cart before the horse: Co-evolution of the Universe and observers as an explanatory hypothesis. Foundations of Science, 23(3):427–442, September 2018. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-017-9532-0.

Longo:2018:HFD

[691] Giuseppe Longo. How future depends on past and rare events in systems of life. Foundations of Science, 23(3):443–474, September 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-017-9535-x.

Gornitz:2018:QTN

[692] Thomas Görnitz. Quantum theory and the nature of consciousness. Foundations of Science, 23(3):475–510, September 2018. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-017-9536-9.

Aerts:2018:RTR

[693] Diederik Aerts. Relativity theory refounded. Foundations of Science, 23 (3):511–547, September 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-017-9538-7;   http://link.springer.com/content/
pdf/10.1007/s10699-017-9538-7.pdf.

dePaz:2018:PGF

[694] María de Paz. Poincaré on generalizations and facts: Construction or translation? Foundations of Science, 23(3):549–558, September 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-017-9539-6.

Bair:2018:FDW

[695] Jacques Bair, Mikhail G. Katz, and David Sherry. Fermat’s Dilemma: Why did he keep mum on infinitesimals? And the European theological context. Foundations of Science, 23(3):559–595, September 2018. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-017-9542-y.

Radelet-deGrave:2018:CII

[696] Patricia Radelet-de Grave. Comments on indivisibles and infinitesimals: a response to David Sherry, by Amir Alexander: In view of the original book. Foundations of Science, 23(4):597–602, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-017-9540-0. See [685,

684].

Stencel:2018:HRM

[697] Adrian Stencel and Agnieszka M. Proszewska. How research on microbiomes is changing biology: a discussion on the concept of the organism. Foundations of Science, 23(4):603–620, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
017-9543-x;   http://link.springer.com/content/pdf/10.1007/
s10699-017-9543-x.pdf.

deRonde:2018:QSR

[698] Christian de Ronde. Quantum superpositions and the representation of physical reality beyond measurement outcomes and mathematical structures. Foundations of Science, 23(4):621–648, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-017-9541-z.

Bathfield:2018:WZP

[699] Maël Bathfield. Why Zeno’s paradoxes of motion are actually about immobility. Foundations of Science, 23(4):649–679, December 2018. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-017-9544-9.

Trlifajova:2018:BIQ

[700] Katerina Trlifajová. Bolzano’s infinite quantities. Foundations of Science, 23(4):681–704, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-018-9549-z.

Romero-Maltrana:2018:PCO

[701] Diego Romero-Maltrana, Federico Benitez, and Cristian Soto. A proposal for a coherent ontology of fundamental entities. Foundations of Science, 23(4):705–717, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-018-9548-0.

Lenhard:2018:AMP

[702] Johannes Lenhard and Michael Otte. The applicability of mathematics as a philosophical problem: Mathematization as exploration. Foundations of Science, 23(4):719–737, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-018-9546-2.

deBoer:2018:PIM

[703] Bas de Boer, Hedwig Te Molder, and Peter-Paul Verbeek. The perspective of the instruments: Mediating collectivity. Foundations of Science, 23(4):739–755, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-018-9545-3;   http://link.springer.com/
content/pdf/10.1007/s10699-018-9545-3.pdf.

Arguelles:2018:HIQ

[704] Jonito Aerts Arguëlles. The heart of an image: Quantum superposition and entanglement in visual perception. Foundations of Science, 23(4): 757–778, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9547-1.

Genova:2018:DPF

[705] Gonzalo Génova and Ignacio Quintanilla Navarro. Discovering the principle of finality in computational machines. Foundations of Science, 23 (4):779–794, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9552-4.

Romero:2018:OTS

[706] Gustavo E. Romero. Outline of a theory of scientific aesthetics. Foundations of Science, 23(4):795–807, December 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-018-9551-5.

Boi:2019:SME

[707] Luciano Boi. Some mathematical, epistemological, and historical reflections on the relationship between geometry and reality, space–time theory and the geometrization of theoretical physics, from Riemann to Weyl and beyond. Foundations of Science, 24(1):1–38, March 2019. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-018-9550-6.

Neves:2019:INP

[708] Juliano C. S. Neves. Infinities as natural places. Foundations of Science, 24(1):39–49, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9556-0.

Krause:2019:IRF

[709] Décio Krause and Jonas R. Becker Arenhart. Is identity really so fundamental? Foundations of Science, 24(1):51–71, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-018-9553-3.

Mets:2019:PCD

[710] Ave Mets. A philosophical critique of the distinction of representational and pragmatic measurements on the example of the periodic system of chemical elements. Foundations of Science, 24(1):73–93, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-018-9567-x.

Kuna:2019:SES

[711] Maciej Kuna. Systems, environments, and soliton rate equations: Toward realistic modeling. Foundations of Science, 24(1):95–132, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9568-9;   http://link.springer.com/content/pdf/10.1007/
s10699-018-9568-9.pdf.

Lombardi:2019:DBI

[712] Olimpia Lombardi and María J. Ferreira Ruiz. Distinguishing between inter-domain and intra-domain emergence. Foundations of Science, 24 (1):133–151, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9554-2.

Sergeyev:2019:IGB

[713] Yaroslav D. Sergeyev. Independence of the grossone-based infinity methodology from non-standard analysis and comments upon logical fallacies in some texts asserting the opposite. Foundations of Science, 24 (1):153–170, March 2019. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9566-y.

Stewart:2019:OLM

[714] John E. Stewart. The origins of life: The managed-metabolism hypothesis. Foundations of Science, 24(1):171–195, March 2019. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-018-9563-1;
 http://link.springer.com/content/pdf/10.1007/s10699-018-9563-
1.pdf.

Krause:2019:DNS

[715] Décio Krause. Does Newtonian space provide identity to quantum systems? Foundations of Science, 24(2):197–215, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-018-9561-3.

Aerts:2019:QTM

[716] Diederik Aerts, Lester Beltran, Suzette Geriente, Massimiliano Sassoli de Bianchi, Sandro Sozzo, Rembrandt Van Sprundel, and Tomas Veloz. Quantum theory methods as a possible alternative for the double-blind gold standard of evidence-based medicine: Outlining a new research program. Foundations of Science, 24(2):217–225, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9572-0;   http://link.springer.com/content/pdf/10.1007/
s10699-018-9572-0.pdf.

Beltran:2019:QEC

[717] Lester Beltran and Suzette Geriente. Quantum entanglement in corpuses of documents. Foundations of Science, 24(2):227–246, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9570-2;   http://link.springer.com/content/pdf/10.1007/
s10699-018-9570-2.pdf.

Crespo:2019:LNN

[718] Ricardo F. Crespo. Liberal naturalism and non-epistemic values. Foundations of Science, 24(2):247–273, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-018-9565-z.

Sanchez-Canizares:2019:CFW

[719] Javier Sánchez-Cañizares. Classicality first: Why Zurek’s existential interpretation of quantum mechanics implies Copenhagen. Foundations of Science, 24(2):275–285, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-018-9574-y.

Zenczykowski:2019:QHE

[720] Piotr Zenczykowski. Quarks, hadrons, and emergent spacetime. Foundations of Science, 24(2):287–305, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-018-9562-2;   http://link.springer.
com/content/pdf/10.1007/s10699-018-9562-2.pdf.

Solari:2019:CVN

[721] H. G. Solari and M. A. Natiello. A constructivist view of Newton’s mechanics. Foundations of Science, 24(2):307–341, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9573-z;   http://link.springer.com/content/pdf/10.1007/
s10699-018-9573-z.pdf.

Rodriguez-Vellando:2019:LED

[722] Pablo Rodríguez-Vellando. … and so Euler discovered differential equations. Foundations of Science, 24(2):343–374, June 2019. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-018-9571-1.

Drechsel:2019:FQM

[723] Paul Drechsel. Foundation of quantum mechanics: Once again. Foundations of Science, 24(2):375–389, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-018-9555-1.

Arfini:2019:IPM

[724] Selene Arfini, Claudia Casadio, and Lorenzo Magnani. Ignorancepreserving mental models thought experiments as abductive metaphors. Foundations of Science, 24(2):391–409, June 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-018-9564-0.

Ippoliti:2019:HIM

[725] Emiliano Ippoliti. Heuristics and inferential microstructures: The path to quaternions. Foundations of Science, 24(3):411–425, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-018-9576-9.

Nescolarde-Selva:2019:ISE

[726] Josué-Antonio Nescolarde-Selva, José-Luis Usó-Doménech, and Miguel Lloret-Climent. Impure systems and ecological models (II): Components and thermodynamics. Foundations of Science, 24(3):427–455, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9575-x.

Zwart:2019:FTP

[727] Hub Zwart. Fabricated truths and the pathos of proximity: What would be a Nietzschean philosophy of contemporary technoscience? Foundations of Science, 24(3):457–482, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09599-3;   http://link.
springer.com/content/pdf/10.1007/s10699-019-09599-3.pdf.

Bunge:2019:IP

[728] Mario Bunge. Inverse problems. Foundations of Science, 24(3):483– 525, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-09577-1.

Jho:2019:IAC

[729] Hunkoog Jho. Interdisciplinary approach to combine science and art: Understanding of the paintings of René Magritte from the viewpoint of quantum mechanics. Foundations of Science, 24(3):527–540, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
019-09600-z.

Peck:2019:RBD

[730] Steven L. Peck. The rumors of Bergson’s demise may have been exaggerated: Novelty, complexity, and emergence in biological evolution. Foundations of Science, 24(3):541–557, September 2019. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09598-4.

Weber:2019:SEV

[731] Erik Weber, Dingmar van Eck, and Julie Mennes. On the structure and epistemic value of function ascriptions in biology and engineering sciences. Foundations of Science, 24(3):559–581, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-019-09584-w.

Cellucci:2019:DM

[732] Carlo Cellucci. Diagrams in mathematics. Foundations of Science, 24(3): 583–604, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-019-09583-x.

Park:2019:DRS

[733] Seungbae Park. In defense of realism and selectivism from Lyons’s objections. Foundations of Science, 24(4):605–615, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-019-09614-7.

Wu:2019:WL

[734] Victoria Wu and Vuk Uskoković. Waiting for Aπαταω : 250 years later. Foundations of Science, 24(4):617–640, December 2019. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09602-x.

Uso-Domenech:2019:MLN

[735] J. L. Usó-Doménech and J. A. Nescolarde-Selva. Mathematical logic of notions and concepts. Foundations of Science, 24(4):641–655, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
019-09604-9.

Hohol:2019:CAG

[736] Mateusz Hohol and Marcin Miłkowski. Cognitive artifacts for geometric reasoning. Foundations of Science, 24(4):657–680, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09603-w;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09603-w.pdf.

Andreas:2019:HIB

[737] Holger Andreas and Lorenzo Casini. Hypothetical interventions and belief changes. Foundations of Science, 24(4):681–704, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-019-09605-8.

Lee:2019:WSF

[738] Clarissa Ai Ling Lee. What science fiction can demonstrate about novelty in the context of discovery and scientific creativity. Foundations of Science, 24(4):705–725, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-019-09615-6.

Uso-Domenech:2019:AOA

[739] J. L. Usó-Doménech, J. A. Nescolarde-Selva, and H. Gash. Argumentum ontologicum and argumentum ornithologicum: Anselm of Canterbury and Jorge Luis Borges. Foundations of Science, 24(4):727–749, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
019-09601-y.

Magazu:2019:VML

[740] Salvatore Magazù, Nella Coletta, and Federica Migliardo. The Vitruvian man of Leonardo da Vinci as a representation of an operational approach to knowledge. Foundations of Science, 24(4):751–773, December 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-019-09616-5.

Kertesz:2019:FHG

[741] András Kertész. Fallacies in the historiography of generative linguistics. Foundations of Science, 24(4):775–801, December 2019. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09612-9;   http://link.
springer.com/content/pdf/10.1007/s10699-019-09612-9.pdf.

Aerts:2020:PSI

[742] Diederik Aerts, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz. Preface of the special issue International Symposium “Worlds of Entanglement”. Foundations of Science, 25(1):1–4, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09636-1;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09636-1.pdf.

Aerts:2020:CIQ

[743] Diederik Aerts, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz. On the conceptuality interpretation of quantum and relativity theories. Foundations of Science, 25(1):5–54, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
018-9557-z;   http://link.springer.com/content/pdf/10.1007/
s10699-018-9557-z.pdf.

deRonde:2020:UOQ

[744] Christian de Ronde. Unscrambling the omelette of quantum contextuality (Part I): Preexistent properties or measurement outcomes? Foundations of Science, 25(1):55–76, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-019-09578-8.

deBianchi:2020:UAE

[745] Massimiliano Sassoli de Bianchi. Using abstract elastic membranes to learn about quantum measurements. Foundations of Science, 25 (1):77–85, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-018-9558-y.

Sergioli:2020:TMT

[746] Giuseppe Sergioli. Towards a multi target quantum computational logic. Foundations of Science, 25(1):87–104, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-018-9569-8.

Garola:2020:EIQ

[747] Claudio Garola. An epistemic interpretation of quantum probability via contextuality. Foundations of Science, 25(1):105–120, March 2020. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-018-9560-4.

Prakash:2020:ISW

[748] Chetan Prakash. On invention of structure in the world: Interfaces and conscious agents. Foundations of Science, 25(1):121–134, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-019-09579-7.

Tesar:2020:HDS

[749] Jakub Tesar. How do social norms and expectations about others influence individual behavior? Foundations of Science, 25(1):135–150, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10699-
019-09582-y.

Gunji:2020:DCB

[750] Yukio-Pegio Gunji and Kyoko Nakamura. Dancing chief in the brain or consciousness as an entanglement. Foundations of Science, 25(1):151– 184, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09585-9;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09585-9.pdf.

Gobbo:2020:EDQ

[751] Federico Gobbo and Federica Russo. Epistemic diversity and the question of lingua franca in science and philosophy. Foundations of Science, 25(1):185–207, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.com/

article/10.1007/s10699-019-09631-6;   http://link.springer.
com/content/pdf/10.1007/s10699-019-09631-6.pdf.

Naudts:2020:ECF

[752] Jan Naudts. Emergent Coulomb forces in reducible quantum electrodynamics. Foundations of Science, 25(1):209–245, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-019-09607-6.

Nakamura:2020:EAC

[753] Kyoko Nakamura and Yukio Pegio Gunji. Entanglement of art coefficient, or creativity. Foundations of Science, 25(1):247–257, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09586-8;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09586-8.pdf.

Veloz:2020:CSD

[754] Tomas Veloz. The complexity–stability debate, chemical organization theory, and the identification of non-classical structures in ecology. Foundations of Science, 25(1):259–273, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-019-09639-y.

Svozil:2020:TFG

[755] Karl Svozil. Towards fractal gravity. Foundations of Science, 25(1):275– 280, March 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09609-4;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09609-4.pdf.

Shoshani:2020:AS

[756] Yakir Shoshani and Asher Yahalom. Apriorics and structuralism. Foundations of Science, 25(2):281–296, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-019-09617-4.

DeHaro:2020:SPL

[757] Sebastian De Haro. Science and philosophy: A love–hate relationship. Foundations of Science, 25(2):297–314, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09619-2;   http://link.
springer.com/content/pdf/10.1007/s10699-019-09619-2.pdf.

Wang:2020:ASS

[758] Xijia Wang. Axiomatization of the symbols system of classic of changes: The marriage of Oriental mysticism and Western scientific tradition. Foundations of Science, 25(2):315–325, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.

springer.com/article/10.1007/s10699-019-09624-5.

Etesi:2020:PPS

[759] Gábor Etesi. On a possibly pure set–theoretic contribution to black hole entropy. Foundations of Science, 25(2):327–340, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-019-09625-4;   http://link.springer.com/content/pdf/
10.1007/s10699-019-09625-4.pdf.

Issajeva:2020:CTM

[760] Jelena Issajeva. Can theories of mental representation adequately explain mental imagery? Foundations of Science, 25(2):341–355, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-019-09613-8.

Facoetti:2020:DDC

[761] Marta Facoetti. Donald Davidson’s critiques of conceptual relativism applied to non-adaptationist evolutionary epistemology and refuted. Foundations of Science, 25(2):357–374, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-019-09606-7.

Barzegar:2020:IBQ

[762] Ali Barzegar, Afshin Shafiee, and Mostafa Taqavi. ”It from bit” and quantum mechanics. Foundations of Science, 25(2):375–384, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

http://link.springer.com/article/10.1007/s10699-019-09644-1.

Uso-Domenech:2020:OAI

[763] J. L. Usó-Doménech, J. A. Nescolarde-Selva, and Hugh Gash. Ontological argument and infinity in Spinoza’s thought. Foundations of Science, 25(2):385–400, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-019-09627-2.

Shomar:2020:CR

[764] Towfic Shomar. Complementarity revisited. Foundations of Science, 25(2):401–424, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL http://link.springer.com/article/10.
1007/s10699-019-09641-4.

Barreiro:2020:TCU

[765] C. Barreiro, Jose M. Barreiro, J. A. Lara, D. Lizcano, M. A. Martínez, and J. Pazos. The third construct of the universe: Information. Foundations of Science, 25(2):425–440, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-019-09630-7.

Osiejewicz:2020:TLC

[766] Joanna Osiejewicz. Transnational legal communication: Towards comprehensible and consistent law. Foundations of Science, 25(2):441–475, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL http://link.springer.com/article/10.1007/
s10699-020-09655-3;   http://link.springer.com/content/pdf/
10.1007/s10699-020-09655-3.pdf.

Pisano:2020:TTL

[767] Raffaele Pisano. A tale of Tartaglia’s Libro Sesto & La Gionta in Quesiti et Inventioni Diverse (1546–1554): Exploring the historical and cultural foundations. Foundations of Science, 25(2):477–505, June 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-019-09635-2.

Dua:2020:SDR

[768] Mikhael Dua. Scientific discovery and its rationality: Michael Polanyi’s epistemological exposition. Foundations of Science, 25(3):507–518, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-019-09621-8.

Livadas:2020:TOI

[769] Stathis Livadas. Is there an ontology of infinity? Foundations of Science, 25(3):519–540, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09669-x.

Symons:2020:WTN

[770] John Symons and Jack K. Horner. Why there is no general solution to the problem of software verification. Foundations of Science, 25(3): 541–557, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09611-w.

Lee:2020:RIC

[771] Chunghyoung Lee. A retrocausal interpretation of classical collision between rigid bodies. Foundations of Science, 25(3):559–571, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09623-6.

Dillern:2020:AKM

[772] Thomas Dillern. The act of knowing: Michael Polanyi meets contemporary natural science. Foundations of Science, 25(3):573–585, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-019-09626-3.

Uso-Domenech:2020:TAT

[773] J. L. Usó-Doménech, J. A. Nescolarde-Selva, and H. Gash. Theorizing about theories and mathematical existence. Foundations of Science, 25 (3):587–595, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09648-2.

Kuzemsky:2020:STL

[774] A. L. Kuzemsky. In search of time lost: Asymmetry of time and irreversibility in natural processes. Foundations of Science, 25(3):597– 645, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09658-0.

Kouneiher:2020:MDT

[775] Joseph Kouneiher and Newton da Costa. The mathematical descriptions of truth and change. Foundations of Science, 25(3):647–670, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-020-09686-w.

Bussotti:2020:HFD

[776] Paolo Bussotti and Raffaele Pisano. Historical and foundational details on the method of infinite descent: Every prime number of the form 4n + 1 is the sum of two squares. Foundations of Science, 25(3): 671–702, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09642-3.

Coelho:2020:CSN

[777] R. Lopes Coelho. A comment on Solari and Natiello’s constructivist view of Newton’s mechanics. Foundations of Science, 25(3):703–710, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-020-09651-7.

Rizzuti:2020:OAT

[778] B. F. Rizzuti, L. M. Gaio, and C. Duarte. Operational approach to the topological structure of the physical space. Foundations of Science, 25(3): 711–735, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09650-8.

Kvasz:2020:CUT

[779] Ladislav Kvasz. Cognitive unity of Thales’ mathematics. Foundations of Science, 25(3):737–753, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09622-7.

Aerts:2020:QSC

[780] Diederik Aerts and Lester Beltran. Quantum structure in cognition: Human language as a boson gas of entangled words. Foundations of Science, 25(3):755–802, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09633-4.

Javanmardi:2020:EHC

[781] Ehsan Javanmardi and Sifeng Liu. Exploring the human cognitive capacity in understanding systems: a grey systems theory perspective. Foundations of Science, 25(3):803–825, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-019-09618-3.

DelSanto:2020:DPG

[782] Flavio Del Santo and Chiara Cardelli. Demolishing prejudices to get to the foundations: A criterion of demarcation for fundamentality. Foundations of Science, 25(3):827–843, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-019-09629-0.

Luk:2020:WDW

[783] Robert W. P. Luk. What do we mean by “true” in scientific realism? Foundations of Science, 25(3):845–856, September 2020. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-019-09634-3.

Neves:2020:PDS

[784] Juliano C. S. Neves. Proposal for a degree of scientificity in cosmology. Foundations of Science, 25(3):857–878, September 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-019-09620-9.

Krol:2020:QMF

[785] Jerzy Król and Torsten Asselmeyer-Maluga. Quantum mechanics, formalization and the cosmological constant problem. Foundations of Science, 25(4):879–904, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09646-z.

Javanmardi:2020:EPP

[786] Ehsan Javanmardi, Sifeng Liu, and Naiming Xie. Exploring the philosophical paradigm of grey systems theory as a postmodern theory. Foundations of Science, 25(4):905–925, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-019-09640-5.

Gruber:2020:PTD

[787] Ronald P. Gruber, Carlos Montemayor, and Richard A. Block. From physical time to a dualistic model of human time. Foundations of Science, 25(4):927–954, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09670-4.

Soto:2020:LSR

[788] Cristian Soto and Diego Romero-Maltrana. Local selective realism: Shifting from classical to quantum electrodynamics. Foundations of Science, 25(4):955–970, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09663-3.

Czachor:2020:LAL

[789] Marek Czachor. A loophole of all ‘loophole-free’ Bell-type theorems. Foundations of Science, 25(4):971–985, December 2020. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:

//link.springer.com/article/10.1007/s10699-020-09666-0. See

comment [848].

Lambare:2020:IAF

[790] Justo Pastor Lambare. Incisive approach to Fermi–Walker transport. Foundations of Science, 25(4):987–1001, December 2020. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09679-9.

Lara:2020:SIF

[791] Juan A. Lara and Shadi Aljawarneh. Special issue on the foundations of software science and computation structures. Foundations of Science, 25(4):1003–1008, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09587-7.

Mohammed:2020:HEG

[792] Tareq Abed Mohammed, Oguz Bayat, and Shaymaa Alhayali. Hybrid efficient genetic algorithm for big data feature selection problems. Foundations of Science, 25(4):1009–1025, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-019-09588-6.

Vangipuram:2020:KSS

[793] Radhakrishna Vangipuram, P. V. Kumar, and Khalaf Khatatneh. Krishna Sudarsana — a Z-space interest measure for mining similarity profiled temporal association patterns. Foundations of Science, 25(4): 1027–1048, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09590-y.

Nagaraja:2020:UID

[794] Arun Nagaraja, B. Uma, and Rajesh Kumar Gunupudi. UTTAMA: An intrusion detection system based on feature clustering and feature transformation. Foundations of Science, 25(4):1049–1075, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09589-5.

Kotte:2020:SFF

[795] Vinay Kumar Kotte, Srinivasan Rajavelu, and Elijah Blessing Rajsingh. A similarity function for feature pattern clustering and high dimensional text document classification. Foundations of Science, 25(4): 1077–1094, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09592-w.

Yassein:2020:HRT

[796] Muneer Bani Yassein, Shadi Aljawarneh, and Yarub Wahsheh. Hybrid real-time protection system for online social networks. Foundations of Science, 25(4):1095–1124, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-019-09595-7.

Reddicherla:2020:SNU

[797] Vikram Raju Reddicherla, Umashankar Rawat, and Kumkum Garg. Securing NEMO using a bilinear pairing-based 3-party key exchange (3PKE-NEMO) in heterogeneous networks. Foundations of Science, 25(4):1125–1146, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.com/

article/10.1007/s10699-019-09591-x.

Aljawarneh:2020:UUL

[798] Shadi A. Aljawarneh, Vangipuram Radhakrishna, and John William Atwood. Ultimate: Unearthing latent time profiled temporal associations. Foundations of Science, 25(4):1147–1171, December 2020. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-019-09594-8.

delaPena:2020:NAC

[799] F. David de la Peña, Juan A. Lara, and Juan Pazos. A new approach to computing using informons and holons: Towards a theory of computing science. Foundations of Science, 25(4):1173–1201, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09597-5.

Anguera:2020:TGU

[800] Aurea Anguera, Juan A. Lara, and F. David de la Peña. Turing: The great unknown. Foundations of Science, 25(4):1203–1225, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09596-6.

AlHayali:2020:IEE

[801] Shaymaa Al Hayali, Javad Rahebi, and Oguz Bayat. Increasing energy efficiency in wireless sensor networks using GA-ANFIS to choose a cluster head and assess routing and weighted trusts to demodulate attacker nodes. Foundations of Science, 25(4):1227–1246, December 2020. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09593-9.

Aerts:2021:PSI

[802] Diederik Aerts, Massimiliano Sassoli de Bianchi, and Tomas Veloz. Preface of the special issue: International Symposium “Worlds of Entanglement” — Second Part. Foundations of Science, 26(1):1–4, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/

s10699-021-09785-2. See correction [803].

Aerts:2021:CPS

[803] Diederik Aerts, Massimiliano Sassoli de Bianchi, and Tomas Veloz. Correction to: Preface of the special Issue: International Symposium “Worlds of Entanglement” — Second Part. Foundations of Science, 26(1):5, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09793-2. See [802].

Schaffer:2021:OTI

[804] Kathryn Schaffer and Gabriela Barreto Lemos. Obliterating thingness: An introduction to the “what’ and the “so what” of quantum physics. Foundations of Science, 26(1):7–26, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-019-09608-5.

Aerts:2021:MHD

[805] Diederik Aerts, Massimiliano Sassoli de Bianchi, and Tomas Veloz. Modeling human decision-making: An overview of the Brussels quantum approach. Foundations of Science, 26(1):27–54, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-018-9559-x.

Langenhove:2021:ESR

[806] Luk Van Langenhove. The entanglement of the social realm: Towards a quantum theory inspired ontology for the social sciences. Foundations of Science, 26(1):55–73, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09717-6.

Czachor:2021:NNM

[807] Marek Czachor. Non-Newtonian mathematics instead of non-Newtonian physics: Dark matter and dark energy from a mismatch of arithmetics. Foundations of Science, 26(1):75–95, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09687-9.

Farouki:2021:EBD

[808] Nayla Farouki and Philippe Grangier. The Einstein–Bohr debate: Finding a common ground of understanding? Foundations of Science, 26 (1):97–101, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09716-7.

Sozzo:2021:RAT

[809] Sandro Sozzo. Representing attitudes towards ambiguity in Hilbert space: Foundations and applications. Foundations of Science, 26(1): 103–128, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09718-5.

Lawless:2021:TEI

[810] W. F. Lawless. Towards an epistemology of interdependence among the orthogonal roles in human-machine teams. Foundations of Science, 26 (1):129–142, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09632-5.

deBianchi:2021:NSR

[811] Massimiliano Sassoli de Bianchi. A non-spatial reality. Foundations of Science, 26(1):143–170, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09719-4.

Bowman:2021:HTE

[812] Norah Bowman. Here/there/everywhere: Quantum models for decolonizing Canadian state onto-epistemology. Foundations of Science, 26 (1):171–186, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09610-x.

Kordes:2021:TEN

[813] Urban Kordes and Ema Demsar. Towards the epistemology of the non-trivial: Research characteristics connecting quantum mechanics and first-person inquiry. Foundations of Science, 26(1):187–216, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-019-09638-z.

Etesi:2021:ASL

[814] Gábor Etesi. Antimatter and the Second Law of Thermodynamics. Foundations of Science, 26(2):217–224, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09652-6.

Heylighen:2021:FAF

[815] Francis Heylighen and Katarina Petrović. Foundations of ArtScience: Formulating the problem. Foundations of Science, 26(2):225–244, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09660-6.

Tozzi:2021:TAI

[816] Arturo Tozzi and James F. Peters. A topological approach to infinity in physics and biophysics. Foundations of Science, 26(2):245–255, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09674-0.

Mazloom:2021:EPF

[817] Roham Mazloom. Evidence of the physiologic functions of the gastrointestinal tract as a complex system. Foundations of Science, 26 (2):257–274, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09656-2.

Mizrahi:2021:TGS

[818] Moti Mizrahi. Is truth the gold standard of inquiry? A comment on Elgin’s argument against veritism. Foundations of Science, 26 (2):275–280, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09671-3.

Lamza:2021:SPK

[819] Lukasz Lamza. Superorganisms of the protist kingdom: a new level of biological organization. Foundations of Science, 26(2):281–300, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09688-8.

DeRisi:2021:ECN

[820] Vincenzo De Risi. Euclid’s common notions and the Theory of Equivalence. Foundations of Science, 26(2):301–324, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09694-w.

Camina:2021:BOE

[821] Eduardo Camina, Javier Bernacer, and Francisco Guell. Belief operationalization for empirical research in psychological sciences. Foundations of Science, 26(2):325–340, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09722-9.

Pique:2021:ECP

[822] Pilar Piqué. Ernst Cassirer’s philosophy of culture: an economic assessment of scope and limitations. Foundations of Science, 26(2):341– 354, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09678-w.

Zambon:2021:NIR

[823] Alfio Zambon and Mariana Córdoba. Nanomaterials and intertheoretical relations: Macro and nanochemistry as emergent levels. Foundations of Science, 26(2):355–370, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09723-8.

Javanmardi:2021:EPF

[824] Ehsan Javanmardi, Sifeng Liu, and Naiming Xie. Exploring the philosophical foundations of grey systems theory: Subjective processes, information extraction and knowledge formation. Foundations of Science, 26(2):371–404, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09690-0.

Cordoba:2021:BTP

[825] Mariana Córdoba. The biological turn on personal identity: The role of science as a response to children’s appropriation in Argentinian dictatorship (1976–1983). Foundations of Science, 26(2):405– 427, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-019-09628-1.

Yan:2021:BMR

[826] Chunling Yan. Bohmian mechanics: Realism and the “Box” experiment. Foundations of Science, 26(2):429–451, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09661-5.

Nescolarde-Selva:2021:BKF

[827] J. Nescolarde-Selva, J. L. Usó-Doménech, and H. Gash. Belief, knowledge and faith: a logical modal theory. Foundations of Science, 26 (2):453–474, June 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09677-x.

Chen:2021:IEA

[828] Zheng Chen and Yu He. Imitation or efficiency? Aesthetics of artificial intelligence. Foundations of Science, 26(3):475–486, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09682-0.

Dua:2021:RRH

[829] Mikhael Dua. Radical reflection in human sciences, Calvin Schrag’s epistemological proposal. Foundations of Science, 26(3):487–501, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09693-x.

East:2021:IKF

[830] Robert East and Lawrence Ang. Insight: The key to faster progress in science. Foundations of Science, 26(3):503–514, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09692-y.

Valore:2021:OCS

[831] Paolo Valore, M. G. Dainotti, and Oskar Kopczyński. Ontological categorizations and selection biases in cosmology: The case of extra galactic objects. Foundations of Science, 26(3):515–529, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09699-5.

Pisano:2021:IIS

[832] Raffaele Pisano, Paulo Mauricio, and Philippe Vincent. Introduction to IDTC special issue: Joule’s bicentenary history of science, foundations and nature of science. Foundations of Science, 26(3):531– 551, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09680-2.

Agassi:2021:SEP

[833] Joseph Agassi. Science education: Principles. Foundations of Science, 26 (3):553–558, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09643-2.

Pitts:2021:CEM

[834] J. Brian Pitts. Conservation of energy: Missing features in its nature and justification and why they matter. Foundations of Science, 26(3): 559–584, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09657-1.

Drago:2021:JEE

[835] Antonino Drago. Joule’s experiment as an event triggering a formalization of a Baconian science till up to an alternative theory to Newton’s one. Foundations of Science, 26(3):585–605, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09667-z.

Coelho:2021:ECP

[836] R. Lopes Coelho. On the energy concept problem: Experiments and interpretations. Foundations of Science, 26(3):607–624, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09675-z.

Martins:2021:JEH

[837] R. A. Martins and A. P. B. Silva. Joule’s experiments on the heat evolved by metallic conductors of electricity. Foundations of Science, 26(3): 625–701, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09681-1.

Kreinovich:2021:JNC

[838] Vladik Kreinovich and Olga Kosheleva. Joule’s Nineteenth Century energy conservation meta-law and the Twentieth Century physics (quantum mechanics and relativity): Twenty-first century analysis. Foundations of Science, 26(3):703–725, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09659-z.

Pisano:2021:HFP

[839] Raffaele Pisano, Philippe Vincent, and Mateja Ploj Virtic. Historical foundations of physics & applied technology as dynamic frameworks in pre-service STEM. Foundations of Science, 26(3):727–756, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-020-09662-4.

Lusito:2021:SOA

[840] Fabio Lusito. Science outside academies: an Italian case of “Scientific mediation” — from Joule’s seminal experience to Lucio Lombardo Radice’s contemporary attempt. Foundations of Science, 26(3):757– 790, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09665-1.

Bachtold:2021:IJP

[841] Manuel Bächtold. Introducing Joule’s paddle wheel experiment in the teaching of energy: Why and how? Foundations of Science, 26(3): 791–805, September 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09664-2.

Bielecki:2021:SCC

[842] Andrzej Bielecki. The systemic concept of contextual truth. Foundations of Science, 26(4):807–824, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09713-w.

Werth:2021:HWT

[843] Alexander J. Werth and Douglas Allchin. How we think about human nature: Cognitive errors and concrete remedies. Foundations of Science, 26 (4):825–846, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09726-5.

Andersson:2021:DDD

[844] E. Andersson and S. Abdelmalek. Dyscalculia/dyslexia: a dichotomy? Foundations of Science, 26(4):847–858, December 2021. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09698-6.

Boucher:2021:BTR

[845] Sandy C. Boucher. Biological teleology, reductionism, and verbal disputes. Foundations of Science, 26(4):859–880, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09728-3.

DelDin:2021:SMT

[846] Guido Del Din. Some model-theoretic remarks on the Ramsey sentence, with a closer look at Ketland’s argument. Foundations of Science, 26(4): 881–900, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09778-1.

Uzan:2021:MBC

[847] Pierre Uzan. Mind-body connection and causation: Conceptual and experimental advances. Foundations of Science, 26(4):901–915, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09709-6.

Lambare:2021:CLA

[848] Justo Pastor Lambare. Comment on “A loophole of all ”loopholefree“ Bell-type theorems”. Foundations of Science, 26(4):917–924, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09695-9. See [789].

Schinckus:2021:SFI

[849] Christophe Schinckus. The Santa Fe Institute and econophysics: a possible genealogy? Foundations of Science, 26(4):925–945, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09714-9.

Surov:2021:QCT

[850] Ilya A. Surov. Quantum cognitive triad: Semantic geometry of context representation. Foundations of Science, 26(4):947–975, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09712-x.

Mansfield:2021:PSR

[851] Daniel F. Mansfield. Plimpton 322: a study of rectangles. Foundations of Science, 26(4):977–1005, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09806-0.

Barzegar:2021:ZIR

[852] Ali Barzegar, Mostafa Taqavi, and Afshin Shafiee. Zeilinger on information and reality. Foundations of Science, 26(4):1007–1019, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09696-8.

Weber:2021:IUD

[853] Erik Weber, Henk W. de Regt, and Dingmar van Eck. Investigating the unity and disunity of scientific explanation. Foundations of Science, 26(4):1021–1024, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09704-x.

Rice:2021:HRU

[854] Collin Rice and Yasha Rohwer. How to reconcile a unified account of explanation with explanatory diversity. Foundations of Science, 26(4): 1025–1047, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09647-y.

Lawler:2021:MEV

[855] Insa Lawler and Emily Sullivan. Model explanation versus modelinduced explanation. Foundations of Science, 26(4):1049–1074, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09649-1.

Rosales:2021:SET

[856] Alirio Rosales and Adam Morton. Scientific explanation and tradeoffs between explanatory virtues. Foundations of Science, 26(4):1075– 1087, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-019-09645-0.

Gervais:2021:MEC

[857] Raoul Gervais. The multiplicity of explanation in cognitive science. Foundations of Science, 26(4):1089–1104, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09653-5.

Frans:2021:UUE

[858] Joachim Frans. Unificatory understanding and explanatory proofs. Foundations of Science, 26(4):1105–1127, December 2021. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09654-4.

Eckstein:2022:EPP

[859] Michał Eckstein and Paweł Horodecki. The experiment paradox in physics. Foundations of Science, 27(1):1–15, March 2022. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09711-y.

Uso-Domenech:2022:CWB

[860] J. L. Usó-Doménech and J. Nescolarde-Selva. Correction to: What are belief systems? Foundations of Science, 27(1):17, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:

//link.springer.com/article/10.1007/s10699-021-09824-y. See

[565].

Maldonado:2022:LHS

[861] Carlos Eduardo Maldonado and Eduardo Villar Concha. A landmark in the history of science. Foundations of Science, 27(1):19–25, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/

s10699-020-09721-w. See correction [914].

deRonde:2022:ATP

[862] C. de Ronde and C. Massri. Against the tyranny of ‘pure states’ in quantum theory. Foundations of Science, 27(1):27–41, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09720-x.

Cardone:2022:LVT

[863] Fabio Cardone, Gianni Albertini, and Domenico Bassani. Lorentz violation in torsional antenna. Foundations of Science, 27(1):43–55, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09710-z.

Leporini:2022:HCL

[864] Roberto Leporini. Holistic and compositional logics based on the Bertini gate. Foundations of Science, 27(1):57–75, March 2022. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09703-y.

Sanchez-Canizares:2022:FCI

[865] Javier Sánchez-Cañizares. Formal causation in integrated information theory: an answer to the intrinsicality problem. Foundations of Science, 27(1):77–94, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09775-w.

Lemmens:2022:RTA

[866] Pieter Lemmens and Yoni Van Den Eede. Rethinking technology in the Anthropocene: Guest Editors’ introduction. Foundations of Science, 27 (1):95–105, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09772-z.

Scharff:2022:BET

[867] Robert C. Scharff. Before empirical turns and transcendental inquiry: Pre-philosophical considerations. Foundations of Science, 27 (1):107–124, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09734-5.

DePreester:2022:STI

[868] Helena De Preester. Subjectivity and transcendental illusions in the Anthropocene. Foundations of Science, 27(1):125–140, March 2022. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

https://link.springer.com/article/10.1007/s10699-020-09733-
6.

Hui:2022:CE

[869] Yuk Hui. For a cosmotechnical event. Foundations of Science, 27 (1):141–154, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09729-2.

Smith:2022:TEP

[870] Dominic Smith. Taking exception: Philosophy of technology as a multidimensional problem space. Foundations of Science, 27(1):155–170, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09735-4.

Lemmens:2022:TTB

[871] Pieter Lemmens. Thinking technology big again. reconsidering the question of the transcendental and ‘technology with a capital T’ in the light of the Anthropocene. Foundations of Science, 27(1):171–187, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09732-7.

Wellner:2022:DII

[872] Galit Wellner. Digital imagination: Ihde’s and Stiegler’s concepts of imagination. Foundations of Science, 27(1):189–204, March 2022. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

https://link.springer.com/article/10.1007/s10699-020-09737-

Zwart:2022:PDT

[873] Hub Zwart. Practicing dialectics of technoscience during the Anthropocene. Foundations of Science, 27(1):205–224, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:

//link.springer.com/article/10.1007/s10699-020-09738-1. See

commentary [921].

VanDenEede:2022:TTN

[874] Yoni Van Den Eede. Thing-transcendentality: Navigating the interval of “technology” and “Technology”. Foundations of Science, 27(1): 225–243, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09749-y.

Ihde:2022:HIG

[875] Don Ihde. From Heideggerian industrial gigantism to nanoscale technologies. Foundations of Science, 27(1):245–257, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09731-8.

Coeckelbergh:2022:ETL

[876] Mark Coeckelbergh. Earth, technology, language: a contribution to holistic and transcendental revisions after the artifactual turn. Foundations of Science, 27(1):259–270, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09730-9.

Stiegler:2022:OTC

[877] Bernard Stiegler. The ordeal of truth: Causes and quasi-causes in the Entropocene. Foundations of Science, 27(1):271–280, March 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09736-3.

Belu:2022:BPT

[878] Dana S. Belu. Back to the phenomenology of technical life. Foundations of Science, 27(2):281–285, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09759-w.

deWee:2022:CPA

[879] Guswin de Wee. Comparative policy analysis and the science of conceptual systems: a candidate pathway to a common variable. Foundations of Science, 27(2):287–304, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09782-5.

Glazebrook:2022:WNP

[880] Tricia Glazebrook. What now in philosophy of technology? Ethics, time, and Poiêsis in crisis thinking. Foundations of Science, 27(2): 305–310, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09760-3.

Agnati:2022:MRS

[881] Luigi F. Agnati, Deanna Anderlini, and Guido Maura. Man is a “Rope” stretched between virosphere and humanoid robots: On the urgent need of an ethical code for ecosystem survival. Foundations of Science, 27 (2):311–325, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09796-z.

Franklin:2022:MSN

[882] James Franklin. Mathematics as a science of non-abstract reality: Aristotelian realist philosophies of mathematics. Foundations of Science, 27(2):327–344, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09786-1.

Botin:2022:WHM

[883] Lars Botin. Why Heidegger makes sense in contemporary philosophy of technology. Foundations of Science, 27(2):345–350, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09755-0.

Furnell:2022:FDT

[884] Miles W. Furnell. FRAME dynamics: a theory of general evolution. Foundations of Science, 27(2):351–370, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09795-0.

Melogno:2022:EIH

[885] Pablo Melogno. From externalism to internalism: The historiographical development of Thomas Kuhn. Foundations of Science, 27(2): 371–385, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09801-5.

Ruyant:2022:RBM

[886] Quentin Ruyant. On the relationship between modelling practices and interpretive stances in quantum mechanics. Foundations of Science, 27 (2):387–405, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09774-x.

Liberati:2022:RED

[887] Nicola Liberati. Reflections on an externalized digital imagination. Foundations of Science, 27(2):407–410, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09768-9.

Blok:2022:EMA

[888] Vincent Blok. The morendo of the Anthropocene. Foundations of Science, 27(2):411–415, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.com/

article/10.1007/s10699-020-09763-0.

Scharff:2022:BOT

[889] Robert C. Scharff. Before one takes empirical or transcendental positions. Foundations of Science, 27(2):417–425, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09744-3.

Luk:2022:WIR

[890] Robert W. P. Luk. Why is information retrieval a scientific discipline? Foundations of Science, 27(2):427–453, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09685-x.

Saint-Mont:2022:ILA

[891] Uwe Saint-Mont. Induction: a logical analysis. Foundations of Science, 27(2):455–487, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09683-z.

Tozzi:2022:BIO

[892] Arturo Tozzi. Are borders inside or outside? Foundations of Science, 27(2):489–505, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09708-7.

Lally:2022:PPT

[893] Róisín Lally. Post-phenomenology, transduction, and speculative fabulations. Foundations of Science, 27(2):507–514, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09765-y.

Schick:2022:TA

[894] Johannes F. M. Schick. On technical alterity. Foundations of Science, 27(2):515–520, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09771-0.

Barker:2022:TAM

[895] Timothy Barker. The transcendental and the agonistic: a media philosophy perspective. Foundations of Science, 27(2):521–525, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09761-2.

Delhotel:2022:UTE

[896] Jean-Michel Delhôtel. Unspeakable transport — what quantum teleportation might be, and what it more probably is. Foundations of Science, 27(2):527–548, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09779-0.

Besmer:2022:TET

[897] Kirk M. Besmer. On turning away from “the empirical turn” (but not necessarily towards the transcendental). Foundations of Science, 27(2):549–554, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09757-y.

Uso-Domenech:2022:DHL

[898] J. L. Usó-Doménech, J. A. Nescolarde-Selva, and H. Gash. Dialectical Hegelian logic and physical quantity and quality. Foundations of Science, 27(2):555–572, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09790-5.

Pandilovski:2022:PLT

[899] Melentie Pandilovski. Perspectives on living and thinking vectors of the Anthropocene. Foundations of Science, 27(2):573–580, June 2022. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

https://link.springer.com/article/10.1007/s10699-020-09770-
1.

DePreester:2022:OTH

[900] Helena De Preester. The object that technology is not and how we can relate to it. Foundations of Science, 27(2):581–585, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09743-4.

Blaasjo:2022:OIP

[901] Viktor Blåsjö. Operationalism: an interpretation of the philosophy of Ancient Greek geometry. Foundations of Science, 27(2):587–708, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09791-4.

Rosenberger:2022:LD

[902] Robert Rosenberger. Localizations of dystopia. Foundations of Science, 27(2):709–715, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09756-z.

DeGroot:2022:PSI

[903] T. M. M. De Groot and P. Meurs. Preface of the special issue: Worldviews and health-related stigma. Foundations of Science, 27 (2):717–718, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09842-4.

Reimer-Kirkham:2022:MAH

[904] Sheryl Reimer-Kirkham, Barbara Astle, and Emma Strobell. Mothering, albinism and human rights: The disproportionate impact of healthrelated stigma in Tanzania. Foundations of Science, 27(2):719–740, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09701-0.

Stroeken:2022:SBN

[905] Koen Stroeken. Stigmatized bodies near Lake Victoria: a cultural analysis of institutions. Foundations of Science, 27(2):741–751, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09684-y.

Roozen:2022:USS

[906] Sylvia Roozen, Sarah E. Stutterheim, and Leopold M. G. Curfs. Understanding the social stigma of fetal alcohol spectrum disorders: From theory to interventions. Foundations of Science, 27(2):753–771, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09676-y.

Forsyth:2022:CSR

[907] Miranda Forsyth and Philip Gibbs. Causal stories and the role of worldviews in analysing responses to sorcery accusations and related violence. Foundations of Science, 27(2):773–784, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09727-4.

DeGroot:2022:CVE

[908] T. M. M. De Groot, P. Meurs, and R. M. H. Peters. Contact versus education: an explorative comparison between the contact and education strategy considering albinism related stigma in Tanzanian high schools. Foundations of Science, 27(2):785–803, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09788-z.

Hartog:2022:USR

[909] Kim Hartog, Ruth M. H. Peters, and Mark J. D. Jordans. Understanding stigmatisation: Results of a qualitative formative study with adolescents and adults in DR Congo. Foundations of Science, 27 (2):805–828, June 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09706-9.

Ginoux:2022:AED

[910] Jean-Marc Ginoux. Albert Einstein and the doubling of the deflection of light. Foundations of Science, 27(3):829–850, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-021-09783-4.

Ihde:2022:PET

[911] Don Ihde. Postphenomenology, the empirical turn and “Transcendentality”. Foundations of Science, 27(3):851–854, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09741-6.

Bell:2022:BRE

[912] Andrew Bell, Bryn Davies, and Habib Ammari. Bernhard Riemann, the ear, and an atom of consciousness. Foundations of Science, 27(3): 855–873, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09813-1.

Muller:2022:LGC

[913] Christopher John Müller. Are language games also confidence tricks? Technology as embodied power and collective disempowerment. Foundations of Science, 27(3):875–880, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09752-3.

Maldonado:2022:CLH

[914] Carlos Eduardo Maldonado and Eduardo Villar Concha. Correction to: A landmark in the history of science. Foundations of Science, 27(3): 881–883, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09776-9. See [861].

Gornitz:2022:QCE

[915] Thomas Görnitz and Uwe Schomäcker. Quantumbit cosmology explains effects of rotation curves of galaxies. Foundations of Science, 27(3): 885–914, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09808-y.

Wiltse:2022:MIR

[916] Heather Wiltse. Maneuvering in the interval: Reflections on immanent entanglements. Foundations of Science, 27(3):915–920, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09766-x.

Krol:2022:OPM

[917] Zbigniew Król and Józef Lubacz. Ontological pluralism and multiquantificational ontology. Foundations of Science, 27(3):921–940, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-021-09810-4.

Gunton:2022:GTO

[918] Richard M. Gunton, Marinus D. Stafleu, and Michael J. Reiss. A general theory of objectivity: Contributions from the reformational philosophy tradition. Foundations of Science, 27(3):941–955, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09809-x.

Ippoliti:2022:WFN

[919] Emiliano Ippoliti. Why finance needs philosophy (and vice versa): Some epistemic and methodological issues. Foundations of Science, 27(3): 957–974, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09804-2.

Romele:2022:TTS

[920] Alberto Romele. The transcendental of technology is said in many ways. Foundations of Science, 27(3):975–980, September 2022. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09758-x.

Rheinberger:2022:CPD

[921] Hans-Jörg Rheinberger. Commentary to “Practicing dialectics of technoscience during the anthropocene” by Hub Zwart. Foundations of Science, 27(3):981–985, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09773-y. See [873].

Delgado:2022:OSM

[922] Henrique Estides Delgado. On the ontological status of mechanisms and processes in the social world. Foundations of Science, 27(3):987– 1000, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09787-0.

Wilson:2022:DAC

[923] Alexander Wilson. Desperate acts and compromises. Foundations of Science, 27(3):1001–1008, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09764-z.

Zwier:2022:AE

[924] Jochem Zwier. Accepting the exceptional? Foundations of Science, 27(3):1009–1014, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.com/

article/10.1007/s10699-020-09762-1.

Gremmen:2022:DMP

[925] Bart Gremmen. Dialectical methodology of the praxis of biology. Foundations of Science, 27(3):1015–1020, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09769-8.

Lemmens:2022:MBS

[926] Pieter Lemmens. In memory of Bernard Stiegler (1 April 1952– 5 August 2020). Foundations of Science, 27(3):1021–1028, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09746-1.

Lara:2022:PMC

[927] Juan A. Lara, Juan Pazos, Aurea Anguera de Sojo, and Shadi Aljawarneh. The paternity of the modern computer. Foundations of Science, 27(3):1029–1040, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09797-y.

Zwart:2022:EHT

[928] Hub Zwart. The empirical and the holistic turn: a Hegelian dialectics of technoscience revisited. Foundations of Science, 27(3):1041– 1048, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09748-z.

Neyrat:2022:CTD

[929] Frédéric Neyrat. Cosmos and technology (Dasein’s planetary condition). Foundations of Science, 27(3):1049–1054, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09753-2.

Croce:2022:HDE

[930] Mariano Croce and Emilia Margoni. How does an entity acquire identity? Reassembling relativistic physics with actor–network theory. Foundations of Science, 27(3):1055–1071, September 2022. CODEN FOSCFI.

ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09800-6.

Sterpetti:2022:MEE

[931] Fabio Sterpetti. Mathematical explanations in evolutionary biology or naturalism? A challenge for the statisticalist. Foundations of Science, 27(3):1073–1105, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.com/

article/10.1007/s10699-021-09818-w.

Pinillos:2022:NPD

[932] Alger Sans Pinillos. Neglected pragmatism: Discussing abduction to dissolute classical dichotomies. Foundations of Science, 27(3):1107– 1125, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09817-x.

Pisano:2022:CFR

[933] Raffaele Pisano and Paolo Bussotti. Conceptual frameworks on the relationship between physics–mathematics in the Newton Principia Geneva Edition (1822). Foundations of Science, 27(3):1127–1182, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09820-2.

Alvarado:2022:CSS

[934] Ramón Alvarado. Computer simulations as scientific instruments. Foundations of Science, 27(3):1183–1205, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09812-2.

Jarnicki:2022:ESN

[935] Paweł Jarnicki. Stimmung/Nastrój as content of modern science: On musical metaphors in Ludwik Fleck’s theory of thought styles and thought collectives. Foundations of Science, 27(3):1207–1228, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09792-3.

Falk:2022:POA

[936] Johannes Falk, Edwin Eichler, Katja Windt, and Marc-Thorsten Hütt. Physics is organized around transformations connecting contextures in a polycontextural world. Foundations of Science, 27(3):1229–1251, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-021-09814-0.

Mohrhoff:2022:QO

[937] U. J. Mohrhoff. A QBist ontology. Foundations of Science, 27(3):1253– 1277, September 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09802-4.

Vaccari:2022:CTB

[938] Andrés Vaccari. Cosmotechnical thought between substantivism and the empirical turn. Foundations of Science, 27(4):1279–1284, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09754-1.

Dolev:2022:WPW

[939] Yuval Dolev. Why physics is not wrong on temporal directionality, and why this is not necessarily good news for physicalism. Foundations of Science, 27(4):1285–1300, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09816-y.

VanDenEede:2022:EBS

[940] Yoni Van Den Eede. Exploring in between small and big: On algorithmic mediation, the importance of relation, and the terms of debate. Foundations of Science, 27(4):1301–1305, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09750-5.

Lemmens:2022:TTD

[941] Pieter Lemmens. Technologizing the transcendental, not discarding it. Foundations of Science, 27(4):1307–1315, December 2022. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-020-09742-5.

Smith:2022:NES

[942] Dominic Smith. Necessity, entailment, shared agonism. Foundations of Science, 27(4):1317–1325, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09745-2.

Bielecki:2022:IES

[943] Andrzej Bielecki and Michael Schmittel. The information encoded in structures: Theory and application to molecular cybernetics. Foundations of Science, 27(4):1327–1345, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09830-8.

Zwart:2022:LDK

[944] Hub Zwart. Lacan’s dialectics of knowledge production: The four discourses as a detour to Hegel. Foundations of Science, 27(4):1347– 1370, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09832-6.

Campbell:2022:ULM

[945] Melinda Campbell. Unfolding the layers of mind and world: Wellner’s posthuman digital imagination. Foundations of Science, 27(4): 1371–1380, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09767-w.

Etesi:2022:PBH

[946] Gábor Etesi. Primordial black holes from collapsing antimatter. Foundations of Science, 27(4):1381–1400, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09831-7.

deBoer:2022:DOW

[947] Bas de Boer. Does an ontic whole exist?: Conditions of possibility and technology use. Foundations of Science, 27(4):1401–1407, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09751-4.

Hui:2022:SBC

[948] Yuk Hui. This strange being called the Cosmos. Foundations of Science, 27(4):1409–1414, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09740-7.

Coeckelbergh:2022:DTH

[949] Mark Coeckelbergh. Defamiliarizing technology, habituation, and the need for a structuralist approach. Foundations of Science, 27(4):1415– 1420, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09739-0.

Rao:2022:UEP

[950] Mithun Bantwal Rao. The use of examples in philosophy of technology. Foundations of Science, 27(4):1421–1443, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-021-09819-9.

Wellner:2022:DIF

[951] Galit Wellner. Digital imagination, fantasy, AI art. Foundations of Science, 27(4):1445–1451, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09747-0.

Kotowski:2022:DPD

[952] Mateusz Kotowski and Krzysztof Szlachcic. Deconstructing the phantom: Duhem and the scientific realism debate. Foundations of Science, 27(4): 1453–1475, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09823-z.

Tozzi:2022:GMI

[953] Arturo Tozzi, Alexander Yurkin, and James F. Peters. A geometric milieu inside the brain. Foundations of Science, 27(4):1477–1488, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09798-x.

Seddon:2022:NCA

[954] Peter B. Seddon. Nature chose abduction: Support from brain research for Lipton’s theory of inference to the best explanation. Foundations of Science, 27(4):1489–1505, December 2022. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09811-3.

Cordovil:2023:ROS

[955] João L. Cordovil, Gil C. Santos, and John Symons. Reconciling ontic structural realism and ontological emergence. Foundations of Science, 28(1):1–20, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09828-8.

deOliveira:2023:QBD

[956] Michael de Oliveira and Luis Soares Barbosa. Quantum Bayesian decision–making. Foundations of Science, 28(1):21–41, March 2023. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

https://link.springer.com/article/10.1007/s10699-021-09781-
6.

Beltran:2023:QBE

[957] Lester Beltran. Quantum Bose–Einstein statistics for indistinguishable concepts in human language. Foundations of Science, 28(1):43– 55, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-021-09794-1.

Ginoux:2023:FPI

[958] Jean-Marc Ginoux and Franck Jovanovic. Frisch’s propagation–impulse model: a comprehensive mathematical analysis. Foundations of Science, 28(1):57–84, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09827-9.

Heylighen:2023:ESB

[959] Francis Heylighen. Entanglement, symmetry breaking and collapse: Correspondences between quantum and self-organizing dynamics. Foundations of Science, 28(1):85–107, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09780-7.

Giuntini:2023:CTS

[960] Roberto Giuntini, Pierluigi Graziani, Giuseppe Sergioli, and Sandro Sozzo. Connecting things in the setting of foundations and philosophy of science: Introduction to the special issue. Foundations of Science, 28 (1):109–114, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09845-1.

Salibra:2023:CLE

[961] A. Salibra, A. Bucciarelli, A. Ledda, and F. Paoli. Classical logic with n truth values as a symmetric many-valued logic. Foundations of Science, 28(1):115–142, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09697-7.

Boem:2023:CCE

[962] F. Boem, S. Bonzio, B. Osimani, and A. Sacco. The Cochrane case: an epistemic analysis on decision-making and trust in science in the age of information. Foundations of Science, 28(1):143–158, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09668-y.

Tarozzi:2023:NTC

[963] Gino Tarozzi and Giovanni Macchia. No-thing and causality in realistic non–standard interpretations of the quantum mechanical wave function: Ex Nihilo Aliquid? Foundations of Science, 28(1):159–184, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09807-z.

Conti:2023:RPE

[964] Ludovica Conti. Russell’s Paradox and Free zig zag solutions. Foundations of Science, 28(1):185–203, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09689-7.

Negri:2023:AAL

[965] Sara Negri and Edi Pavlović. Alternative axiomatization for logics of agency in a G3 calculus. Foundations of Science, 28(1):205–224, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-021-09784-3.

Tagliaferri:2023:RCO

[966] Mirko Tagliaferri. Reviewing the case of online interpersonal trust. Foundations of Science, 28(1):225–254, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09836-2.

Chis-Ciure:2023:FHP

[967] Robert Chis-Ciure and Francesco Ellia. Facing up to the hard problem of consciousness as an integrated information theorist. Foundations of Science, 28(1):255–271, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-020-09724-7.

Florio:2023:WMK

[968] Ciro De Florio and Vincenzo Fano. Williamson on the margins of knowledge: a criticism. Foundations of Science, 28(1):273–285, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09715-8.

Friedman:2023:MPH

[969] Sy-David Friedman and Claudio Ternullo. Maximality principles in the hyperuniverse programme. Foundations of Science, 28(1):287–305, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09707-8.

Boniolo:2023:MBM

[970] Giovanni Boniolo, Marcello D’Agostino, Mario Piazza, and Gabriele Pulcini. Molecular biology meets logic: Context-sensitiveness in focus. Foundations of Science, 28(1):307–325, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09789-y.

Giunti:2023:DPT

[971] Marco Giunti. Dynamical phenomena and their models: Truth and empirical correctness. Foundations of Science, 28(1):327–375, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09700-1.

Alai:2023:SRM

[972] Mario Alai. Scientific realism, metaphysical antirealism and the no miracle arguments. Foundations of Science, 28(1):377–400, March 2023. CO-DEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL

https://link.springer.com/article/10.1007/s10699-020-09691-
z.

Bianchini:2023:NDA

[973] Francesco Bianchini. A new definition of “Artificial” for two artificial sciences. Foundations of Science, 28(1):401–417, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-021-09799-w.

Fazio:2023:FCC

[974] D. Fazio, A. Ledda, and F. Paoli. On Finch’s conditions for the completion of orthomodular posets. Foundations of Science, 28(1):

419–440, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09702-z.

Nathan:2023:CVC

[975] Marco J. Nathan. Causation vs. causal explanation: Which is more fundamental? Foundations of Science, 28(1):441–454, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-020-09672-2.

Corti:2023:SRR

[976] Alberto Corti. Scientific realism without reality? What happens when metaphysics is left out. Foundations of Science, 28(1):455–475, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-020-09705-w.

Muller:2023:IQP

[977] F. A. Muller. The influence of quantum physics on philosophy. Foundations of Science, 28(1):477–488, March 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-020-09725-6.

Anta:2023:IMC

[978] Javier Anta. Ignorance, milk and coffee: Can epistemic states be causally-explanatorily relevant in statistical mechanics? Foundations of Science, 28(2):489–505, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09803-3.

Uzan:2023:BRR

[979] Pierre Uzan. Bayesian rationality revisited: Integrating order effects. Foundations of Science, 28(2):507–528, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09838-0.

Yan:2023:URI

[980] Chunling Yan. Underdetermination: a realist interpretation of quantum mechanics and Bohmian mechanics. Foundations of Science, 28 (2):529–550, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09839-z.

Drago:2023:DBE

[981] Antonino Drago. Dirac’s book The Principles of Quantum Mechanics as an alternative way of organizing a theory. Foundations of Science, 28(2):551–574, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09835-3.

Krol:2023:RWQ

[982] Jerzy Król, Krzysztof Bielas, and Torsten Asselmeyer-Maluga. Random world and quantum mechanics. Foundations of Science, 28(2): 575–625, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09852-2.

Uskokovic:2023:CTN

[983] Evangelina Uskoković, Theo Uskoković, Victoria Wu, and Vuk Uskoković. Chutes too narrow: The Brazil nut effect and the blessings of the fall. Foundations of Science, 28(2):627–708, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09892-8.

Fortin:2023:OSM

[984] Sebastian Fortin, Martín Labarca, and Olimpia Lombardi. On the ontological status of molecular structure: Is it possible to reconcile molecular chemistry with quantum mechanics? Foundations of Science, 28 (2):709–725, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09834-4.

Cabaret:2023:EPW

[985] D.-M. Cabaret, T. Grandou, G.-M. Grange, and E. Perrier. Elementary particles: What are they? Substances, elements and primary matter. Foundations of Science, 28(2):727–753, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-021-09826-w.

Kim:2023:IBI

[986] Namjoong Kim. Imprecise Bayesianism and inference to the best explanation. Foundations of Science, 28(2):755–781, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-022-09841-5.

Bosschaert:2023:EET

[987] Mariska Thalitha Bosschaert and Vincent Blok. The ‘empirical’ in the empirical turn: a critical analysis. Foundations of Science, 28 (2):783–804, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09840-6.

Ginoux:2023:CFP

[988] Jean-Marc Ginoux and Franck Jovanovic. Correction to Frisch’s propagation–impulse model: a comprehensive Mathematical Analysis. Foundations of Science, 28(3):805–807, September 2023. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-022-09895-5.

Alamuti:2023:CRE

[989] Masoud Mohammadi Alamuti. Critical rationalism: an epistemological critique. Foundations of Science, 28(3):809–840, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09847-z.

Mykhailov:2023:STI

[990] Dmytro Mykhailov and Nicola Liberati. A study of technological intentionality in C++ and generative adversarial model: Phenomenological and postphenomenological perspectives. Foundations of Science, 28(3): 841–857, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09833-5.

Froes:2023:ERB

[991] Maira M. Fróes and Agamenon R. E. Oliveira. Entangling and rupture of body and mind for building of the modern science: Lessons from da Vinci and Descartes. Foundations of Science, 28(3):859– 884, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09874-w.

Arroyo:2023:PQM

[992] Raoni Wohnrath Arroyo and Jonas R. Becker Arenhart. The powers of quantum mechanics: a metametaphysical discussion of the “Logos Approach”. Foundations of Science, 28(3):885–910, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09837-1.

Uso-Domenech:2023:PAR

[993] J. L. Usó-Doménech, J. A. Nescolarde-Selva, and H. Gash. Paradoxical aspects of the Russellian conception of existence. Foundations of Science, 28(3):911–925, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-021-09822-0.

Nemati:2023:ACS

[994] Farshad Nemati. Anthropomorphism in the context of scientific discovery: Implications for comparative cognition. Foundations of Science, 28 (3):927–945, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09821-1.

Bussotti:2023:PFA

[995] Paolo Bussotti. Parmenides, the founder of abstract geometry: Enriques interpreter of the eleatic thought. Foundations of Science, 28(3): 947–975, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09854-0.

Rigat:2023:WPI

[996] Fabio Rigat. Why probability isn’t magic. Foundations of Science, 28(3): 977–985, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-021-09815-z.

Blok:2023:OTB

[997] Vincent Blok. The ontology of technology beyond anthropocentrism and determinism: The role of technologies in the constitution of the (post)Anthropocene world. Foundations of Science, 28(3):987–1005, September 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL https://link.springer.com/article/10.
1007/s10699-022-09829-1.

Ferro:2023:PDK

[998] Floriana Ferro and Luca Taddio. Phenomenology and digital knowledge: Introduction to the special issue. Foundations of Science, 28(4): 1007–1010, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-023-09905-0.

Cantone:2023:SBP

[999] Damiano Cantone. The simulated body: a preliminary investigation into the relationship between neuroscientific studies, phenomenology and virtual reality. Foundations of Science, 28(4):1011–1020, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09849-x.

Taddio:2023:PDW

[1000] Luca Taddio. From perception to the digital world: phenomenological observations. Foundations of Science, 28(4):1021–1034, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09846-0.

Buongiorno:2023:CAE

[1001] Federica Buongiorno. Can algorithms be embodied? A phenomenological perspective on the relationship between algorithimic thinking and the life–world. Foundations of Science, 28(4):1035–1045, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09855-z.

Zhok:2023:PC

[1002] Andrea Zhok. Phenomenology and complexity. Foundations of Science, 28(4):1047–1058, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-022-09865-x.

Lanfredini:2023:DAP

[1003] Roberta Lanfredini. Digital and analogue phenomenology. Foundations of Science, 28(4):1059–1070, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09867-9.

Ferro:2023:PRD

[1004] Floriana Ferro. Perceptual relations in digital environments. Foundations of Science, 28(4):1071–1084, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09853-1.

Possati:2023:TPS

[1005] Luca M. Possati. From Turing to Peirce. a semiotic interpretation of computation. Foundations of Science, 28(4):1085–1110, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/

s10699-022-09878-6. See correction [1010].

Cali:2023:PES

[1006] Carmelo Calì. Philosophical, experimental and synthetic phenomenology: The study of perception for biological, artificial agents and environments. Foundations of Science, 28(4):1111–1124, December 2023. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-022-09869-7.

Osler:2023:SEO

[1007] Lucy Osler and Dan Zahavi. Sociality and embodiment: Online communication during and after Covid-19. Foundations of Science, 28(4): 1125–1142, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09861-1.

Cusano:2023:CSN

[1008] Nicoletta Cusano. Cobot and sobot: For a new ontology of collaborative and social robots. Foundations of Science, 28(4):1143–1155, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09860-2.

Tagliagambe:2023:PDW

[1009] Silvano Tagliagambe. Phenomenology and the digital world: Problems and perspectives. Foundations of Science, 28(4):1157–1174, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09863-z.

Possati:2023:CTP

[1010] Luca M. Possati. Correction: From Turing to Peirce. a semiotic interpretation of computation. Foundations of Science, 28(4):1175, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09884-8. See [1005].

Harman:2023:MIO

[1011] Graham Harman. Merely intentional objects: a defense. Foundations of Science, 28(4):1177–1183, December 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-022-09862-0.

Ceccarelli:2024:ISI

[1012] Marco Ceccarelli and Rafael López-García. Introduction to the special issue: Findings of history of mechanism science. Foundations of Science, 29(1):1–4, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-023-09911-2.

Jauregui-Correa:2024:ASM

[1013] Juan Carlos Jauregui-Correa. Analysis of a stamp mill of Mexico’s antique mines. Foundations of Science, 29(1):5–17, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https:/

/link.springer.com/article/10.1007/s10699-023-09898-w.

Rubio:2024:AFT

[1014] H. Rubio, A. Bustos, C. Castejon, and J. Meneses. Analysis of the first treatise on machine elements: Codex Madrid I. Foundations of Science, 29(1):19–40, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09896-4.

Dorado-Vicente:2024:MWU

[1015] R. Dorado-Vicente, R. López-García, J. M. Quero-Nieves, and G. Medina-Sánchez. Mechanical wits used in the America colonization: Engineering assessment. Foundations of Science, 29(1):41–58, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471

(electronic). URL https://link.springer.com/article/10.1007/
s10699-022-09877-7.

Ambrosetti:2024:FRB

[1016] N. Ambrosetti. Fighting with rotating blades, boomerangs, and crushing punches: a history of mecha from a robotics point of view. Foundations of Science, 29(1):59–85, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.springer.

com/article/10.1007/s10699-022-09887-5.

Seyhan:2024:EHP

[1017] Irem Aslan Seyhan. The early history of the pulleys and crane systems. Foundations of Science, 29(1):87–103, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09888-4.

Medina-Sanchez:2024:HAS

[1018] G. Medina-Sánchez, J. Moreno-Buesa, R. Dorado-Vicente, and R. López-García. Historical approach and scale reconstruction of two medieval mechanisms from “The Book of Secrets”. Foundations of Science, 29 (1):105–124, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09885-7.

Mottola:2024:NHE

[1019] Giovanni Mottola and Marco Cocconcelli. Nomograms in the history and education of machine mechanics. Foundations of Science, 29(1): 125–155, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09890-w.

Navarro-Arcas:2024:MSA

[1020] A. Navarro-Arcas, S. M. Marco Lozano, and Emilio Velasco-Sánchez. Mechanical and structural artefacts used in “The Mystery of Elche”. Foundations of Science, 29(1):157–183, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09893-7.

Fortes-Garrido:2024:CAW

[1021] J. C. Fortes-Garrido, A. M. Rodríguez-Pérez, J. A. Hernández-Torres, J. J. Caparrós-Mancera, J. M. Dávila-Martín, and J. Castilla-Gutiérrez. Comparative analysis of water extraction mechanism in Roman mines. Foundations of Science, 29(1):185–203, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://link.

springer.com/article/10.1007/s10699-022-09891-9.

Gallozzi:2024:BMH

[1022] Arturo Gallozzi and Rodolfo Maria Strollo. Between mechanics and harmony: The drawing of Lissajous curves. Foundations of Science, 29 (1):205–224, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-023-09906-z.

Zrnic:2024:HBE

[1023] Nenad Zrnić, Milos Dordević, and Vlada Gasić. Historical background and evolution of belt conveyors. Foundations of Science, 29 (1):225–255, March 2024. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-022-09894-6.